A Complete Bibliography of Publications in Biometrika for the decade 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

(0, ∞)^{p} [WS99]. C_{p} [FS97]. χ^{2} [JT91]. D [BS94, DW98]. E [DK91, DH94]. F [JT91]. ∞ [PS97b]. J [LTCR90]. k [Dou97, Lam98]. l_{q} [OS93]. M [Ber94, MR93b, Zho92]. ARMA(p, q) [KW97]. n^{−1} [CdPF91]. P [CGS95, BC91b, MMS92]. p^{∗} [Vid95]. t [DDM94, FS99, HSH96, JT91, KTDMB97, PS96].

-distributed [KTDMB97]. -distribution [PS96]. -Estimation [Zho92]. -Estimators [Ber94, MR93b]. -Formula [Vid95]. -Optimal [BS94, DW98, DK91, DH94]. -Solution [HSH96, DDM94]. -Spherical [OS93]. -Values [CGS95, BC91b]. -variate [MMS92].

2000a [DM99].

73 [Dav90b, Fin90, Roc90]. 74 [Pos90]. 76 [CP91, Hal90a, ST91, TKK91]. 77 [FC91, MM99, Rud91, Sch94]. 78 [RB92, Tho93]. 79 [Gle96, JS93, PC93].

80 [Fir95, RLS94, Zel96]. 81 [GT95, Lie96, McC96, PP95]. 83 [BT99, MKGL98]. 84 [Kus99, MO05, UM99]. 85 [DM99]. 86 [Kos01]. 87i [Dav90b]. 88b [Roc90]. 88e [Fin90].

90f [CP91, Pos90]. 90g [Hal90a]. 91a [Sch94]. 91c [ST91, TKK91]. 91f [FC91, Rud91]. 91i [MM99]. 92k [Tho93]. 93a [RB92]. 93f [Gle96]. 93j [PC93]. 94a [JS93]. 94b [Fir95]. 94e [Zel96]. 94g [RLS94]. 95a [PP95]. 96d [McC96]. 96f [Lie96]. 98a [BT99]. 98d [MKGL98]. 98k [Kus99]. 99c [UM99].

Aberration [WZ93]. Abilities [CM99]. Absolute [Mor92]. Abundance [BC98]. accelerate [LRW98]. Accelerated [Jon97, Lin92, LWY98, Rob92, RT92a]. Accept [Per99]. Accept-Reject [Per99]. Accommodate [Gri95]. Accommodation [Sha90]. Accounting [Raf96]. Accuracy [TLP95]. Accurate [DF91, DMY92, DE92, Llo96a, WC99a]. Adaptation [DBL97, DJ94]. Adaptive [Büh99, Don91, HT98, MS97, Tho91b]. Adding [MO97, MO05]. Additive [HK96, KL98, LY94, LOY98, LH96, Lin97b, MS99, MS94, So99, YB97]. Additivity [CLT95]. Addressing [FR99]. Adequacy [CB97, Gli99, KW97, Li91b]. Adjusted [BNC94, BN95a, KS97, ZT97, BNM93]. Adjusting [Alh90, GW98, Lin92]. Adjustment [Cor98, Kak96, MC91, Sev99, WJU97, Won91]. Adjustments [DS93, SA93]. Affects [FK96]. Affine [BMM95]. Affine-Resolvable [BMM95]. After [CW96, HS97, Woo92, Zha92]. Against [ALP98, LL92]. Age [Ahm98, BM93, EH90]. Age-Specific [BM93]. Aggregate [Car97, PS95b]. Aggregation [ZHH^{+}98]. Ahmad [Kum93]. AIC [FS97, HT91, HT98]. AIDS [BM93, Din96, GLV99, Vis96]. Akaike [BN95b]. Akaike-Type [BN95b]. Algorithm [AA94, BB99, DMS98, LR94, Luc94b, Mal95, MR93a, NZ99, SA93, KTDMB97, LRW98]. Algorithms [JJL91, RT96]. Aliased [Wu93]. All-or-None [FR99]. Allocation [GK93, RZ91, Wan91]. Allows [GH90a]. Alternating [CZD93]. Alternative [MS90a, MR99, MW90, Rin96]. Alternatives [ALP98, LL92, MW92, Yao93]. Amendment [ST91]. Amendments [Ano04, Azz97, BT99, Bly95, CP91, Dav90b, DM99, Fin90, Fir95, FC91, GT95, Hal90a, JS93, Kos01, Kus99, Lie96, MM99, MKGL98, MO05, McC96, PP95, Pos90, PC93, RLS94, RB92, Roc90, Sch94, Tar90, Tho93, TKK91, UM99, Whi97, Wis90, Zel96]. Amount [Wan99]. Analyses [HL99, JVDW94, KW94, Smi91, WSL90, XL99]. Analysing [FL93]. Analysis [AC95, AG97, BT95, Cha91a, CS98, CL99b, CWY95, CG98, CW96, CV91b, Dro98, DM92b, ESM95, FK98, FR99, FGS91, Fuk99, Gam92, GKR98, Gle92, GR95, Gra94b, Gri95, HFC90, Jun96, KB90, KBB90, KR92, Lee90, Lev90, LY94, MW94, MD95b, MKL97, MB92b, Mor91, Mor93, MW90, NZ99, Nie99a, Oma91, PD99, Pat91, Pau90, PS95a, PH97, RI93, Rin96, RV91, Sam97, Sch93, SP96, SJDT93, Sch90b, Sch97, SP97, Shi97, STM93, ST98, TLH98, Wan96, WYL90, WW93, YWH98, ZH91, Zha94, ZHH^{+}98, BAM95, Bin97, BHB92a, Dua93, Gle96, Mij90a]. Analytical [DMY93]. Ancillarity [Sev93]. Application [AM94, Azz94, Azz97, BHD98, BB92, DY91, DBL97, DGS94, Din96, EGS96, FG93, GLV99, Gol91, Luc96, MO97, ML95, Nak90, RS95, RW98, TLP95, Vis96, Fry95, MO05]. Applications [AGG97, CW96, DM91, GFG99, LTW99, LWK94, Pou99, Sch97]. Approach [BK96, BH90, BT95, BFV99, Buo96, EGS96, Fon90, Gam92, GG98, GS95, GL94, Hir91, Hsi96, LTCR90, MWC97, NJ93, Pau90, Sam97, Swe95a, WC99b, Won92, Woo92, GR98]. Approaches [AS98, VS93]. Approximate [Bel90, Dav86, Dav90b, DF91, DMY92, DS93, HL95, HSH96, KB90, Lav95, PP99, PR92, Raf96, Sev94, Swe95a, TKK89, TKK91, DDM94]. Approximating [WL96]. Approximation [BH94b, CW90, DY91, Lie94a, Lie94b, Liu98, Luc96, Rab93, Sev98a, Von96, Wol93, Lie96, Mij90a, Mij90b, MR93b]. Approximations [BB90, BHB92b, CFR94, DFF90, DM91, DMY93, HN95, Jen92, JFR94, Kuo99, OSS91, Ska93, Spa91, Swe95b, Wan92, BHB92a]. Arbitrarily [Luc94a]. ARIMA [Oda93]. Arising [LRW93]. ARMA [GZW97, PKK90]. Armed [LL92]. Arrays [BT96b, Che98, Sit93, Wu91]. Ascertainment [HT96]. Asking [Kuk90]. Asmussen [MM90, MM99]. Assays [Dav90a]. Assess [NJ93]. Assessing [CJ99, HP96b, Sla99, WC92a]. Assessment [AG97, HSH96]. Associate [GD97]. Associated [HHLR94]. Association [Bly94, LLH91, Shi98, ZHH^{+}98, Bly95]. Associations [BRL96]. Assumed [ZK95]. Assumption [Tsa90]. Asymptotic [AHT92, BHD98, Bro90, CBD94, Din96, FV98, GH90b, KS91, Kom96, Li92, RV91, Ste90a, Wal92]. Asymptotically [Ahm93, DC91, HKT90, Pig90, Ste93]. Asymptotics [BNL94, CNF95, MD93, PP94, Sch97, Swe92, PP95]. Attributable [DB94, Llo96b]. Augmentation [LWK94, MV99, WT90]. Author [Ano92a]. Autocorrelated [Azz94, Azz97, PK98]. Autocorrelation [Mon94]. Autocorrelations [Li92, NA93, VB91]. Automatic [Chi92, Paw96]. Autonomous [AM94]. Autoradiography [AA94]. Autoregression [CLT95, GZW97, Lie94b, Wan92, WL97a, Lie96]. Autoregression-Based [GZW97]. Autoregressions [BK95]. Autoregressive [Ahn93, BBO98, CT98, HN93, HT99, Kav91, KAW92, Luc94b, Ma97, Nie97, PF93, PS97a, Pos87, Rah99, HST90, Pos90]. Autoregressive-Integrated-Moving [Ahn93]. Autoregressive-Moving [Kav91, Luc94b, PS97a, Pos87, Pos90]. Auxiliary [CQ93, Kuk93, RKM90, RP95, ZP95]. Average [Ahn93, Fin86, GZW94, Had95, Hal89, Kav91, KAW92, Luc94b, Ma97, PS97a, Pos87, Fin90, Hal90a, Pos90].

B [Cox90, FS90, Kem90, Poc90]. Back [Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d]. Backcalculation [Din96]. Balance [LD91]. Balanced [GD97, ME98, MB93, RS99, Sit93, UM90, Udd90, Wu91, DK91]. Bands [HL99, LFW94]. Bandwidth [BHP98, Chi90, Chi91, Chi92, Hal90b, HSJM91, HGK92, RT97, WW90, WMP95]. Bandwidths [Chi90]. Bartlett [BHB92a, CS95, Cor98, DS93, Kak96, LK90, Mij90a, MC91, Nie97, Won91]. Bartlett-Type [Kak96, MC91]. Based [Ber90, Buc91, CP95, DDNN93, DO99, EH93, FL93, GZW97, Gou90, GT94, HKT90, HSJM91, HPT99, Hil90, Hir91, HKH92, HH93, HZ95, JN93, KW94, Kle93, LTCR90, Li98, LWY93, LN95, Nie99b, PWY94, QZ97, Rom90, Sat96, Sev90, Sit93, ST93, Swe92, TGF90, Vid95, WYL90, Wu91, YWL92, Zha99, ZK95, Zim91, GT95]. Bases [PW96]. Bayes [Efr93, FG93, LT92, MR92a, MB92a, O’N92, PY90, PR94b, Raf96, REKB98]. Bayesian [RLS94, AC95, BBW99, BGM96, BFV99, BM96, Cha91a, CS98, DG95, Dat96, DMS98, DW98, DF91, DM91, DS93, Dup95, EGS96, Fon90, FRW99, GK91, GR94, GS95, GR98, Gre95, Gri95, HS93, HL97, Kim91, Len91, Loh95, MY97, MS92a, MB92b, MEW96, MR97, OS93, PdBP95, PS95a, RLS93, RH99, Smi91, Swe95a, Swe95b, WJU97]. Bayesian-Frequentist [BBW99]. Begg [Cox90, FS90, Kem90, Poc90]. Behaviour [JF99, Wol90]. Behrens [Kim92]. Beliefs [GW98]. Bernoulli [FC91, FC90a, WNG94]. Better [Ahm98, EH90]. Between [CR92, EK96, Hon96, HP96b, Par97a, SD92, Sun94, dLA91, BNM93, Bin97, Far90, MR93b]. Beyond [Per99]. Bias [Alh90, BL95, BG96, CH98, DH92, DM95, Fir93, Fir95, GLV99, GGI92, HN93, Hal90b, HT91, JLN95, LK90, MRMRMP99, mN93a, NJ93, Neu99, NA93, PG91, PD97b, Ros92, Sil97, YD90]. Bias-Corrected [mN93a]. Biased [Beg90b, GK93, Jon91, RZ91, Wan96]. Bilinear [Gab98]. Binary [AC95, BG92, CZD93, CC95, CW92, CW94b, DB94, ESM95, FL93, LLH91, LD98, NL96, Neu99, RI93, TW97, Vla90, ZH91, ZP90]. Binning [HPT98]. Binomial [CH90, NM95, PW90, Tar79, Tar90, Wis68, Wis90]. Bioassay [GK91, RLS93, RLS94]. Bioequivalence [HSH96, HH93, HHLR94, WHD99, DDM94]. Biological [DBL97]. Biometrika [BT99, CP91, Dav90b, DM99, Fin90, Fir95, FC91, Gle96, GT95, Hal90a, JS93, Kos01, Kus99, Lie96, MM99, MKGL98, MO05, McC96, PP95, Pos90, PC93, RLS94, RB92, Roc90, Rud91, Sch94, ST91, Tho93, TKK91, UM99, Zel96, Ano90e, Ano92g]. Biplots [Gow90, Gow92]. Birth [OSS91]. Bivariate [CFG97, DBL97, LY93, Nie99a, OM92, Paw96, PW90, Shi98, TC98, Vis96, WW97, HP96a]. Blackwellisation [CR96]. Block [Ano99e, Bos96, BM96, DM96, GL92, GD97, ME92, UM90, Udd90, UM97, UM99]. Blocking [HHJ95]. Bonferroni [Ano04, Rom90, SC96, WJU97]. Bootstrap [BH94b, DY91, DHW92, DH92, DMY92, DH91, FR96, HHJ95, HPT99, HZ95, JFR94, YD90]. Bootstrapping [New96]. Boscovich [Far90]. Both [ALP98]. Boundaries [Pro99]. Boundary [WT90]. Bounded [PS95a]. Bounds [JS92, JS93, LRW93, Liu96c, RR93]. Box [HW93, Dua93]. Box-Counting [HW93]. Breaking [CG93]. Buckley [Cur96].

C. [Cox90, FS90, Kem90, Poc90]. Calculating [WT90]. Calculation [TW97]. Calculations [Efr93, JT91]. Calibration [BO91, FV98, Oma98]. Cancer [Kor93]. Canonical [GP95, GP97, Mor93, NF95, Nie99a, Rin96, Sch90b]. Capture [CJ91, Dup95, GR92, Llo92b, Llo92a, Smi91, WC92b, YHL96]. Capture-Recapture [CJ91, Dup95, GR92, Smi91, WC92b, YHL96]. Capturing [JF99]. CAR [STS99]. Carcinogenicity [MW92]. Carlo [BC91b, BH94b, BR99b, CK96, CS98, DF99, DK97, FMS97, Gam98, Gre95]. CART [DMS98]. Case [CL99b, HPZF99, LTCR90, LB95, MR97, QZC97, QZ97, Qin98, Rab97, Sam97, SJDT93, SW97, SZ99, TX93, WC93, WW93, WC92a, Whi95, Whi97, Wil91, XL99, Zha99, ZHH^{+}98]. Case-Cohort [CL99b]. Case-Control [CL99b, HPZF99, LTCR90, LB95, MR97, QZC97, QZ97, Qin98, Rab97, Sam97, SJDT93, SW97, WC93, WW93, Whi95, Whi97, Wil91, XL99, Zha99, ZHH^{+}98]. Cases [CJ99, Gho99]. Categorical [AL93, Glo96, LR98, MD95b, PdBP95, STM93, ST98, VS93]. Cauchy [McC92]. Causal [CW95, Daw95, FGS95, Fre95, HN93, IR95, Pea95a, Pea95b, Rob95, Ros95, Sha95, Sob95]. Cause [MW92, Lam98]. Cause-of-Death [MW92]. cause-specific [Lam98]. Causes [FRY98]. Censored [BML91, BSD96, CWY95, Cur96, DF91, ES90a, FYW98, GG94, HS90, KL98, LSY99, MZ92, MW90, Par97a, PR94a, PC92, RTA95, Sat96, SZ95, UW92, VH98, WYL90, Zho92, Zim91, Fry95, PC93]. Censoring [Cha91a, HG90, HL96, LY93, LRW96, PdBP95, RR95, TC98, WW97, WW98, YWH98]. Central [RT96]. Certain [BT96b]. Chain [CK96, DF99, Gam98, MB92a, Rud90, STC97, BR99b, FMS97, Gre95, Rud91]. Chains [BT98]. Change [BT96a, FC90b, GH90b, Hen90, Loa91, ME98, MW90, Rus91, Yao93, BT99]. Change-Over [ME98, Rus91]. Change-Point [Hen90, MW90]. Change-Points [Yao93]. Changed [FC90a, WNG94, FC91]. Changepoint [JF99, LTC97]. Changes [MW90]. Characterisation [EK96]. Characteristic [LTW99, Pep97, VB96]. Characterization [AT90]. Characterizing [RT92b]. Checking [FGH91, Gli99, LWY93, SL96]. Chi [CdPF91, GL97, HKH92, Zha99]. Chi-Squared [CdPF91, GL97, HKH92, Zha99]. Choice [BFV99, GE90, GG98, HGK92, Sun94, GR98]. Choose [Chi90]. Circular [KN94, Riv97]. Class [Ber94, DM91, GH90a, GD97, Glo96, GFG99, Jon97, Lit94, NM95, Pig90, Ste91, DK91, Lam98]. Classes [Sil97]. Classic [CBD94, FBCN97]. Classification [CM98, O’N92, Vla90]. Clinical [Beg90b, GD96, HZ97, Kor93, KF93, LKT96, Lin91, LL92, Pat91, RI93, SK97, TWF96, TW97, Wan91, WSL90]. Closed [Had95, MY97]. Cluster [MS97, RJ90, Tho91b]. Clustered [QZC97, BP99]. Clustering [BRL96, BC98]. Coefficient [HRWY98, KT99, KS91, Nag91, Rah99]. Coefficients [MM97b]. Coherence [Ait90]. Coherent [DDNN93]. Cohort [CL99b, XL99]. Coin [Beg90b]. Collapsibility [MM90, MM99]. Collection [BGM96, VH98]. Collinearity [SD92]. Column [JS92, LD91, SB91, JS93]. Columns [UM90, Udd90]. Combinations [ABS99]. Combined [FR99, ZHH^{+}98]. Combining [CM99, KC92, MS92a, MW97]. Comments [Kum93]. Common [mN93a, NF95, Pig90, Sch91b, Sch99, Dou97]. Commonality [EH93]. comparative [DDM94]. Comparing [CH90, GL92, Hir91, LRW96, MMLH92, PG93, VB96, ZL90]. Comparison [CES92, GGI92, HS90, HKH92, LTW99, UW92, VS93, Wak90, XL99]. Comparisons [LWK94, SWT90, Ste90b]. Competing [Dew92, FK96, FRY98, GR90, GR95, ZK95, Lam98]. Complete [GK94]. Completely [CL99a, BAM95]. Complex [RS99]. complexity [FMS97]. Complications [FR99]. Component [BL95, HKH92, KR92, Lin97a, MS90b, Nay91, Sch91b, Sch99, SG97, SC92]. Components [AAL90, Sch97, Shi97, ST99]. Compositions [ABS99]. Compound [Wal92]. Computation [Buc94, DGS94, GG94, Liu98, Gre95]. Computational [FMS97]. Computing [Bea91]. Condition [Jef90]. Conditional [BH90, Bel90, BK95, BB99, DF91, DM91, DMY93, FYT96, FY98, FGS91, Lev90, LT92, LI96b, Liu98, McC92, MV99, MRMRMP99, Mor91, MC91, Pro97, Sev90, Sev92, Swe95a, WL96, WL97b, WL97a]. Conditionality [LM98]. Conditionally [CK96]. Conditioned [EL97]. Conditioning [PP99, Sev94]. Conditions [Kos99, WW97, Kos01]. Conference [EN95]. Confidence [BK96, Che96, CL98, CW96, DMY92, DE92, Efr93, HHLR94, HL99, KSZ91, LY99, LFW94, Liu96c, Mon97, RR93, Rin96, Ros92, Sch90b, TW97]. Confounders [DM95]. Confounding [PW96, Rob92]. Conjugate [BFV99, GP97]. Conservative [Ano04, SC96]. Consistency [GG94, Hu97, SO96, Wes91]. Consistent [JN93, ZT97]. Constancy [CHK95]. Constrained [CS98, HM93, SG97, Wan91, WP92]. Constraints [RLS93, Sil92, RLS94]. Constructing [ME92]. Construction [BMM95, JR90, Liu96c, Rus91, Wil90, Won92, Wu93]. Constructions [UM90, Udd90]. Contact [HN90]. contacts [Far90]. Content [GP95, WT90]. Contingency [MM90, dLvdH91, MM99]. Continual [OR98, SO96]. Continuity [Duf90]. Continuity-Corrected [Duf90]. Continuous [CT98, FOS94, Kom93, LR98, Llo96b, Vla90, YHL96]. Continuum [Ste90b]. Contoured [GR94]. Contrast [Mit93, Ree91, Dua93]. Contrast-Type [Mit93]. Contrasts [Con92]. Control [CL99b, HPZF99, LTCR90, LB95, MR97, QZC97, QZ97, Qin98, Rab97, Sam97, SJDT93, SW97, WC93, WW93, Whi95, Whi97, Wil91, XL99, Zha99, ZHH^{+}98]. Controlled [RN90]. Convergence [HT95b, LR94, RT96, BR99b, KTDMB97]. Convergent [JJL91]. Convex [ABS99]. Convexity [Mit93]. Convolution [HT97]. Copula [ZK95]. Copulas [CFG97]. Corpuscle [VH90]. Corrected [CW96, Duf90, HT91, Nak90, mN93a, PG91]. Correcting [Buo96]. Correction [BL95, LK90, Lev90, Nie97]. Corrections [Ano04, Azz97, Bly95, CS95, CP91, Dav90b, Fir95, FC91, Gle96, GT95, JS93, Kos01, Lie96, MKGL98, MO05, McC96, Pos90, PC93, RLS94, Rud91, Sch94, Tar90, Tho93, TKK91, Whi97, Wis90, Zel96, BT99, DM99, Fin90, Hal90a, Kus99, MM99, PP95, RB92, Roc90, UM99]. Correlated [Ano99e, CP95, GD96, HGK92, HPZF99, Kun91, LLH91, LD98, RJ90, SL96, UM97, ZP90, UM99]. Correlation [ALP98, Bel90, Cro95, JA90, KS91, KN94, Kle93, Nie99a, Pau90, Sch97, VR90, Wat90, Wol90]. Correlations [Smy94]. Cost [KF93]. Count [JLCSS99]. Counter [LB95]. Counter-Matching [LB95]. Counting [HW93, LWY98, Wel94]. Counts [Ber90]. Covariables [FR91]. Covariance [Hon96, Hu97, KBB90, KN94, LWK94, Pou99, PC92, PC93, SG97, ZPB91]. Covariances [NL98]. Covariate [KS97, RP95, WD96a, ZP95]. Covariate-Adjusted [KS97]. Covariates [HL96, Lin92, LI96b, NJ93, Orm98, Pai96, PT97, RT92a, VS93, VH98]. Coverage [MR99]. Cox [BNM93, Bin92, CL99b, DGS94, Dua93, LWY93, MM97b, MC91, PT97, SL96]. Cramér [ÖH97]. Criteria [CBD94, DO99, FBCN97, HH93, HT95b, KK96]. Criterion [Anr99, BN95b, HT91, Pau98, QGG96, Sun94, VS95, Mij90b]. Cross [AM94, BCN94, DH92, Kun91, Paw96, PH97, Smy94, ST98, VS95]. Cross-Correlations [Smy94]. Cross-Over [Kun91]. Cross-Sectional [AM94]. Cross-Spectral [ST98]. Cross-Spectrum [Paw96]. Cross-Validation [DH92, VS95]. Cross-Validatory [BCN94]. Crosses [CM99, DM96, GD97, GK94, Muk97]. Crossover [BT95, FG93, Gri95]. Crossvalidatory [HT98]. Cumulants [Li98]. Cumulative [DDNN93, HKH92, LWY93, Ska93, SB96]. Cure [RR93]. Current [LOY98]. Curvature [Gua95, SC93]. Curve [CHT99, ES90b, FR91, Hal90b, HMT95, HPT98, MEW96, RLS93, RLS94]. Curved [BN95a]. Curves [FGH91, Hsi96, LTW99, LFW94, Pep97, VB96, ZL90]. Cusp [Wan95]. Cusum [Buc91, EH93].

D [Roc90]. Data [AM94, Azz94, Azz97, BRL96, BML91, BM93, Bin92, BSD96, Buc94, BCN94, CP95, CZD93, Car97, CM91, CL99a, CWY95, CH90, CH99, Cur96, Dav90a, Dav92, DF91, DM92b, Dup95, ES90a, FJ98, FYW98, GG94, Gle92, GR95, Gol91, Gra94a, Gri95, HSJM91, HHJ95, Hei94, HRWY98, HPZF99, HL96, JA90, Jon91, JC92, JLCSS99, Kas90, KR92, KF93, Lan92, LOY98, LSY99, LLH91, Lit94, LR98, LTC97, MW94, MS99, MD95a, MR92b, MV99, MKL97, Mor91, MW90, ML95, NZ99, PD99, Par97a, Pat91, Pau90, PdBP95, PS97a, Pep92, PR94a, Pou99, PC92, PS95b, QZC97, QZ97, RTA95, Rab97, RI93, Rah99, RKM90, RS92, RS95, RS99, RP95, RJ90, Sat96, SW97, STH94, SZ95, SL96, SD99b, TLH98, UW92, VH98, VB91, Wan96, Wat90, WT90]. Data [WYL90, WW93, Whi95, Whi97, Wil91, WC92b, Yin90, ZH91, Zha99, ZHH^{+}98, ZP95, BP99, Dou97, Fah94, Fry95, Gle96, HP96a, PC93]. Data-Based [HSJM91]. Data-Translated [Kas90]. Databases [REKB98]. Death [MW92, Fry95]. Debugging [CMY93, Gou90]. Decentred [Riv97]. Decision [BFV99]. Deck [RS92]. Decomposable [GG99, RW98]. Decomposition [AS98, AGG97, Wes97]. Deconvolution [SB96]. Defectives [Gri94, Wri92]. Defined [Jor93, Tsa94]. Definition [Nag91]. Degrees [BR99a]. Delay [HT96]. Delayed [VH98]. Deletion [PQ96]. Densities [FYT96, TKK89, WL92, TKK91]. Density [Ait90, BHHJ98, Che96, Chi91, Chi92, CH99, DG93, HSJM91, Jon91, JLN95, Len91, Qin98, SES90, SW91, Vid95, WG95, WT90, Wes91]. Dependence [DM90, GH90a, GGR95, HPZF99, Jon96]. Dependency [HP96b]. Dependent [BCN94, HHJ95, LRW96, ME92, ME98, MM97b, PD99, RT97, Rob92, RT92a, RR95, WW98, YWH98, ZK95]. Derivative [CHT99, MKGL96, MKGL98]. Derived [Cox93, Mor91]. Deriving [EJ92]. Design [Beg90b, BG92, Cha91b, CH90, DM92a, DC91, DW98, EGS96, HPT98, LLH90, Loh95, MMB97, ZH91, dLA91]. Design-Unbiased [DC91]. Designing [TW90]. Designs [AR96, Ano99e, BMM95, BT95, BS94, BM96, CR93, CM99, CS91, DW96, DW98, DM96, EN95, GL92, GD97, Hu98, JS92, Kun91, Kus97, LD91, ME92, ME98, MB93, O’B92, PK98, PD97b, PR92, RN90, Rus91, SWT90, SB91, UM90, Udd90, UM97, WJ96, Won92, Wu93, WZ93, BAM95, DK91, DH94, JS93, Kus99, UM99]. details [Far90]. Detecting [Buc91, CM97, Lef91, Pan93, Ros92, WNG94]. Detection [HZ97, PG93, Wan95, Zel93, Zel96]. Determination [AT90, DF99, GG99, Nag91, Pos87, Gre95, Pos90]. Determining [HH93]. Deterministic [Ahn93]. Deviance [Li93, QGG96]. Deviations [Mor92, New96]. diabetes [Fry95]. Diagnosis [FADD94]. Diagnostic [HS93, MRMRMP99, Pep97]. Diagnostics [CJP92, GT94, HW91, KT99, PQ96, WG91, GT95]. Diagrams [CW95, Daw95, FGS95, Fre95, IR95, Pea95a, Pea95b, Rob95, Ros95, Sha95, Sob95]. Diallel [CM99, DM96, GD97, GK94, Muk97]. Difference [BK96, BNL94, CR92, HKT90, HN95, Pro97]. Difference-Based [HKT90]. Differenced [Oda93]. Differences [Cha91b, Gra94a]. Differencing [Li91a, Luc96]. Different [MS92a]. Differential [GL97, SO98]. Diffusion [PR94b]. Dimension [HW93, OK91, WC99b]. Dimension-Reduced [WC99b]. Dimensional [Gow90]. Dimensions [FLR90]. Directed [BNC94, BN95a, SBSP99]. Direction [AM94, CH93]. Discrepancy [Cor98]. Discrete [FJ98, Fah94, Gol91, Liu96a, PR94b]. Discrete-State [Liu96a]. Discretization [Chi91]. Discretized [Buc94]. Discriminant [CV91b, DHR99, HFC90, O’N92, RV91, Wak90]. Discriminating [dLA91]. Discrimination [CF91, DH92, EK96, Vla90]. Discussion [Beg90a, Cox90, CW95, Daw95, FS90, FGS95, Fre95, IR95, Kem90, Poc90, Rob95, Ros95, Sha95, Sob95]. Discussions [Pea95b]. Disease [Gra94b, Mal95, PS95b, SZ99, TLP95, VJP97, Zel93, Zel96]. Diseases [GP95, HN90]. Dispersion [BL95, FBCN97, GG97, MM91, New96]. Distance [BHD98, Zim91, ÖH97]. Distance-Based [Zim91]. Distributed [Fin86, Fin90, KTDMB97]. Distribution [Ahm98, AD96, BHD98, Ber90, BSD96, BHP98, Bro90, CDH92, CdPF91, CW94b, DMY93, EH90, GP97, HT96, HL98, Kor93, Kuk93, KS96, KW97, MN93b, NM95, NHK92, OM92, PF93, RKM90, Shi95, Ska93, ST93, SB96, STS99, Tar79, Tar90, VJP97, VH98, VB96, WD96b, ZT97, BP99, CH93, PS96]. Distribution-Free [VB96]. Distributional [JF99]. Distributions [Ahm93, BH94a, BB90, CBD94, DT99, EK96, GGR95, GH90b, Kom96, Kuo99, LRW93, LRW96, LSY99, MO97, MM91, MR92a, MMLH92, PW90, Riv94, RT92b, Sla99, STS99, Taw90, Vid98, WS99, MO05]. Disturbance [Koo93]. Divergence [BHHJ98, Bly94, Bly95]. Double [BO91, DMY92]. Drift [GS91]. Drop [MKL97]. Drop-Out [MKL97]. Dual [GKR98, GL92]. Due [NJ93, Neu99, DM95]. Duration [WW98]. Durbin [Rab93]. Dynamic [Büh99, Fah94, Gam92, Gam98, HW91, Pan93]. Dynamical [FYT96].

Early [HZ97, Zel93, Zel96]. ECM [LR94, MR93a]. ECME [KTDMB97, LR94]. Edge [JSAHM98]. Editorial [Daw92, Daw93]. Editorship [Ano90e, Ano92g]. Edwards [MM90, MM99]. Effect [BML91, Che91, Chi91, Dav92, LTCR90, MB93, PY90, RZ91, RT92b]. Effective [CQ93, WMP95]. Effectiveness [KF93, ZL90]. Effects [AL93, Bea91, Car97, DG93, FH91, JVDW94, Lin97a, Loh95, MMB97, NHK92, NL96, New96, Oma91, OJ93, Sch91a, SPP96, Von96, WD96a]. Efficiency [JS92, Kus97, Llo94, NL96, Neu99, RV91, Slu91, SD99b, DK91, JS93, Kus99]. efficiency-balanced [DK91]. Efficient [Ano99e, BHHJ98, Bos96, DGS94, FY98, GSC95, HT95b, Lin97b, MV99, Nie99b, PG91, Pig90, PR92, QF91, Qui99, Rab97, SD99a, UM97, VJP97, VH98, UM99]. Efficiently [GL97]. Eigenprojections [Sch97]. Electron [AA94, JVDW94]. Elimination [Lis93, Sev94, Til96]. Ellipsoidally [LR98]. Elliptically [GR94]. Empirical [Bag98, CQ93, Che96, Cor98, CW95, Daw95, FGS95, Fre95, GP95, HFC90, Hsi96, IR95, LM98, LM99, LT92, MB92a, Mon97, Pea95a, Pea95b, Rob95, Ros95, Sev99, Sha95, ST93, Sob95, MR93b]. Empty [SB91]. Endpoints [Mül91]. Enhancements [CHT99]. Enhancing [HT97]. Entropy [CDL94, WG95]. Envelope [STM93]. Epidemic [Din96, Yao93]. Equality [MM91, Lam98]. Equation [AM94, BNL94, Buo96, Jef90, Lee90, Li98, Pau90]. Equations [CP95, CL99a, Cox93, GL97, HM98, HZ95, Lie98, LLH91, PQ96, SO98, Zha94]. Equineighboured [Udd90]. Equivalence [BK96, Bel90, HW97]. Error [BHO99, Buo96, Car97, CMR99, Che91, Chi91, DH92, FGS91, Jef90, O’N92, WD96a, WG91]. Errors [Ano99e, GR94, Hal89, KAW92, Kun91, Li92, ME92, MR97, Nak90, RT97, Sch93, SP96, UM97, ZH91, Hal90a, KTDMB97, UM99].

Errors-in-Variables [Nak90, SP96]. Estimate [OSS91, SW94]. Estimated [Con90]. Estimates [Bea91, Bro90, CC95, CW90, CV91b, DH92, Fir93, HRWY98, OK91, SES90, SW91, TX93, ZK95, Fir95]. Estimating [AM94, BM93, Buo96, CP95, Cha92b, CL99a, Cox93, GL97, God91, HL95, HM98, HZ95, HT96, Kav91, Kor93, Kuk93, Li98, Lie98, LN95, LLH91, Llo96b, MZ92, NL98, Nay91, PWY94, Pau90, PD97b, PQ96, QT91, Rab93, RKM90, RZ91, TLP95, WW90, Zha94]. Estimation [Ait90, AA94, BC91a, BHHJ98, Bax94, BC98, BG96, CFG97, CMF98, Cha91b, CMY93, CQ93, Che96, Chi91, Chi92, CH99, Dav90a, DP98, Dou97, Dro98, Dup95, DK97, EF90, Fah94, FYT96, FY98, FADD94, FC90b, GZW97, GE90, GGR95, GR92, GLV99, GK93, GW96, HKT90, HM90, HSJM91, HMT95, HPT98, HN90, HM93, HFC90, HG90, HPZF99, HL96, HL98, HT98, Jef90, Jon91, JLN95, KN94, Kle93, LSY99, LH96, Lin97b, LH99, Llo92b, Luc97, MY97, MS90b, MR93a, MS90c, NL96, Nie99b, Pau90, PG91, Paw96, Per95, PS97b, PC92, QF91, Qui99, Rab97, RS92, RS95, Ree91, Rob92, RT92a, RR95, Sch91a, SGW93, Sev92, SG97, She98, SZ99, SO98, Ste93, Sun94, TKT91, TWF96, VJP97, VH90, VH98, VR90, Vis96, WC93, WW98, Wel94, Wes91, WC92b, Woo92]. Estimation [Zho92, Zim91, BR99b, Fry95, PC93]. Estimative [Vid98]. Estimator [BC91a, BML91, BSD96, Car97, CT98, Cha92b, GZW94, GGI92, JF99, Len91, Lie94b, LY93, MD95a, mN93a, WMP95, WD96b, ZT97, Lie96]. Estimators [Ber94, CDH92, CHT99, CM98, Cur96, DC91, Din96, HN93, Hal90b, HW93, HS90, Hor90, Lie98, LWK94, MB92a, Mül91, PP92, PK91, Pig90, SB96, Ste91, Sun94, SD99b, TC98, UW92, WW97, WP92, CH93, HST90, MR93b, ÖH97]. Evaluate [HZ97]. Evaluation [DCCL94, LM98]. Event [FK98]. Events [LSY99]. Evidence [DM98, Gri94, Wri92, DM99]. Exact [Bha96, BB99, JT91, KS92, Luc94b, Ma97, PS97a, PP99]. Examinations [Par97b, Zel93, Zel96]. Example [FLR90]. Exceedances [STC97]. Excess [Sas96]. Exchangeability [Sug93]. Exchangeable [GW98]. Existence [NM95]. Expansion [WL92, LRW98]. Expansions [AHT92, Wan93, Wat90]. Experiment [VB96]. Experimental [DM92a, ZH91, dLA91]. Experiments [CJ91, Llo92b, Llo96a, Loh95, MS92a, MW97, PD97a, WP92, YHL96]. Explained [Sch90a, Sch92, Sch94]. Explanatory [Wel94]. Explicit [WP92]. Exploratory [Sun91]. Exponential [BN95a, BH94a, CFR94, CP89, CW94b, DE92, ES90a, GD90, GP97, MO97, McC94, McC96, SY96, TX93, YB91, ZP90, CP91, MO05]. Exposure [DB94, TW90]. Extended [XTL99]. Extending [Gri95]. Extension [LR94, MM90, BP99, DDM94, MM99]. Extensions [LR98]. External [MD95b]. Extra [MN93b, Mor91]. Extra-Multinomial [Mor91]. Extra-Poisson [Mor91]. Extrabinomial [Rud90, Rud91]. Extrapolation [DW96, SB96]. Extreme [CFG97, LT96, OM92, Shi95, Taw90]. Extremes [BT98].

Factor [JS92, Llo96b, MD95b, MB93, Sta97, JS93]. Factorial [CS91].

Factors [DB94, MR92a, PY90, PR94b, PS95b, Raf96, Rob92, WZ93]. Failure [BRL96, CP95, CMY93, Dew92, FK96, Gli99, GR90, GR95, HP96b, HPZF99, Jon97, KF93, LRW96, LWY98, Orm98, PC92, PH97, Rob92, RT92a, SL96, Tsa90, ZP95, HP96a, PC93]. Failure-Time [BRL96]. Failures [FRY98]. Familial [Gle92, KR92, Pau90, Gle96]. Families [BH94a, BD95, DE92, GD90, GGR95, MO97, McC94, MO05, McC96]. Family [CP89, GP97, HPZF99, MO97, SY96, WS99, Whi95, Whi97, ZHH^{+}98, CP91, MO05]. Fast [Buc94, DCCL94, DMY92, Luc94b, Luc96, QF91, Qui99]. Faster [LR94]. Feasibility [SW91]. Fewer [Rus91]. Field [GH90a, WI98]. Fields [AM94, Die91, FMS97]. Filtering [GR94, WC99b]. Finding [LLH90, SZ99]. Finite [CDH92, CSB92, CQ93, CDL94, DC91, Kuk93, KS96, KW97, MB92a, MN93b, MS90c, PK98, SB96, Wan93, WD96b, Wat90, BP99, FMS97]. Finite-Sample [KS96, KW97]. First [AT90, Cha91b, Had95, Li98, Lie94b, Wan92, Lie96]. First-Order [Lie94b, Wan92, Lie96]. Fisher [CH93, Kim92, O’N92, OM92, Pro97, Shi95]. Fit [Ahm93, Bag98, BD95, EH93, KL98, MK91, QZ97, Ree93, RB91, Shi98, Tar79, YB97, Zha99, RB92, Tar90]. Fits [Mor92]. Fitting [Ber93, Bin92, ES90b, MM97b, MEW96, NHK92, SW97, Wil91]. Fixed [DW98]. Following [EF90, Liu96c, LH99, TWF96]. Forecasting [SD99a]. Forecasts [Liu96b]. Forensic [DM98, REKB98, DM99]. forerunner [PS96]. Form [Had95, KS92]. Forms [Kuo99, Lie94a, Par97a]. Formula [FRW99, Vid95]. Formulae [ME92]. Four [GGI92, WZ93]. Four-Level [WZ93]. Fractal [HW93, OK91]. Fractional [Luc96, NA93]. Fractionally [Oda93]. Frailty [DBL97, FK96, Gli99]. Framework [Hil90, JN93, MR93a, Pep97, Swe95b]. Free [JR90, VB96]. Freedom [BR99a]. Frequency [JSAHM98, QT91, QF91, Qui99, WC92b]. Frequentist [BBW99, DG95, Dat96, FRW99, MD93, SY96]. Front [Ano90f, Ano90g, Ano90h, Ano90i, Ano91e, Ano91f, Ano91g, Ano91h, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano93e, Ano93f, Ano93g, Ano93h, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g, Ano95h, Ano96e, Ano96f, Ano96g, Ano96h, Ano97e, Ano97f, Ano97g, Ano97h, Ano98e, Ano98f, Ano98g, Ano98h, Ano99f, Ano99g, Ano99h, Ano99i]. Full [XL99]. Function [BSD96, CDH92, Had95, HT98, JJL91, Jon96, Kor93, LK90, Li93, LTW99, LY93, Llo96b, Ma97, MS92b, Nak90, O’N92, PC92, Pro99, QT91, Sev98a, ST93, SB96, TC98, Vis96, WD96b, WW97, Wat90, PC93]. functional [Dou97]. Functionals [BN97]. Functions [BHP98, CB97, Cox93, CB94, Dat96, Dav90a, DDNN93, DM91, DMY93, FY98, God91, HL95, Jor93, Kuk93, MG94, NL98, OJ93, Par97b, PWY94, RKM90, Sev98b, Sil97, Ska93, TKK89, WG95, Wak90, WL92, Bin97, TKK91]. Further [Gri94, KW97, Sch92, Far90].

G [PS97b]. Gamma [BHO99, Gli99, WI98]. Gap [LSY99]. Gauss [LP94]. Gaussian [Fin90, Ber94, CK96, Die91, DK97, Fin86, GG99, KS96, PD97b, RW98, She94, SP97, Sta97]. General [BML91, Bea91, BN95b, FRW99, GL94, GFG99, HL95, Hei94, HM98, Hil90, Kos99, LD91, LR98, Luc96, MR93a, Nag91, PdBP95, Sha90, WW93, Kos01]. Generalisation [Qui99, HP96a]. Generalised [BL95, BS94, CL99a, CBD94, CNF95, Cro95, FBCN97, Gab98, Gam98, KK96, Lin97a, ÖH97, PW96, PQ96, QGG96, Raf96, SD99b, VS95, GR98]. Generalized [Ber93, BB90, BHB92b, Gow92, LLH91, MG94, MS90b, Mor92, Nak90, NJ93, RJ90, Sch91a]. Generating [LD98]. Geometric [NJ93, RT96]. Gibbs [CK94, GR92, LWK94, Liu96a]. Gill [Cha92a]. Gittins [Wan91]. Globally [JJL91]. Good [Rus91]. Goodness [Ahm93, Bag98, BD95, EH93, KL98, MK91, QZ97, Ree93, RB91, Shi98, Tar79, Tar90, Zha99, RB92]. Goodness-of-Fit [Ahm93, Bag98, EH93, KL98, QZ97, Ree93, Shi98, Zha99]. Graphical [DF99, DP98, GG99, RW98, Sta97, WC92a]. Gröbner [PW96]. Group [BB92, CGS95, CL98, EJ92, EF90, LH99, PD97b, SPP96]. Grouped [FG93, Par97a, STH94]. Groups [Sch91b]. Growth [FR91].

Haenszel [HP96a, LP93]. Haenszel-type [LP93]. Hannan [Pos90, Pos87]. Hard [HT98]. Harmonic [Bro90]. Hastings [RT96]. Having [CdPF91]. Hazard [CP95, FK98, GP95, Hen90, Loa91, MW90, MWC97, Nie99b, OJ93, PH97, UW92, Lam98]. Hazards [AF95, Bin92, DS95, EK96, GT94, KC92, LFW94, LOY98, PT97, Sas96, Sat96, Sch90a, Sch92, Wan96, GT95, Sch94]. Heavily [BML91]. Hermite [LP94]. Heterogeneity [Dav92, LP93, Orm98, REKB98]. Heterogeneous [KBB90, Llo92b]. Heteroscedastic [CB94, DW98, WL97b]. Heteroscedasticity [CHT98, WL97a]. Heuristic [Efr97]. Hidden [Che98]. Hierarchical [DF99, HL97, REKB98, Sha90, STS99]. High [CHT99, FLR90, Nie99b, WC99a]. High-Derivative [CHT99]. High-Order [WC99a]. High-Quality [Nie99b]. Higher [Kak96, MD93, PP94, PP95]. Higher-Order [PP94, PP95]. Highest [WT90]. History [AG97, FK98, Bin97, Far90, PS96, Sti92]. HIV [BM93, Ber90, FADD94, TLP95]. Holgate [Bin97]. Homogeneity [LLH90, Llo92a, VV96, Wis68, Wis90]. Hot [RS92]. Hotelling [Mij90b]. Human [Zha94]. Hyperbolic [Smy94]. Hypotheses [BBW99, Con90, MR92a, Sil92, ZH91]. Hypothesis [DS95, Efr97, EF90, RJ90].

Ideal [DJ94]. Identically [Fin86, Fin90]. Identifiability [GL94]. Identification [AM94, AT90, DM98, GZW97, PKK90, RH99, Sha90, Sta97, DM99]. Ignorability [Hei94]. II [ES90a]. illness [Fry95]. illness-death [Fry95]. Image [Buc94, NR99]. Imaging [GS95]. Immunes [MZ92]. Imperfect [DM98, DM99]. Importance [BB99, DH91, GW96]. Improved [Ano04, CP89, CP91, EJ92, HS93, HST90, SC96]. Improving [PP99]. Imputation [BR99a, RS92, RS95, WR98]. Imputed [MR92b].

Inadmissibility [BC91a]. Incidence [BM93, DDNN93, MW92]. Including [GK91]. Incomplete [Bos96, GD97, Hei94, Lan92, Lin91, LI96b, Lit94, MD95a, PD99, Pat91, PS97a, PS97b, UM90, Udd90, YWH98]. Incomplete-Data [Hei94]. Incompletely [VS93]. Incorporating [Gra94b]. Increase [KF93]. Increment [PD97b]. Independence [Hon96, LT96, MK91, ST93, Tsa90, Sti92]. Independent [Fin86, PD97b, Fin90, Bin97]. Index [Ano92a, XTL99]. Indicator [SK97]. Indicators [DBL97]. Indices [Wan91]. Indirectly [Kuk90]. Individual [HW97]. Individuals [FADD94]. Inequality [Rom90, Sil92]. Infected [FADD94]. Infection [BM93, FADD94]. Infectious [GP95, HN90]. Inference [CG93, CJ91, DG95, Dat96, DFF90, DF91, DM91, DS93, FGS91, FRW99, HS93, LTW99, LT92, Loa91, McC92, NL98, NR99, OC94, OS93, Pep92, Sat96, Sev98b, Sug93, Sun94, Swe95a, YHL96, Zha92, FS99]. Inferences [Beg90b, KB90, LRW93, Qin98, YB91]. Influence [BD91b, CJ99, CV91b, Jor93, MRMRMP99, OJ93, PS95a, SD92, Shi97, SC93, WC92a, WG91]. Influential [Lef91]. Influenza [DM90]. Information [Ano90j, Ano90k, Ano91i, Ano91j, Ano92m, Ano92n, Ano93i, Ano94i, Ano94j, Ano95i, Ano96i, Ano97i, Ano97j, Ano98i, Ano99j, Anr99, AKW93, CQ93, Con92, ES90a, Gra94b, HL96, KK96, Kuk93, Lan92, Liu98, MW97, MKGL96, Nie99b, OM92, PK91, PS97b, RKM90, Shi95, SB98, dLA91, HST90, MKGL98]. Informative [PdBP95]. Informativeness [TLP95]. Initial [Bea91]. Insensitivity [Orm98]. Instrumental [Li95]. Integral [Wol90, SW94]. Integrated [Ahn93]. Integration [LN95]. Intensity [Par97b, Slu91, Zim91]. Intention [FR99]. Intention-to-Treat [FR99]. Interaction [DB94]. Interactions [Tho91a, Wu93, Tho93]. Intercept [WW93]. Interclass [KS91, KN94]. Intercorrelated [HT99]. Interim [HL99, KW94, WSL90]. Interpolation [AHT92, GS91, HT97]. Interpretation [Qui99]. Interval [BSD96, GG94, HFC90, RTA95, Sat96, TWF96, TW97, Fry95]. Interval-Censored [RTA95, Fry95]. Intervals [BK96, Che96, CL98, CW96, DMY92, DE92, Efr93, HPT99, HHLR94, LY99]. Intraclass [Pau90]. Intrinsic [BK95]. Invariant [BNC94, BN95a]. Inverse [AS98]. Invertibility [Oda93]. Inverting [WG95]. Investigation [HFC90]. Isotonic [DHR99]. Isserlis [RW98]. Iterated [BH94b]. Iteratively [Jor93, Tsa94].

J. [CW95, Daw95, FGS95, Fre95, IR95, Rob95, Ros95, Sha95, Sob95]. Jackknife [PP92, RS92, SW94]. James [Cur96]. Jeffreys [Kas90]. Joint [HL98, Pou99]. Jorgensen [Tsa94]. Jump [Wan95, Gre95]. June [Far90].

Kalman [GR94, WC99b]. Kaplan [SW94]. Karl [Sti92]. Kernel [Chi91, Chi92, CM98, GE90, Gra94b, HSJM91, HGK92, Jon91, Kuk93, LN95, MM97a, Mül91, RT97, SES90, VH90, Wel94, Wes91]. Known [GLL92, LH96]. Kriging [MKGL96, MKGL98]. Kullback [GR98, HST90]. Kurtosis [KBB90].

Lag [ZL90]. Lagrange [Li91a]. Landmark [DM92b, MW94]. Laplace [Lie94a, Von96, WG95, Wol93, WL92]. Large [Chr94, New96, PF93, WR98]. Large-Sample [WR98]. late [Bin97]. Latency [Ber90]. Lattice [WJ96]. Lawley [Mij90b]. Least [CT98, Chr94, HN90, Lie94b, Mor92, Pro97, QGG96, Lie96]. Least-Absolute-Deviations [Mor92]. Least-Squares [HN90]. Left [UW92]. Leibler [GR98, HST90]. Length [Efr97, Jon91, RB91, RB92, Wan96]. Length-Biased [Wan96]. Level [CS91, WZ93]. Levels [BRL96, FLR90, Sun91]. Leverage [SC93]. Life [Ahm92, AG97, Bax94, EK96, Kor93, Kum93, Lin92, OD90]. Lifetables [MWC97]. Lifetime [HL96, VJP97]. Likelihood [AA94, Bag98, BC91a, BN91, BN95a, BSD96, CG93, CFR94, CQ93, Che96, CS95, CP89, CW90, CR92, CB94, Dav86, DCCL94, DP98, Die91, DK97, Efr93, Fah94, Fir93, FL93, FC90b, GG94, GLV99, GH90b, Gou90, GL94, Had95, HPT99, HL95, Hen90, HM93, JF99, JJL91, Jen92, Kas90, LTCR90, Lav95, LM98, LM99, LY99, Lev90, Li93, Luc94b, Luc96, LTC97, Ma97, MD95a, MR92a, MS92b, MR92b, MR93a, Mon97, MC91, ML95, mN93a, Nie99a, NR99, OK91, OSS91, Pai96, Pau90, PS97a, PP94, SBSP99, Sch93, SP96, SW97, Sev90, Sev98a, Sev98b, Sev99, She98, SP97, STH94, SA93, Swe95b, TX93, WP92, Won91, BNM93, CP91, Dav90b, FH91, Fir95, MR93b, PP95]. Likelihood-Based [FL93, Gou90, Sev90]. Likelihood-Ratio [Nie99a]. Likelihoods [BNC94, Bel90, DHW92, Fra90, Luc94a, MB92b, Wat90]. Limit [GGI92, RT96]. Limiting [CT98]. Limits [Vid98]. Line [BC98]. Linear [ABS99, Ber93, BB90, BB99, BL95, BS94, Cha92a, Che99, CH98, CJP92, CBD94, CNF95, CV91b, Cro95, DF99, DF91, FG93, FBCN97, FYW98, FLR90, FS97, Gam98, GSC95, HW91, HS97, HZ95, HT95b, Lan92, Lie98, Lin97a, Luc96, MG94, MS90b, Mon94, Mor92, MS90c, Nak90, NJ93, O’N92, Par97a, Pau98, PP92, PG93, PR94a, QGG96, Raf96, Ree91, RJ90, RT92b, RV91, Sch91a, Sil92, SD99b, VS95, Wak90, WYL90, WG91, XTL99, Yin90, YWL92, Zha92, Zho92, DH94, Dou97, GR98]. Linearity [CW94a, HT95a, Mit93, SF98]. Link [MG94]. Linked [HG90]. Links [LH96]. Local [Bly94, Bly95, CH98, FBCN97, HMT95, Jon96, SD92, Sev93, Shi97, SC93]. Locally [CH90, VH98]. Locate [Gri94, Wri92]. Locating [FC90a, FC91]. Location [BHD98, BD95, FOS94, Liu96c, LR98, Pol91, Sev92]. Location-Scale [BD95, FOS94]. Log [BN91, BB99, DF99, GR90, HFC90, Lan92, Lin92, MMLH92, mN93a, Slu91, ZL90]. Log-Linear [BB99, DF99]. Logistic [Alh90, Azz94, Azz97, BH90, BD91b, CZD93, Chr94, CF91, Duf90, HL97, KC92, Lev90, NHK92, QZ97, RV91, SJDT93, TW90, VS93, WC93, Whi95, Whi97, Zha99]. Logrank [KS97]. Long [Ber93, Ber94, BT96a, BBO98, GH90a, Jun96, Luc96, PD99, RT97, BT99]. Long-Memory [Ber93, Ber94, BT96a, BBO98, BT99]. Long-Range [PD99, RT97]. Long-Term [Jun96]. Longitudinal [FL93, HRWY98, JA90, JLCSS99, LKT96, MS99, MKL97, ML95, Pou99, Rah99, SD99b, TLH98].

Look [DM92a]. Loss [ES90a, GG98, Neu99]. Lower [RR93].

M [Roc90, PS97b]. M/G/ [PS97b]. Magnetic [GS95]. Mantel [HP96a, LP93]. Many [LT92, SBSP99, VV96]. Marginal [Bel90, DFF90, DM91, ESM95, LN95, LLH90, MV99, STH94, SL96, TKK89, TLH98, ZK95, TKK91]. Marginalization [Wes91]. Mark [HL98]. Marker [Nie99b]. Markers [DBL97, JN93, Mal95]. Markov [Rud91, BT98, BR99b, CK96, CSB92, DF99, FMS97, Fry95, Gam98, Gre95, JC92, Li91b, MB92a, Rud90, STC97]. Martingale [LWY93, Llo92b, Llo94, TGF90]. Martingale-Based [LWY93, TGF90]. Masked [FRY98]. Mass [MM97a]. Matched [GKR98, LTCR90]. Matching [Gho99, LB95, MG97, MR99, RT92b, ST98]. Matrices [Con92, Hu97, Sch97]. Matrix [Cro95, Lan92, LLH90, Liu98, RW98, ZPB91]. Matter [Ano90a, Ano90b, Ano90c, Ano90d, Ano90f, Ano90g, Ano90h, Ano90i, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano99a, Ano99b, Ano99c, Ano99d, Ano99f, Ano99g, Ano99h, Ano99i]. Matusita [BHD98]. Maximal [HKH92]. Maximization [STH94]. Maximize [CDL94]. Maximizing [JJL91]. Maximum [AA94, BC91a, BSD96, CW90, DP98, DK97, Fir93, FC90b, GG94, GLV99, GH90b, JF99, MD95a, MR93a, mN93a, OK91, OSS91, Pau90, SW97, She98, TX93, WG95, WP92, Fir95]. Mean [Ahm92, BD91a, BS94, DY91, DC91, GH90b, Kum93, MS90c, Pol91, Pou99, RP95, Sch90b, Wil90, WS94]. Mean-Covariance [Pou99]. Means [BHO99, NL98, PK91, Riv94]. Measure [Bag98, EK96, LLH91, Ste91, Ste93]. Measurement [Buo96, CR93, Car97, CMR99, Che91, FG93, Sch93, SP97, WD96a, WG91, Kus99]. Measurements [Kus97, MD95b, Pat91, SP96, SPP96, TLH98, WSL90, YWH98]. Measures [Azz94, Azz97, Cro95, FYT96, Gua95, Hir91, TW90]. Measuring [PY90, REKB98]. Median [Chr94, KW94, CH93]. median-type [CH93]. Medical [Par97b, Pep97]. Meier [SW94]. Memory [Ber93, Ber94, BT96a, BBO98, BT99]. Method [Bea91, BM93, BCN94, DT98, DF91, EJ92, FL93, Gra94b, JLN95, Kuk93, LB95, Li93, LN95, MO97, OR98, PWY94, RL99, RP95, SO96, SO98, WG95, MO05]. Methodology [MS92a, Nak90]. Methods [CS98, CL98, FG93, HT97, HPT99, HFC90, Llo94, MY97, NR99, Pau98, PdBP95, PG93, RT92b, TBK95, WC99a, WL96, ZPB91, BAM95, FMS97, SM93]. Metropolis [RT96]. Microscope [AA94]. Minimal [Udd90]. Minimax [Dro98, Won92]. Minimaxity [Ste90a]. Minimising [BHHJ98]. Minimum [GL97, GG98, WZ93]. MINQUE [Kle93]. Misclassification [CH90, CV91b, ZH91]. Misclassified [CC95, Neu99]. Mises [CH93, ÖH97]. Missing [CL99a, Dew92, DP98, FR99, GR90, GR95, Gri95, Luc97, NZ99, Pai96, PT97, PK91, RS95, RP95, TLH98]. Misspecification [AF95, NHK92, SO96, XL99]. Mitochondrial [JVDW94]. Mixed [Bea91, BL95, CW92, DG93, GSC95, LR98, NM95, NHK92, NL96, Oma91, Pos87, Vla90, Von96, WD96a, Wol93, Wu91, Pos90]. Mixed-Effects [NHK92, NL96, Von96]. Mixture [GLL92, KC92, Lit94, MN93b, NHK92, BP99]. Mixtures [McC94, MEW96, McC96]. Modality [Die91]. Model [AL93, AF95, AGG97, AG97, BHD98, Bel90, BBO98, Ber90, BD91b, BGM96, Bro90, BN95b, BM96, CM91, CT98, CL99b, CV91a, Cro90, DG93, DF99, EK96, FK98, FOS94, FYW98, FGH91, FR91, Fuk99, GG98, GG99, Gli99, GR98, Had95, HW91, Hen90, Hil90, Hsi96, HZ95, HT95b, JC92, Jon97, JLCSS99, KT99, KK96, KC92, Lan92, LI96a, Li91b, Lin92, LWY93, LFW94, LY94, LLH90, LI96b, Liu96b, LR98, Llo92a, Loh95, Ma97, MS99, MM97b, MS94, MR97, NL96, Oma91, PT97, PS97a, Pol91, QGG96, Raf96, Rob92, RT92a, Rud90, Sat96, SO96, Shi98, Sil97, Sil92, Slu91, Smi91, Sta97, VS95, WL97b, XL99, YWL92, YB97, ZP90, Dua93, Gre95, HST90, KW97, Lam98, Rud91, SL96]. Model-Based [Hil90]. Modelling [BRL96, CZD93, Fah94, HT99, HG90, Kom93, Li92, MN93b, Pep97, Smy94, Taw90, BP99]. Models [AM94, Ahn93, AC95, BN97, BN95a, Bea91, Ber93, Ber94, Bin92, BB90, BB99, BT98, BL95, BC98, BS94, CK94, CK96, CMF98, CFR94, CWY95, CG98, CJP92, CP89, CBD94, CJ91, CW90, CW92, CNF95, CB94, Cro95, DBL97, DCCL94, De 97, DF99, DW98, DO99, DF91, DP98, DK97, EGS96, FK96, FBCN97, FR96, GZW94, GZW97, Gam98, GG97, GH90a, GSC95, GP95, GLV99, Glo96, GLL92, Gol91, Gua95, HN90, Hei94, HRWY98, Hu97, HT91, KT99, KL98, KS92, Koo93, Lee90, Li91b, Lin97a, LWY98, LH96, Lin97b, LI96b, Lit94, MG94, McC92, MS92b, MMB97, Mon94, Mon97, Mor92, MW90, Nak90, NM95, NHK92, NJ93, Nie97, NR99, O’B92, Orm98, OS93, Pai96, Pan93, Pau98, PP92, PG93, PP94, Pou99, PKK90, QZ97]. Models [Qin98, Raf96, Rah99, RL99, Ree91, RP95, REKB98, RJ90, RW98, SBSP99, Sch91a, SW97, Sha90, SG97, SZ99, STC97, SC92, ST99, Spa91, Ste90b, STS99, SD99b, TGF90, TLH98, TBK95, Von96, WD96a, WW93, WG91, Wil91, Wol93, WI98, XTL99, YB91, Zha92, Zha99, Zho92, dJS95, dLvdH91, dLA91, CP91, DH94, FS99, GR98, KTDMB97, LP93, PP95]. Modification [Ahm93, Mal95]. Modified [BN91, Buo96, CdPF91, CR92, Cro90, FS97, GR90, Jen92, KS96, LTCR90, Li95, Liu96a, Llo92b, Pos87, RS99, Rom90, SBSP99, Sev98a, Vid98, BNM93, Pos90]. Modulated [OC94]. Moment [HH93, Ste91, Ste93]. Moment-Based [HH93]. Moments [Bea91, GP97, Lie94a]. Monitoring [GD96, GFG99]. Monotone [Kak96, LR94, TLH98]. Monte [BC91b, BH94b, BR99b, CK96, CS98, DF99, DK97, FMS97, Gam98, Gre95]. Monti [KW97]. Morel [BP99]. MOSUM [CHK95]. Movement [Dup95]. Moving [Ahn93, Fin86, GZW94, Had95, Hal89, Kav91, KAW92, Luc94b, Ma97, PS97a, Pos87, Fin90, Hal90a, Pos90]. MR0836445 [Roc90]. MR0855892 [Dav90b]. MR0855914 [Fin90]. MR0919846 [Pos90]. MR0991422 [Hal90a]. MR0991426 [CP91]. MR1040637 [TKK91]. MR1041430 [ST91]. MR1603936 [MO05]. Multi [Sit93, TX93]. Multi-Arrays [Sit93]. Multi-Step [TX93]. Multicentre [RI93]. Multidimensional [RT96]. Multilevel [Gol91, Lee90]. Multimodality [SZ99]. Multinomial [BB92, CJ91, Din96, MN93b, Mor91, BP99]. Multiparameter [CFR94]. Multiperiod [Liu96b]. Multiple [BRL96, BR99a, CPV98, Dat96, FC90b, HM98, HKH92, LL92, Smi91, WR98, KTDMB97]. Multiple-Armed [LL92]. Multiplicative [FK98, FH91, Oma91, RZ91, Slu91, WW93]. Multiplicative-Intercept [WW93]. Multiplier [Li91a]. Multiply [MR92b]. Multiply-Imputed [MR92b]. Multiresponse [Gua95, KB90]. Multitaper [WMP95]. Multivariate [AAL90, AD96, BFV99, CZD93, CG98, DT99, ESM95, FG93, FOS94, FS99, FGS91, FS97, GGR95, Gli99, Glo96, JLCSS99, Kak96, KR92, LT96, Lin91, Ma97, MEW96, Orm98, PC92, PH97, SW91, Shi95, Sla99, Taw90, WS99, WHD99, WL97b, BHB92a, Kim92, Mij90a, MR93b, PC93].

Nagaraj [BP99]. Nanda [BHB92a, Mij90a]. Near [LT96, Mül91]. Neighbour [MB93, BAM95]. Nested [AR96, BBW99, LB95, Sam97, UM90, Udd90]. Networks [MS97]. Neumann [EH93]. Neyman [BC98, SM93]. no [BT99, CP91, Dav90b, DM99, Fin90, Fir95, FC91, Gle96, GT95, Hal90a, JS93, Kos01, Kus99, Lie96, MM99, MKGL98, MO05, McC96, PP95, Pos90, PC93, RLS94, RB92, Roc90, Rud91, Sch94, ST91, Tho93, TKK91, UM99, Zel96]. Nodes [SB91]. Noise [ALP98, NA93, So99, TKT91]. Non [BFV99, DK97, Fin86, Gho99, O’N92, RW98, She94, SP97, TLH98, Fin90]. Non-Bayes [O’N92]. Non-Conjugate [BFV99]. Non-Decomposable [RW98]. Non-Gaussian [DK97, Fin86, She94, SP97, Fin90]. Non-Monotone [TLH98]. Non-Regular [Gho99]. Noncircular [Wan92]. Noncompliance [FR99]. None [FR99]. Nonignorable [TLH98]. Noninformative [GG97, YB91]. Noniterative [GZW94, Pig90]. Nonlinear [AT90, BO91, Büh99, CLT95, CP89, DG93, FYT96, Gol91, Hu98, JVDW94, Li92, MR92a, O’B92, SC92, SC93, TKK89, Von96, WD96a, Wol93, WL92, CP91, DH94, TKK91]. Nonlinearity [HR99]. Nonnormal [BD91a, Dav90a, FLR90]. Nonnormality [Sla99]. Nonnull [CBD94]. Nonparameter [MW97]. Nonparametric [AKW93, BT95, BSD96, CFG97, CMR99, Cha91b, Che96, CHT99, CC95, DHR99, FK98, FGH91, Fry95, GK91, HKT90, HM90, HT97, HPT98, HK96, HT95a, HRWY98, HL98, JC92, KAW92, LY99, Len91, Lin91, LL92, LY93, LSY99, LN95, Lin97b, MD95a, MW90, NZ99, PG91, RLS93, RLS94, SES90, SGW93, Sil97, ST93, SM93, TC98, Vis96, WW97, WW98, Wel94]. Nonrandom [MKL97]. Nonregular [CB94, Cro90, Smi94]. Nonresponders [PR94a]. Nonresponse [Alh90]. Nonseasonal [ALP98].

Nonstable [SZ99]. Nonstationary [BBO98, DCCL94, Luc94b, Luc97]. Normal [Ahm93, AD96, Bel90, EGS96, GSC95, GLL92, Kuo99, Lit94, Liu98, MEW96, Pau98, RT92b, Mij90a, Mij90b]. Normality [Din96, MMS92]. Note [Böh94, BSD96, BNL94, BD95, Bro90, CM91, Cha92a, Cha92b, CR92, CW94b, CB94, Cur96, Dew92, DM95, GD90, Gle92, Gri94, GL94, HPT98, HR99, HSH96, Kav91, LD98, Nag91, NM95, O’B92, OD90, Par97a, PS95a, Rab97, SD92, SF98, Smy94, Tho91a, Tsa94, TC98, VS95, Vid98, Von96, WD96a, WJ96, WS94, Wol90, Wri92, Yat91, BNM93, Gle96, LP94, Tho93]. Nuisance [God91, LM99, LT92, Lis93, MD93, RL99, SBSP99, Sev93, Sev94, Sev98b]. Null [Ahm93]. Number [DW98, Nay91].

Object [RH99]. Observational [Ros92]. Observations [CV91b, KL98, Lef91, Lin91, Luc94a, ME98, PY90, PR94b, Fry95]. Observed [Fra90, Lan92, VS93]. Obtaining [Lan92]. Occasions [Fon90]. Odds [AL93, HFC90, LLH91, mN93a, Pig90, She98]. Offspring [Kle93, VR90]. Omitted [DM95, NJ93, Orm98]. One [EJ92, FK96, Jef90, Loh95, MB93, PP94, Sil92, WC92b, PP95]. One-Parameter [PP94, PP95]. One-Sided [EJ92, Sil92]. One-Way [Loh95]. Only [FK96]. Onset [VJP97]. Operating [LTW99, Pep97, VB96]. Optimal [BGM96, BS94, CR93, CM99, Cha91b, CH90, DW96, DW98, DM96, EJ92, FD91, GK93, GK94, HKT90, HSJM91, Li98, Loh95, MMB97, MS90c, Muk97, PK98, RN90, SWT90, Ste93, SB91, Wan99, Zel93, Zel96, DK91, DH94]. Optimality [BMM95, Buc91, DO99, Kus97, WJ96, Kus99]. Optimised [GW96]. Optimum [Mül91, dLA91]. Order [Anr99, AT90, Cha91b, CFR94, CdPF91, CNF95, Had95, HN95, Kak96, Kav91, LRW93, Lie94b, MD93, MG97, PF93, PP94, Pos87, Wan92, WC99a, Lie96, PP95, Pos90]. Ordered [AL93, BB92, GK91, LL92, SPP96]. Ordering [BGM96]. Orders [JR90]. Ordinal [MKL97]. Orthogonal [BT96b, Che98, Dro98, Sit93, Wil90, WS94, Wu91]. Orthogonality [God91, ST99]. Outbreaks [DM90]. Outcome [Pep92, TW90]. Outcomes [FR99]. Outlier [BH90, PG93, Roc86, Roc90]. Outliers [Pan93, Sha90]. Outlying [Luc94a]. Over-Influential [Lef91]. Overdispersed [GD90]. Overdispersion [Böh94].

Paired [Dav92, Gra94a, MW94, Ste90b, VB96]. Pairs [GKR98]. Pairwise [NR99]. Panels [HT99]. Paper [Cox90, FS90, Kem90, Poc90, Bin97]. Parameter [BT96a, Bro90, CM97, CMF98, CS98, CHK95, Cox93, EF90, GL94, GP97, LRW98, MO97, MD93, PP94, SD92, Sev93, Sev98b, SY96, Swe92, Wil90, BT99, MO05, PP95]. Parameter-Based [Swe92]. Parameter-Redundant [CMF98]. Parameterisation [Pou99]. Parameterization [HS93]. Parameters [Anr99, Bel90, CP95, DFF90, GGR95, God91, KBB90, Kim91, LM99, LT92, Lis93, Liu96c, RL99, RJ90, SBSP99, Sev94, WS94, WL92, DH94, Dua93]. Parametric [Ait90, CG93, CHT99, Dat96, LI96b, MS94, SZ99, TBK95, WR98]. Parametrisations [GSC95]. Parent [Kle93, VR90]. Parent-Offspring [Kle93, VR90]. Partial [Fon90, GD97, GL94, HMT95, Muk97, ML95, Sch99, She94, Sug93, SB98]. Partially [GD97, Luc94b, Luc97, Wu93, XTL99]. Partitioning [Büh99]. Partly [MS94]. Pattern [Lit94, ZPB91]. Pattern-Mixture [Lit94]. Patterns [OK91]. Pearl [CW95, Daw95, FGS95, Fre95, IR95, Rob95, Ros95, Sha95, Sob95]. Pearson [Sti92]. Pedigrees [Zha94]. Penalised [Wan99]. Penalized [Fah94]. Percentage [AAL90, Wal92]. Performance [HW93]. Performing [MR92b]. Period [BT95, CR93, Gri95]. Periodic [QT91, VB91]. Periodicities [Wal92]. Periodogram [Lee97]. Perspective [WJU97]. Peskun [Liu96a]. Phase [RS95, ST91]. Philip [Bin97]. Phylogenies [New96]. Pillai [BHB92a, Mij90a]. pitfalls [FS99]. Pivotal [PWY94]. Pivotals [Llo96a]. Planning [HZ97]. Plate [Buc94]. Plots [HM91, HS93]. Point [Gam92, Hen90, Loa91, MW90, NL98, Ste91, Ste93, TWF96, Zim91]. Points [AAL90, BO91, DW98, Flu90, FC90b, Wal92, Yao93]. Poisson [WI98, Böh94, CJ91, Lan92, Mor91, Sig91]. Poisson/Gamma [WI98]. Poland [Bin97]. Polyefficiency [Yat91]. Polygon [JSAHM98]. Polynomial [DW98, ES90b, PF93, Wat90]. Polynomial-Trigonometric [ES90b]. Polytomous [GK91]. Pooled [TLP95]. Pooling [MS92a]. Population [BC91a, CDH92, CDL94, DC91, Din96, Kuk93, MY97, MS90c, SB96, Wan93, WD96b]. Populations [BD91a, CSB92, CQ93, DS95, Hor90, PK98]. Portmanteau [KS96, KW97]. Possibly [Dav90a, Luc97]. Post [SK97]. Post-Randomisation [SK97]. Posterior [GG98, Hu97, MD93, NM95, Pol91, SY96, STS99, WT90, WC92a, WL92]. Poststratification [PK91]. Potency [RLS93, RLS94]. Power [BHHJ98, FBCN97, LM98, Li95, Pro97, WNG94, Dua93]. Powerful [Don91]. Practicable [Len91]. practical [Kim92]. Precision [Con92, Sun94]. Predicted [DO99]. Prediction [CJP92, DBL97, DCCL94, HPT99, JN93, Vid98, YWL92]. Predictions [TX93]. Predictive [BH94a, CJ99, Dav86, EGS96, GG98, HPT99, Kom96, LI96a, QGG96, Vid95, Vla90, Dav90b]. Predictor [Riv97]. preferable [BAM95]. Prelude [CH99]. Prequential [AG97]. Presence [CMR99, Cha91a, FR99, GS91, Hal89, LM99, LRW96, MD95b, MS90c, MD93, Rob92, RR95, Sev93, Sev98b, Wol90, dLA91, Hal90a]. Prevalence [TLP95]. Principal [Flu90, KR92, MS90b, Sch91b, Sch97, Sch99, Shi97]. Principle [PP94, PP95]. Prior [AKW93, Gri94, LI96a, PW90, RLS93, Wri92, dLA91, RLS94]. Priors [BB92, DG95, Dat96, GG97, Gho99, Lis93, Luc94a, MW97, MG97, MR99, SB98, YB91]. Probabilistic [PG93]. Probabilities [CH90, DGS94, DFF90, DM91, Dup95, Fra90, FRW99, LI96a, Llo92b, MR99].

Probability [CV91b, DT98, DT99, Gho99, MM91, MG97, MR99, SD99a, Til96, WG95, Wan92, Woo92, YWL92, Bin97, Far90, PS96, ST91, Sti92, Bin97]. Probable [BHO99]. Probit [CG98, Sch93]. Problem [BB92, EGS96, Hen90, PR94a, VH90, Kim92]. Problems [AS98, CG93, CS98, DH92, Hu98, Wan93, WC99a, SM93]. Procedure [Ano04, CFG97, GGR95, MS90a, Pro97, Rom90, SC96, Til96, VB96]. Procedures [Buc91, DHR99, HKH92, HL97, SZ99, WR98]. Process [CJ99, Hsi96, Nay91, Fry95]. Processes [BBO98, Che99, Gam92, HN93, Kom93, LWY98, Luc94b, Luc96, NL98, Nay91, OC94, Oda93, OSS91, PR94b, Ree93, Ste91, Ste93, Wel94, Zim91]. Product [CM98, GGI92, SM93]. Product-Limit [GGI92]. Profile [BN95a, CR92, LK90, SBSP99, Sev98a, SA93, BNM93]. Prognostic [DBL97]. Programmes [HZ97, SZ99]. Progression [Gra94b]. Projected [MS92b, ML95, WL96]. Projection [Che98, Sun91]. Projections [Sch90b, GR98]. Projective [BT96b]. Propensity [RT92b]. Proper [MB92b, DK91]. Properties [Bha96, BT96b, CDH92, CT98, Che98, Kom96, LM98, Pro99, Riv94, Sev90]. Property [MR99]. Proportion [MZ92]. Proportional [AL93, AF95, Bin92, BS94, DS95, EK96, GP95, GT94, GT95, KC92, LFW94, PT97, Sas96, Sat96, Sch90a, Sch92, She98, Sch94, ST91]. Proportionality [Cha92a]. Proportions [GLL92]. Proposal [Mon94]. Prospective [WW93, Wil91]. Providing [DG95, Dat96]. Pseudolikelihood [Sam97]. Publication [Sil97]. Pulses [Kom93]. Pure [OSS91]. Pursuit [Sun91]. PX [LRW98].

Quadratic [CW94b, Kuo99, Lie94a, McC94, O’B92, Wak90, ZP90, McC96]. quadrature [LP94]. Quality [Kor93, Nie99b, ZT97]. Quantile [GW96, Hor90, KNP94, Kos99, LTW99, BR99b, Kos01]. Quantiles [Lav95, MD93, RKM90, SY96]. Quantitative [CW92]. Quasi [Bax94, FH91, HM93, Li93, Pai96, QGG96, RL99, Sti92]. Quasi-Deviance [QGG96]. quasi-independence [Sti92]. Quasi-Likelihood [HM93, Li93, Pai96, FH91]. Quasi-Score [RL99]. Questions [Kuk90]. Queue [PS97b].

Random [CL99a, DG93, Die91, Fin86, FH91, GH90a, JVDW94, Lin97a, Loh95, MMB97, OK91, Oma98, OJ93, Rah99, Sch91a, WI98, Fin90, FMS97]. Random-Effects [MMB97]. Randomisation [SK97]. Randomised [BM96]. Randomization [BB90, Don91]. Randomized [MS90a]. Randomness [KS96]. Range [GH90a, Liu96c, PD99, RT97, Wat90]. Rank [AAL90, Cha92a, GR90, Lin92, MMLH92, Nie99a, Par97a, PR94a, Sat96, Slu91, Tho91a, WYL90, Yin90, ZL90, dLvdH91, Tho93]. Rank-Based [Sat96]. Rao [CR96, MK91]. Rare [Gri94, TLP95, Wri92]. Rate [DH92, FADD94, HN90, Hen90, Jun96, Loa91, MW92, OJ93, UW92, KTDMB97].

Rates [CMY93, HT95b, MW90, MWC97, O’N92, YHL96, Lam98]. Ratio [BN91, BHO99, CP95, CS95, CP89, Hen90, HFC90, LY99, Lie94a, LLH91, LTC97, MR92b, MC91, mN93a, Nie99a, Pig90, Qin98, Sev99, Ste90a, Won91, Woo92, CP91]. Ratios [HL95, HW97, PH97]. Reassessment [OR98, SO96]. Recapture [CMY93, CJ91, Dup95, GR92, Gou90, Llo92b, Llo92a, Llo94, Llo96a, Smi91, WC92b, YHL96]. Receiver [LTW99, Pep97, VB96]. Recentred [MM97a]. Record [CG93]. Recovery [CM91]. Recursive [ME92, NZ99]. Redistribution [Mal95]. Reduced [CW90, Ste91, Ste93, WC99b, dLvdH91]. Reducing [PD97b, RL99]. Reduction [CH98, Fir93, JLN95, Fir95]. Redundancy [CM97]. Redundant [CMF98]. Reference [BB92, Lis93, SB98]. Region [WT90]. Regions [KSZ91, Mon97, Rin96, Sch90b]. Regression [AC95, Alh90, AF95, AKW93, Azz94, Azz97, BN97, BD91a, BH90, Ber93, Bro90, BFV99, Buo96, BN95b, CHT98, CMR99, Cha91b, Cha92b, Chr94, CC95, CB97, CB94, DGS94, DO99, DF91, Dro98, Duf90, EGS96, ESM95, ES90b, FY98, FLR90, FS97, Gab98, GE90, Glo96, GL94, Gua95, HKT90, HM90, HT97, HK96, HS97, HS90, HM91, HGK92, HL97, Hu97, HT91, HT95b, JVDW94, Jun96, KB90, Kim91, KSZ91, KAW92, KC92, Li91b, LOY98, LN95, LH96, Lin97b, LI96b, MS99, MS90b, MM97b, MMB97, Mit93, Neu99, O’B92, Pai96, Pan93, Pep97, PK91, PH97, QZ97, RTA95, RT97, RP95, Riv97, RJ90, RR95, RV91, Sch93, SP96, Sch90a, Sch92, SW97, She98, Sig91, Smi94, Smy94, Spa91, SC93, SD99b, TW90, VS93, Wan99, Wan96, WYL90, Whi95]. Regression [Whi97, Wil91, YB91, YWL92, Zha92, Zha99, ZP90, ZP95, FS99, KTDMB97, Sch94]. Regressions [CZD93, Jon97]. Regret [Dro98]. Regular [Gho99]. Regularization [TKT91]. Reid [BNM93, MC91]. Reinsch [CES92]. Reject [Per99]. Rejective [Rom90]. Rejoinder [Pea95b]. Related [Pau98]. Relation [Par97a, BNM93]. Relationship [SD92, MR93b]. Relationships [Mor93, Dou97]. Relative [BM93, ES90a, GL94, HM91, HT95b, RV91, Slu91]. Remedy [Chi90]. Renewal [MD95a, OC94]. Renovated [SZ95]. Repeated [AL93, Azz94, Azz97, CR93, Cro95, DM90, FG93, Hir91, HL99, Kus97, LKT96, MD95b, Pat91, RS99, SPP96, Sit93, WSL90, Wu91, YWH98, Kus99]. Replacement [DT98, Fon90, MS97, Til96]. Replicate [SP96]. Replication [Dav90a, RS99]. Replications [Sit93, Wu91]. Representation [Pol91]. Representations [Fin86, Fin90]. Resampling [CL98, DH91, PWY94]. Research [CW95, Daw95, FGS95, Fre95, IR95, Pea95a, Pea95b, Rob95, Ros95, Sha95, Sob95]. Residual [Ahm92, AC95, Cha91a, EK96, HM91, Kum93, Li92, Mon94, OD90, SGW93]. Residuals [Duf90, GT94, HGK92, LWY93, TGF90, GT95]. Resistance [Roc86, Roc90]. Resolvable [BMM95]. Resonance [GS95]. Response [AC95, Beg90a, BG92, Coa91, CW94a, CW92, ESM95, GK91, Gol91, HS90, HG90, Llo96b, MS90a, SF98, SK97]. Responses [AL93, CC95, FL93, GD96, Glo96, NM95, Neu99, TW97]. Restoration [TKT91]. Restricted [Die91, LRW93]. Restriction [Anr99]. Results [KW97, MM90, MS92a, Sch92, MM99]. Retrospective [WC99a]. Reversibility [Che99]. Reversible [Gre95]. Revisited [FOS94, SGW93]. Right [Mal95, Par97a, UW92, VH98]. Right-Censored [VH98]. Ring [CM91]. Ring-Recovery [CM91]. Risk [BM93, DB94, GL94, HM91, KL98, LY94, Llo96b, MS94, PS95b, WW93, YB97]. Risks [Dew92, FK96, FRY98, GR90, GR95, ZK95, Lam98]. Rissanen [Pos90, Pos87]. Rival [dLA91]. Robust [BH94a, BHHJ98, BG92, BN95b, CS91, CF91, DW96, Gra94a, Hor90, Jef90, KS97, OS93, Sil92, WC93, ZPB91]. Robustness [DM92a, FOS94, O’N92, Sev92, TX93]. Rocke [Roc90]. Root [FR96, GS91, Hal89, Nie97, Hal90a]. Roots [Ahn93, Bha96, HM98, Li95, PF93]. Rounding [PR92]. Row [JS92, LD91, SB91, JS93]. Row-Column [JS92, LD91, SB91, JS93]. Rows [UM90, Udd90]. Rule [Gou90, OR98]. Rules [AHT92, CMY93, HHJ95, HMT95, JN93, Kas90, O’N92, Sta97]. Run [JR90].

Sacrifice [WP92]. Saddlepoint [BB90, BHB92a, BHB92b, DY91, HN95, Jen92, JFR94, Kuo99, LK90, Lev90, Lie94b, Lie96, Spa91, Wan93, MR93b]. Sample [BR99a, DDNN93, GFG99, KL98, Kos99, KS96, KW97, MZ92, NA93, PF93, Pep92, Qin98, Sig91, SK97, Sun94, WR98, Kos01, ST91]. Sampled [Kom93]. Sampler [De 97, LWK94, Liu96a]. Samples [Roc86, VV96, WC92b, HST90, Roc90]. Sampling [BB99, CK94, CR96, CSB92, CDL94, DT98, Fon90, GR92, GW96, Gri94, LB95, MS97, MRMRMP99, PK98, Per99, PD97a, QZC97, RN90, RS95, SJDT93, Sug93, Tho91b, Til96, WC99a, Wri92, ST91]. Scalar [DFF90]. Scale [BD95, FOS94, Gra94a]. Scaling [STM93]. Scan [De 97, WNG94]. Scatterplots [SZ95]. Scheduling [Zel93, Zel96]. Schemes [CR96, LWK94, MV99]. Schumacher [Cha92a]. Schwarz [Pau98]. Score [CHT98, CdPF91, CNF95, KS92, LK90, MK91, Nak90, Orm98, RL99, RP95, RT92b, WL96]. Scores [Con90, WL96]. Scott [BC98, SM93]. Screen [BG92]. Screening [EN95, SWT90, SZ99, TLP95]. Screens [BGM96]. Search [HK96]. Seasonal [ALP98, Li91a, VB91]. Seber [CM91]. Secant [Ska93, Smy94]. Second [CNF95, MG97, PF93, Ste91, Ste93]. Second-Order [MG97]. Sectional [AM94]. Seeking [MV99]. Segment [FC90a, WNG94, FC91]. Segregation [Zha94]. Selected [MS97]. Selection [AGG97, BBO98, BHP98, BN95b, Chi91, CPV98, DGS94, Dro98, FR91, Fuk99, GLV99, HSJM91, HT95b, KK96, LI96a, Liu96b, QGG96, RT97, VS95, Zha92, HST90, ST91]. Selector [Chi92, Lee97]. Selectors [Chi90]. Semilinear [KSZ91]. Semiparametric [BN97, BM96, FJ98, GGR95, GLV99, HL97, Jon97, LTW99, LY94, MS99, MS92b, MR97, OC94, Qin98, RL99, RT92a, RJ90, RR95, SK97, YB97]. Sensitive [Kuk90, Luc94a]. Sensitivity [Dua93, FYT96, GKR98, RL99]. Separating [DM90]. Sequence [FC90a, WNG94, FC91]. Sequences [Bos96, ST98]. Sequential [BC91a, BBW99, BC91b, CGS95, CL98, Coa91, CW96, EJ92, EF90, GD96, Hu98, JT91, Lin91, LL92, Lin92, Liu96c, LH99, PD97b, TWF96, TW97, TBK95, WSL90, Woo92]. Sequentially [Rom90]. Serial [ALP98, JA90, LSY99, ST93]. Series [AT90, Büh99, De 97, Dro98, Fin86, Fuk99, HR99, HT95a, HT99, Hon96, HT91, Kom93, KS92, KS96, Lef91, Li91b, Li92, Luc96, Luc97, Mon97, Paw96, SP97, So99, STM93, ST98, Udd90, VB91, WW90, Wes97, Wis68, Wol90, dJS95, Fin90, Wis90]. Sets [MR92b, Dou97]. Setting [NL96]. Settings [Ano99e, UM97, UM99]. Several [Liu96c, MW97, Sch91b, Won91]. Shape [DM92b, MW94, RLS93, RLS94]. Sharp [Wan95]. Sharpening [CH99]. Short [BBO98]. Short- [BBO98]. Shrinkage [CPV98, DJ94, Dro98]. Sibling [Kle93]. Sided [EJ92, Sil92]. Siegel [Wal92]. Sieve [She98]. Signal [Buc91, TKT91]. Signed [BN91, Jen92]. Significance [FLR90, Kim91, LKT96, Sev90, Sun91]. Significant [HK96, Pro97]. silly [BAM95, BAM95]. Simes [Ano04, SC96]. Simple [Anr99, FRW99, GZW94, GZW97, Gra94a, JLN95, Lee97, LY93, LR94, Vid95, Woo92]. Simplified [WW97]. Simpson [Far90]. Simulation [SO98, SB96, dJS95]. Simulations [EL97]. Simultaneous [BBW99, Efr97, GKR98]. Single [BL95, GL92, Sta97, XTL99, ZPB91]. Single-Factor [Sta97]. Single-Index [XTL99]. Size [BC91a, Din96, DM92b, MY97, Sig91, Udd90, WC92b, ST91]. Sizes [PK91]. Skew [AD96]. Skew-Normal [AD96]. Skewed [Hor90, Riv94]. Slopes [Oma98, Won91]. Small [BR99a, Roc86, VV96, HST90, Roc90]. Small-Sample [BR99a]. Smaller [Chi90]. Smooth [BD95, Buc91, DG93, DM91, Mül91]. Smoother [Koo93, dJS95]. Smoothers [MM97a]. Smoothing [BHP98, Buc94, CH98, EH93, FJ98, HMT95, HRWY98, KNP94, Lee97, TX93, Wan99]. Smoothing-Based [EH93]. Solution [HSH96, DDM94, Kim92]. Solutions [WP92]. Some [Ahm93, Ahn93, Bha96, Bos96, Che98, Cro90, GZW97, GR95, Kak96, Li91a, Li91b, OSS91, Sch91b, Sch97, SB91, UM90, Udd90]. Sources [DM90]. Space [CK94, CK96, CJ99, DCCL94, DK97, JLCSS99, Kom93, KS92, Koo93, She94, WC99b]. Space-Time [WC99b]. Spaced [JA90]. Span [Lee97]. Spatial [CJP92, Die91, DJ94, HN93, OK91, Wat90, WI98, Zim91]. Spatially [Ano99e, HT98, UM97, UM99]. Specific [AL93, BM93, CM99, Lam98]. Specification [GP95, LI96a]. Specifications [MS92a]. Specified [Ahm98, EH90]. Speckman [CES92]. Spectral [PG91, STM93, ST98, WMP95]. Spectrum [Paw96]. Spending [Pro99]. Spherical [OS93]. Spline [AKW93, Buc94, HS97, KAW92, Wan99]. Splines [CES92, KNP94, MKGL96, MKGL98]. Splitting [DT98]. Spread [HN90]. Square [AR96, BS94, WJ96]. Squared [CdPF91, GL97, HKH92, Zha99]. Squares [CT98, Chr94, HN90, Lie94b, Lie96]. Stable [BNC94, BN95a, SZ99]. Stage [SWT90]. Standard [Li92]. State [CK94, CK96, CJ99, DCCL94, DK97, JLCSS99, Kom93, KS92, Koo93, Liu96a, Ree93, She94]. State-Space [Kom93]. Stationarity [Kim91]. Stationary [BBO98, Fin86, Hon96, Luc94b, Fin90]. Statistic [BHB92b, CdPF91, HKH92, Jen92, MC91, Sev99, Wal92, WNG94, WC92a, Won91, BHB92a]. Statistical [AS98, Gam92, PD99, Riv94, SO98, WW90, WHD99]. Statistics [Ahm93, Cor98, CP89, Cox93, EL97, GFG99, HN95, Hil90, Jor93, Kak96, LT96, MMLH92, Par97a, Swe95b, Tsa94, WI98, Yin90, ZL90, Bin97, CP91, Far90, PS96, Sti92]. Status [HT96, LOY98]. Step [TX93]. Stochastic [Ber90, BNL94, EL97, FY98, GL97, Hu97, SO98]. Stopping [CMY93, Gou90, OR98]. Strata [Chr94, PK91]. Strategy [Pos87, Ste90a, Pos90]. Stratified [LB95, Tho91b]. Strength [Che98, HP96b]. Strong [Gri94, Hu97, Wri92]. Structural [Lee90]. Structure [KBB90, LWK94, PD97b]. Structured [LN95]. Student [FS99]. Studentized [DY91]. Studies [Bin97, Far90, FG93, HPZF99, LTCR90, Llo94, MR97, PS96, PS95b, Ros92, Sam97, SJDT93, Sti92, WC93, Whi95, ZHH^{+}98, Whi97]. Study [TW90]. Studying [SA93]. Subject [AL93, CH90, HT96, PD97a, VV96, ZH91]. Subject-Specific [AL93]. Subsampling [Fuk99]. Subsequent [FR99]. Subset [CPV98, DGS94]. Subspaces [Sch91b, Sch99]. Successive [Fon90, WW98]. Sufficient [EL97]. Suggested [AHT92]. Sums [LWY93]. Super [Nie99b]. Super-Efficient [Nie99b]. Superimposed [Nay91]. Supersaturated [Wu93]. Supplementary [HL96]. Surprising [Car97]. Surrogate [KF93, Pep92, TW90]. Surveillance [FD91]. Survey [Bin92, MRMRMP99, RKM90, RS92, RS99, Sug93, Sun94]. Surveys [BC98, ST91]. Survival [Ahm98, CW96, Dup95, EH90, Fah94, FRY98, GR95, Gra94b, GFG99, Hsi96, HT96, HL98, JC92, Jun96, KW94, LY93, LFW94, LTC97, MW90, OJ93, SWT90, Shi98, STH94, TGF90, TC98, TBK95, VH98, Vis96, WW97, WYL90, WP92, YWL92, ZT97, ZK95, ZL90, Fry95]. Survival/Sacrifice [WP92]. Survivor [PC92, PC93]. Symbolic [SA93]. Symmetric [LR98]. Symmetries [DT99]. Synergism [DB94]. Synthesis [JVDW94]. Synthetic [GS95]. System [AM94]. systematic [BAM95]. Systems [DDNN93, FYT96, SO98, SD99a].

T4 [Ber90]. Tables [Bax94, MM90, dLvdH91, MM99]. Tail [DFF90, DM91, Fra90, FRW99, Wan92]. Tale [FOS94]. Technique [QF91, Qui99]. Techniques [GZW97]. telescoping [SM93]. Tend [Chi90]. Term [Jun96]. Test [Ahm92, AAL90, Böh94, CL99a, CdPF91, Dew92, DT99, Don91, GR90, GS91, Hen90, HW97, Kak96, Kim91, Kum93, KW97, Liu96c, LH99, MW92, Mit93, Mon94, Nie97, Nie99a, Orm98, QZ97, Rom90, Shi98, ST93, Slu91, Wal92, Woo92, YB97, Zha99, HP96a, Mij90a, MMS92]. Testing [Ahm98, BK96, BT96a, BBW99, CGS95, CFR94, CW96, Con90, CB97, DS95, EH90, EF90, GLL92, Hal89, Hal90a, Hon96, LKT96, Li91b, Lin91, Lin97a, Llo92a, MM91, Mit93, Pep97, RJ90, SG97, SPP96, Tar79, Tar90, Tsa90, TLP95, VV96, VB91, Wis68, Won91, WL97a, Woo92, ZH91, ZPB91, BT99, Wis90]. Tests [Ahn93, ALP98, BH90, Bha96, BB99, BD95, CHT98, Cha92a, CLT95, CS95, CHK95, Coa91, CNF95, Cro90, DDNN93, EJ92, Efr97, EH93, FR96, GH90b, GT94, Gra94a, HR99, HT95a, HHLR94, JC92, KL98, KS97, Kos99, KS96, LTCR90, Li91a, Li95, LL92, Lin92, LTC97, MK91, MR92b, PP99, PR94a, Ree93, RB91, Sch91b, Sev90, Sil92, SK97, WHD99, WYL90, Yao93, GT95, JT91, Kos01, Lam98, RB92]. Their [Mor91]. Theorem [Liu96a]. Theorems [RT96]. Theory [BFV99, Kle93, WR98]. There [Rus91, ZL90]. Thin [Buc94]. Thin-Plate [Buc94]. Third [CFR94]. Three [BT95, Che98, CBD94, FBCN97, Gow90, Li98, VS93]. Three-Dimensional [Gow90]. Three-Period [BT95]. Three-Treatment [BT95]. Threshold [CT98, STC97, WL97a]. Thresholding [HT98]. Time [AT90, BRL96, Ber90, Büh99, CP95, Che99, De 97, Fin86, Fuk99, HR99, HT95a, HT99, Hon96, HRWY98, HP96b, HPZF99, HL98, Jon97, Kom93, KS92, KF93, KS96, Lef91, Li91b, Li92, LRW96, LWY98, LSY99, Llo96a, Luc97, MM97b, Mon97, Orm98, Paw96, PC92, PH97, RR93, Rob92, RT92a, SP97, So99, SL96, STM93, ST98, Tsa90, VB91, WW90, Wes97, WC99b, Wol90, YHL96, ZT97, ZP95, ZL90, dJS95, Fah94, Fin90, HP96a, HT91, PC93]. Time-Dependent [MM97b, Rob92, RT92a]. Time-Reversibility [Che99]. Time-Varying [HRWY98]. Times [Ahm92, Gli99, HL96, KW94, WW98, YWH98, Kum93]. Timing [Par97b]. Total [Mor91]. trace [BHB92a, Mij90a, Mij90b]. Transect [BC98]. Transfer [BN97]. Transform [Tho91a, Tho93]. Transformation [CWY95, FYW98, HPT98, HS97, Hsi96, MWC97, Dua93]. Transformations [ST99]. Transforming [CW94a, SF98]. Transforms [WG95]. Translated [Kas90]. Transmitted [Vis96]. Transport [JVDW94]. Treat [FR99]. Treating [Pan93]. Treatment [BML91, BT95, Dav92, FK96, FR99, FGS91, Kus97, RZ91, ZL90, Kus99]. Treatment-Noncompliance [FR99]. Treatments [CR93, GL92, GK93, Hir91, JR90, Kun91, Rus91]. Trend [CS91, JR90, MS90c]. Trend-Free [JR90]. Trends [Ahn93]. Trial [Pat91, RI93, TWF96, TW97, WSL90]. Trials [Beg90b, CL98, FG93, GD96, GLV99, HZ97, Kor93, KF93, LKT96, Lin91, LL92, SK97, Wan91, DDM94]. Trigonometric [ES90b]. Truncated [HL96, UW92, Woo92, Yin90]. Truncation [Tsa90]. Tukey [Yat91]. Tumour [MW92]. Two [BHO99, CR93, CS91, CH90, DM90, DB94, DDNN93, DS95, EK96, Fon90, GD97, Gri95, GFG99, Hon96, Hsi96, KL98, Kos99, Kos01, Kun91, Kus97, LRW96, MM91, MB93, MMLH92, Par97a, Qin98, RS95, Ree93, SWT90, SK97, SY96, WZ93, dLA91, DH94, Kus99, ST91]. Two-Factor [MB93]. Two-Level [CS91, WZ93]. Two-Parameter [SY96]. Two-Period [CR93, Gri95]. Two-Phase [RS95, ST91]. Two-Sample [DDNN93, GFG99, KL98, Kos99, Qin98, SK97, Kos01]. Two-Stage [SWT90]. Two-State [Ree93]. Two-Treatment [Kus97, Kus99]. Type [BN95b, Dew92, FK96, GR90, Kak96, Mit93, MC91, ZL90, CH93, LP93, ES90a]. Types [GR95].

Ultrastructural [CV91a, Mor93]. Unbiased [Cox93, DC91, Lie98, LH99]. Uncertain [MS92a]. Uncertainty [Raf96]. Unconditional [OJ93]. Unconstrained [Pou99]. Underfitted [HT91]. Understanding [BG96]. Unequal [CMY93, Dav90a, DT98, Til96]. Unequally [JA90]. Unified [BBO98, Won92]. Unifying [BK96]. Uniqueness [Fin86, Fin90]. Unit [Ahn93, Bha96, FR96, GS91, Hal89, Li95, Nie97, Hal90a]. Units [Fon90, Rus91]. Univariate [LY93, Sla99, TC98]. Universally [SB91]. Unknown [Ahm98, FBCN97, MG94]. Unreplicated [CV91a]. Unstable [Coa91]. Upper [Liu96c]. Usage [CQ93]. Use [AA94, Cro95, Von96]. Used [Ahm98, EH90]. Using [Alh90, AGG97, Cha91b, CC95, Cro95, HPZF99, HL99, KT99, KF93, Kuk93, LLH91, Mal95, Mit93, MEW96, PT97, Pep92, PK91, PC92, RKM90, REKB98, RT92b, Sla99, Sta97, STS99, TKT91, VV96, WC92b, ZP90, Zha94, PC93, TW90].

Vaccine [GLV99]. Validation [DH92, Pep92, SJDT93, VS95]. Validatory [BCN94]. Validity [DG95, Dat96, GL94, HW97, MD93, SY96]. Value [CFG97, OM92, Shi95, Taw90]. Values [CGS95, DP98, LT96, Luc97, BC91b]. Variability [DH92]. Variable [Coa91, CW94a, Hal90b, HS90, Jef90, LI96a, SF98, YHL96, Zha92]. Variables [BD91b, BFV99, CW92, Fin86, FC90a, HK96, HL98, Kuo99, Li95, LD98, Mor91, MR97, Nak90, SP96, Vla90, Wel94, Fin90, FC91]. Variance [AAL90, BG96, BS94, BHB92b, Cha91b, Cha92b, Dav90a, DO99, FY98, HKT90, HM90, KS91, Lin97a, ME98, Oma91, PP92, Per95, RS92, RS95, SGW93, SC92, ST99, Sun94, BAM95]. Variance-Balanced [ME98]. Variances [JVDW94, Ree91, VV96]. Variate [Rin96, Sch90b, MMS92]. Variates [HP96b, Mor93, NF95]. Variation [Llo96a, MN93b, Rud90, Sch90a, Sch92, Rud91, Sch94]. Varietal [Con92]. Varying [HRWY98, KT99]. Vector [BNL94, HR99, Luc94b, Luc96, Luc97]. Versions [GGI92]. Versus [Dro98, Mit93]. Vertically [Vis96]. Via [GR92, Hu97, CM98, GR98, Liu98, MR93a, MV99, Par97b, WT90]. Vital [HT96]. Volume [Ano90j, Ano90k, Ano91i, Ano91j, Ano92m, Ano92n, Ano93i, Ano94i, Ano94j, Ano95i, Ano96i, Ano97i, Ano97j, Ano98i, Ano99j].

Wars [Bin97]. Wavelet [AS98, AGG97, CHT98, DJ94, HT98, Per95]. Wavelets [CPV98, Wan95]. WaveShrink [BG96]. Way [Loh95]. Wei [Beg90b]. Weibull [MO05, Cro90, MO97]. Weight [Sil97]. Weighted [CDL94, Chr94, GT94, MMLH92, ZL90, GT95, ÖH97]. Weights [GE90, LRW93]. Welch [DDM94, HSH96]. Whether [Ahm98, EH90]. which [GH90a]. White [ALP98]. Wicksell [VH90]. Wilks [BHB92b]. Winsorized [Riv94]. Within [Slu91, VV96]. Within-Subject [VV96]. Without [DT98, MS97, Til96]. Working [Cro95].

XLII [Far90]. XLIII [Sti92]. XLIV [PS96]. XLV [Bin97].

Yield [Ahm93]. Youden [AR96].

References

Aykroyd:1994:UEA

[AA94] R. G. Aykroyd and C. W. Anderson. Use of the EM algorithm for maximum likelihood estimation in electron microscope autoradiography. Biometrika, 81(1):41–52, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337048.

Amemiya:1990:PPT

[AAL90] Yasuo Amemiya, T. W. Anderson, and Peter A. W. Lewis. Percentage points for a test of rank in multivariate components of variance. Biometrika, 77(3):637–641, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337003.

Aitchison:1999:CLC

[ABS99] J. Aitchison and J. Bacon-Shone. Convex linear combinations of compositions. Biometrika, 86(2):351–364, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673517.

Albert:1995:BRA

[AC95] Jim Albert and Siddhartha Chib. Bayesian residual analysis for binary response regression models. Biometrika, 82(4):747–759, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337342.

Azzalini:1996:MSN

[AD96] A. Azzalini and A. Dalla Valle. The multivariate skew-normal distribution. Biometrika, 83(4):715–726, December 1996. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337278.

Anderson:1995:MMP

[AF95] Garnet L. Anderson and Thomas R. Fleming. Model misspecification in proportional hazards regression. Biometrika, 82(3): 527–541, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337531.

Arjas:1997:PMA

[AG97] Elja Arjas and Dario Gasbarra. On prequential model assessment in life history analysis. Biometrika, 84(3):505–522, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337575.

Antoniadis:1997:MSU

[AGG97] Anestis Antoniadis, Irène Gijbels, and Gérard Grégoire. Model selection using wavelet decomposition and applications. Biometrika, 84(4):751–763, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337649.

Ahmad:1992:NTM

[Ahm92] Ibrahim A. Ahmad. A new test for mean residual life times. Biometrika, 79(2):416–419, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336855. See comments [Kum93].

Ahmad:1993:MSG

[Ahm93] Ibrahim A. Ahmad. Modification of some goodness-of-fit statistics to yield asymptotically normal null distributions. Biometrika, 80(2):466–472, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337217.

Ahmad:1998:TWS

[Ahm98] Ibrahim A. Ahmad. Testing whether a survival distribution is new better than used of an unknown specified age. Biometrika, 85(2):451–456, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337370.

Ahn:1993:STU

[Ahn93] Sung. K. Ahn. Some tests for unit roots in autoregressiveintegrated-moving average models with deterministic trends. Biometrika, 80(4):855–868, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336877.

Anderson:1992:IRS

[AHT92] Niall H. Anderson, Peter Hall, and D. M. Titterington. Interpolation rules suggested by asymptotic expansions. Biometrika, 79(3):651–653, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336796.

Aitchison:1990:CPD

[Ait90] J. Aitchison. On coherence in parametric density estimation. Biometrika, 77(4):905–908, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337114.

Ansley:1993:NSR

[AKW93] Craig F. Ansley, Robert Kohn, and Chi-Ming Wong. Nonparametric spline regression with prior information. Biometrika, 80 (1):75–88, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336758.

Agresti:1993:POM

[AL93] Alan Agresti and Joseph B. Lang. A proportional odds model with subject-specific effects for repeated ordered categorical responses. Biometrika, 80(3):527–534, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337174.

Alho:1990:ANB

[Alh90] Juha M. Alho. Adjusting for nonresponse bias using logistic regression. Biometrika, 77(3):617–624, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337000.

Andrews:1998:TWN

[ALP98] Donald W. K. Andrews, Xuemei Liu, and Werner Ploberger. Tests for white noise against alternatives with both seasonal and nonseasonal serial correlation. Biometrika, 85(3):727–740, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337399.

Abramson:1994:EDF

[AM94] Ian Abramson and Hans-Georg Müller. Estimating direction fields in autonomous equation models, with an application to system identification from cross-sectional data. Biometrika, 81 (4):663–672, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337069.

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. Biometrika, 77(1):??, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336074.

Anonymous:1990:BMb

[Ano90b] Anonymous. Back matter. Biometrika, 77(2):??, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336831.

Anonymous:1990:BMc

[Ano90c] Anonymous. Back matter. Biometrika, 77(3):??, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337013.

Anonymous:1990:BMd

[Ano90d] Anonymous. Back matter. Biometrika, 77(4):??, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337116.

Anonymous:1990:EB

[Ano90e] Anonymous. The editorship of Biometrika. Biometrika, 77(4): ??, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337090.

Anonymous:1990:FMa

[Ano90f] Anonymous. Front matter. Biometrika, 77(1):??, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336043.

Anonymous:1990:FMb

[Ano90g] Anonymous. Front matter. Biometrika, 77(2):??, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336800.

Anonymous:1990:FMc

[Ano90h] Anonymous. Front matter. Biometrika, 77(3):??, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336977.

Anonymous:1990:FMd

[Ano90i] Anonymous. Front matter. Biometrika, 77(4):??, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337089.

Anonymous:1990:VIa

[Ano90j] Anonymous. Volume information. Biometrika, 77(1):??, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336042.

Anonymous:1990:VIb

[Ano90k] Anonymous. Volume information. Biometrika, 77(4):i–viii, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337088.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. Biometrika, 78(1):??, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336919.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. Biometrika, 78(2):??, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337275.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. Biometrika, 78(3):??, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337042.

Anonymous:1991:BMd

[Ano91d] Anonymous. Back matter. Biometrika, 78(4):??, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336951.

Anonymous:1991:FMa

[Ano91e] Anonymous. Front matter. Biometrika, 78(1):??, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336889.

Anonymous:1991:FMb

[Ano91f] Anonymous. Front matter. Biometrika, 78(2):??, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337247.

Anonymous:1991:FMc

[Ano91g] Anonymous. Front matter. Biometrika, 78(3):??, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337014.

Anonymous:1991:FMd

[Ano91h] Anonymous. Front matter. Biometrika, 78(4):??, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336921.

Anonymous:1991:VIa

[Ano91i] Anonymous. Volume information. Biometrika, 78(1):??, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336888.

Anonymous:1991:VIb

[Ano91j] Anonymous. Volume information. Biometrika, 78(4):i–viii, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336920.

Anonymous:1992:AI

[Ano92a] Anonymous. Author index. Biometrika, 79(??):1–62, 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331431.

Anonymous:1992:BMa

[Ano92b] Anonymous. Back matter. Biometrika, 79(??):??, 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331432.

Anonymous:1992:BMb

[Ano92c] Anonymous. Back matter. Biometrika, 79(1):??, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337167.

Anonymous:1992:BMc

[Ano92d] Anonymous. Back matter. Biometrika, 79(2):??, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336861.

Anonymous:1992:BMd

[Ano92e] Anonymous. Back matter. Biometrika, 79(3):??, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336799.

Anonymous:1992:BMe

[Ano92f] Anonymous. Back matter. Biometrika, 79(4):ix–x, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337246.

Anonymous:1992:EB

[Ano92g] Anonymous. The editorship of Biometrika. Biometrika, 79(2): ??, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336833.

Anonymous:1992:FMa

[Ano92h] Anonymous. Front matter. Biometrika, 79(??):??, 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331430.

Anonymous:1992:FMb

[Ano92i] Anonymous. Front matter. Biometrika, 79(1):??, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337141.

Anonymous:1992:FMc

[Ano92j] Anonymous. Front matter. Biometrika, 79(2):??, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336832.

Anonymous:1992:FMd

[Ano92k] Anonymous. Front matter. Biometrika, 79(3):??, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336776.

Anonymous:1992:FMe

[Ano92l] Anonymous. Front matter. Biometrika, 79(4):??, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337221.

Anonymous:1992:VIa

[Ano92m] Anonymous. Volume information. Biometrika, 79(1):??, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337140.

Anonymous:1992:VIb

[Ano92n] Anonymous. Volume information. Biometrika, 79(4):i–viii, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337220.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. Biometrika, 80(1):??, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336775.

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. Biometrika, 80(2):??, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337219.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. Biometrika, 80(3):??, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337195.

Anonymous:1993:BMd

[Ano93d] Anonymous. Back matter. Biometrika, 80(4):??, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336887.

Anonymous:1993:FMa

[Ano93e] Anonymous. Front matter. Biometrika, 80(1):??, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336752.

Anonymous:1993:FMb

[Ano93f] Anonymous. Front matter. Biometrika, 80(2):??, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337196.

Anonymous:1993:FMc

[Ano93g] Anonymous. Front matter. Biometrika, 80(3):??, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337168.

Anonymous:1993:FMd

[Ano93h] Anonymous. Front matter. Biometrika, 80(4):??, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336863.

Anonymous:1993:VI

[Ano93i] Anonymous. Volume information. Biometrika, 80(4):i–viii, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336862.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. Biometrika, 81(1):??, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337064.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. Biometrika, 81(2):??, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336976.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. Biometrika, 81(3):??, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337139.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. Biometrika, 81(4):ix–x, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337087.

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. Biometrika, 81(1):vi, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337044.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. Biometrika, 81(2):??, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336952.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. Biometrika, 81(3):??, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337117.

Anonymous:1994:FMd

[Ano94h] Anonymous. Front matter. Biometrika, 81(4):??, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337066.

Anonymous:1994:VIa

[Ano94i] Anonymous. Volume information. Biometrika, 81(1):??, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337043.

Anonymous:1994:VIb

[Ano94j] Anonymous. Volume information. Biometrika, 81(4):i–viii, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337065.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. Biometrika, 82(1):??, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337646.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. Biometrika, 82(2):??, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337426.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. Biometrika, 82(3):??, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337548.

Anonymous:1995:BMd

[Ano95d] Anonymous. Back matter. Biometrika, 82(4):??, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337354.

Anonymous:1995:FMa

[Ano95e] Anonymous. Front matter. Biometrika, 82(1):??, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337622.

Anonymous:1995:FMb

[Ano95f] Anonymous. Front matter. Biometrika, 82(2):??, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337403.

Anonymous:1995:FMc

[Ano95g] Anonymous. Front matter. Biometrika, 82(3):??, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337525.

Anonymous:1995:FMd

[Ano95h] Anonymous. Front matter. Biometrika, 82(4):??, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337328.

Anonymous:1995:VI

[Ano95i] Anonymous. Volume information. Biometrika, 82(4):i–viii, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337327.

Anonymous:1996:BMa

[Ano96a] Anonymous. Back matter. Biometrika, 83(1):??, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337453.

Anonymous:1996:BMb

[Ano96b] Anonymous. Back matter. Biometrika, 83(2):??, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337621.

Anonymous:1996:BMc

[Ano96c] Anonymous. Back matter. Biometrika, 83(3):??, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337524.

Anonymous:1996:BMd

[Ano96d] Anonymous. Back matter. Biometrika, 83(4):??, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337303.

Anonymous:1996:FMa

[Ano96e] Anonymous. Front matter. Biometrika, 83(1):??, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337427.

Anonymous:1996:FMb

[Ano96f] Anonymous. Front matter. Biometrika, 83(2):??, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337597.

Anonymous:1996:FMc

[Ano96g] Anonymous. Front matter. Biometrika, 83(3):??, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337502.

Anonymous:1996:FMd

[Ano96h] Anonymous. Front matter. Biometrika, 83(4):??, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337277.

Anonymous:1996:VI

[Ano96i] Anonymous. Volume information. Biometrika, 83(4):i–viii, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337276.

Anonymous:1997:BMa

[Ano97a] Anonymous. Back matter. Biometrika, 84(1):??, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337573.

Anonymous:1997:BMb

[Ano97b] Anonymous. Back matter. Biometrika, 84(2):??, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337478.

Anonymous:1997:BMc

[Ano97c] Anonymous. Back matter. Biometrika, 84(3):??, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337596.

Anonymous:1997:BMd

[Ano97d] Anonymous. Back matter. Biometrika, 84(4):??, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337674.

Anonymous:1997:FMa

[Ano97e] Anonymous. Front matter. Biometrika, 84(1):??, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337550.

Anonymous:1997:FMb

[Ano97f] Anonymous. Front matter. Biometrika, 84(2):??, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337454.

Anonymous:1997:FMc

[Ano97g] Anonymous. Front matter. Biometrika, 84(3):??, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337574.

Anonymous:1997:FMd

[Ano97h] Anonymous. Front matter. Biometrika, 84(4):??, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337648.

Anonymous:1997:VIa

[Ano97i] Anonymous. Volume information. Biometrika, 84(1):??, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337549.

Anonymous:1997:VIb

[Ano97j] Anonymous. Volume information. Biometrika, 84(4):i–viii, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337647.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. Biometrika, 85(1):??, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337326.

Anonymous:1998:BMb

[Ano98b] Anonymous. Back matter. Biometrika, 85(2):??, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337380.

Anonymous:1998:BMc

[Ano98c] Anonymous. Back matter. Biometrika, 85(3):??, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337402.

Anonymous:1998:BMd

[Ano98d] Anonymous. Back matter. Biometrika, 85(4):??, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337501.

Anonymous:1998:FMa

[Ano98e] Anonymous. Front matter. Biometrika, 85(1):??, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337304.

Anonymous:1998:FMb

[Ano98f] Anonymous. Front matter. Biometrika, 85(2):??, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337355.

Anonymous:1998:FMc

[Ano98g] Anonymous. Front matter. Biometrika, 85(3):??, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337381.

Anonymous:1998:FMd

[Ano98h] Anonymous. Front matter. Biometrika, 85(4):??, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337480.

Anonymous:1998:VI

[Ano98i] Anonymous. Volume information. Biometrika, 85(4):i–viii, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337479.

Anonymous:1999:BMa

[Ano99a] Anonymous. Back matter. Biometrika, 86(1):??, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673555.

Anonymous:1999:BMb

[Ano99b] Anonymous. Back matter. Biometrika, 86(2):??, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673530.

Anonymous:1999:BMc

[Ano99c] Anonymous. Back matter. Biometrika, 86(3):??, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673670.

Anonymous:1999:BMd

[Ano99d] Anonymous. Back matter. Biometrika, 86(4):??, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673604.

Anonymous:1999:EBD

[Ano99e] Anonymous. Efficient block designs for settings with spatially correlated errors. Biometrika, 86(1):233, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673553.

Anonymous:1999:FMa

[Ano99f] Anonymous. Front matter. Biometrika, 86(1):??, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673531.

Anonymous:1999:FMb

[Ano99g] Anonymous. Front matter. Biometrika, 86(2):??, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673507.

Anonymous:1999:FMc

[Ano99h] Anonymous. Front matter. Biometrika, 86(3):??, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673648.

Anonymous:1999:FMd

[Ano99i] Anonymous. Front matter. Biometrika, 86(4):??, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673581.

Anonymous:1999:VI

[Ano99j] Anonymous. Volume information. Biometrika, 86(4):i–viii, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673580.

Anonymous:2004:AC

[Ano04] Anonymous. Amendments and corrections: “Is the Simes Improved Bonferroni Procedure Conservative?”. Biometrika, 91(1):249–250, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

biomet.oxfordjournals.org/content/91/1/249;   http:/
/www.jstor.org/stable/20441095. See [SC96].

Anraku:1999:ICP

[Anr99] Kazuo Anraku. An information criterion for parameters under a simple order restriction. Biometrika, 86(1):141–152, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673543.

Altan:1996:NYS

[AR96] S. Altan and D. Raghavarao. Nested Youden square designs. Biometrika, 83(1):242–245, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337448.

Abramovich:1998:WDA

[AS98] F. Abramovich and B. W. Silverman. Wavelet decomposition approaches to statistical inverse problems. Biometrika, 85(1): 115–129, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337313.

Auestad:1990:INT

[AT90] Bjørn Auestad and Dag Tjøstheim. Identification of nonlinear time series: First order characterization and order determination. Biometrika, 77(4):669–687, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337091.

Azzalini:1994:LRA

[Azz94] A. Azzalini. Logistic regression for autocorrelated data with application to repeated measures. Biometrika, 81(4):767–775, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337079. See correction [Azz97].

Azzalini:1997:ACL

[Azz97] A. Azzalini. Amendments and corrections: “Logistic Regression for Autocorrelated Data with Application to Repeated Measures”. Biometrika, 84(4):989, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337672. See [Azz94].

Baggerly:1998:ELG

[Bag98] Keith A. Baggerly. Empirical likelihood as a goodness-of-fit measure. Biometrika, 85(3):535–547, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337384.

Bailey:1995:NMP

[BAM95] R. A. Bailey, J.-M. Azaïs, and H. Monod. Are neighbour methods preferable to analysis of variance for completely systematic designs? “Silly designs are silly!”. Biometrika, 82(3):655–659, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337543.

Baxter:1994:EQL

[Bax94] Laurence A. Baxter. Estimation from quasi life tables. Biometrika, 81(3):567–577, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337127.

Booth:1990:RDS

[BB90] James G. Booth and Ronald W. Butler. Randomization distributions and saddlepoint approximations in generalized linear models. Biometrika, 77(4):787–796, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337101.

Berger:1992:OGR

[BB92] James O. Berger and Jose M. Bernardo. Ordered group reference priors with application to the multinomial problem. Biometrika, 79(1):25–37, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337144.

Booth:1999:ISA

[BB99] James G. Booth and Ronald W. Butler. An importance sampling algorithm for exact conditional tests in log-linear models. Biometrika, 86(2):321–332, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673514.

Beran:1998:UMS

[BBO98] J. Beran, R. J. Bhansali, and D. Ocker. On unified model selection for stationary and nonstationary short- and long-memory autoregressive processes. Biometrika, 85(4):921–934, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337493.

Berger:1999:SBF

[BBW99] James O. Berger, Benzion Boukai, and Yinping Wang. Simultaneous Bayesian-frequentist sequential testing of nested hypotheses. Biometrika, 86(1):79–92, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673538.

Bai:1991:IML

[BC91a] Z. D. Bai and Mo Suk Chow. Inadmissibility of the maximum likelihood estimator in the sequential estimation of the size of a population. Biometrika, 78(4):817–823, December 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336933.

Besag:1991:SMC

[BC91b] Julian Besag and Peter Clifford. Sequential Monte Carlo p-values. Biometrika, 78(2):301–304, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337256.

Brown:1998:CAE

[BC98] B. M. Brown and Ann Cowling. Clustering and abundance estimation for Neyman–Scott models and line transect surveys. Biometrika, 85(2):427–438, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337368.

Burman:1994:CVM

[BCN94] Prabir Burman, Edmond Chow, and Deborah Nolan. A crossvalidatory method for dependent data. Biometrika, 81(2):351– 358, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336965.

Beath:1991:RMN

[BD91a] Kenneth J. Beath and Annette J. Dobson. Regression to the mean for nonnormal populations. Biometrika, 78(2):431– 435, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337271.

Bhattacharjee:1991:IVL

[BD91b] S. K. Bhattacharjee and I. R. Dunsmore. The influence of variables in a logistic model. Biometrika, 78(4):851–856, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336937.

Boulerice:1995:NST

[BD95] Bernard Boulerice and Gilles R. Ducharme. A note on smooth tests of goodness of fit for location-scale families. Biometrika, 82(2):437–438, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337422.

Beal:1991:CIE

[Bea91] Stuart L. Beal. Computing initial estimates with mixed effects models: A general method of moments. Biometrika, 78(1):217– 220, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336912.

Begg:1990:RD

[Beg90a] C. B. Begg. Response to discussion. Biometrika, 77(3):484, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2336985. See [Beg90b].

Begg:1990:IWB

[Beg90b] Colin B. Begg. On inferences from Wei’s biased coin design for clinical trials. Biometrika, 77(3):467–478, September 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336980. See discussion

[FS90, Poc90, Kem90, Cox90] and reply [Beg90a].

Bellhouse:1990:EMA

[Bel90] D. R. Bellhouse. On the equivalence of marginal and approximate conditional likelihoods for correlation parameters under a normal model. Biometrika, 77(4):743–746, December 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337097.

Berman:1990:SMD

[Ber90] Simeon M. Berman. A stochastic model for the distribution of HIV latency time based on T4 counts. Biometrika, 77(4):

733–741, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337096.

Beran:1993:FLM

[Ber93] Jan Beran. Fitting long-memory models by generalized linear regression. Biometrika, 80(4):817–822, December 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336873.

Beran:1994:CEG

[Ber94] Jan Beran. On a class of M-estimators for Gaussian longmemory models. Biometrika, 81(4):755–766, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337078.

Brown:1999:CVM

[BFV99] Philip J. Brown, Tom Fearn, and Marina Vannucci. The choice of variables in multivariate regression: A non-conjugate Bayesian decision theory approach. Biometrika, 86(3):635–648, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673659.

Boys:1992:RDS

[BG92] R. J. Boys and K. D. Glazebrook. A robust design of a screen for a binary response. Biometrika, 79(3):643–650, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336795.

Bruce:1996:UWV

[BG96] Andrew G. Bruce and Hong-Ye Gao. Understanding WaveShrink: Variance and bias estimation. Biometrika, 83(4):727–745, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337279.

Boys:1996:BMO

[BGM96] R. J. Boys, K. D. Glazebrook, and C. M. McCrone. A Bayesian model for the optimal ordering of a collection of screens. Biometrika, 83(2):472–476, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337618.

Bedrick:1990:OTL

[BH90] Edward J. Bedrick and Joe R. Hill. Outlier tests for logistic regression: A conditional approach. Biometrika, 77(4):815–827, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337104.

Basu:1994:RPD

[BH94a] Ayanendranath Basu and Ian R. Harris. Robust predictive distributions for exponential families. Biometrika, 81(4):790–794, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337083.

Booth:1994:MCA

[BH94b] James G. Booth and Peter Hall. Monte Carlo approximation and the iterated bootstrap. Biometrika, 81(2):331–340, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336963.

Bhargava:1996:SPE

[Bha96] Alok Bhargava. Some properties of exact tests for unit roots. Biometrika, 83(4):944–949, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337300.

Butler:1992:SAB

[BHB92a] Ronald W. Butler, S. Huzurbazar, and James G. Booth. Saddlepoint approximations for the Bartlett–Nanda–Pillai trace statistic in multivariate analysis. Biometrika, 79(4):705–715, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337226.

Butler:1992:SAG

[BHB92b] Ronald W. Butler, S. Huzurbazar, and James G. Booth. Saddlepoint approximations for the generalized variance and Wilks’ statistic. Biometrika, 79(1):157–169, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337156.

Bar-Hen:1998:ADM

[BHD98] Avner Bar-Hen and Jean J. Daudin. Asymptotic distribution of Matusita’s distance: Application to the location model.

Biometrika, 85(2):477–481, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337374.

Basu:1998:REE

[BHHJ98] Ayanendranath Basu, Ian R. Harris, Nils L. Hjort, and M. C. Jones. Robust and efficient estimation by minimising a density power divergence. Biometrika, 85(3):549–559, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337385.

Booth:1999:PER

[BHO99] James G. Booth, James P. Hobert, and Pamela A. Ohman. On the probable error of the ratio of two gamma means. Biometrika, 86(2):439–452, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673524.

Bowman:1998:BSS

[BHP98] Adrian Bowman, Peter Hall, and Tania Prvan. Bandwidth selection for the smoothing of distribution functions. Biometrika, 85(4):799–808, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337484.

Binder:1992:FCP

[Bin92] David A. Binder. Fitting Cox’s proportional hazards models from survey data. Biometrika, 79(1):139–147, March 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337154.

Bingham:1997:SHP

[Bin97] N. H. Bingham. Studies in the history of probability and statistics. XLV. The late Philip Holgate’s paper: “Independent functions. Probability and analysis in Poland between the Wars”. Biometrika, 84(1):159–173, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337563. With an introduction by N.

H. Bingham.

Besag:1995:CIA

[BK95] Julian Besag and Charles Kooperberg. On conditional and intrinsic autoregressions. Biometrika, 82(4):733–746, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337341.

Bauer:1996:UAC

[BK96] Peter Bauer and Meinhard Kieser. A unifying approach for confidence intervals and testing of equivalence and difference. Biometrika, 83(4):934–937, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337298.

Breslow:1995:BCG

[BL95] Norman E. Breslow and Xihong Lin. Bias correction in generalised linear mixed models with a single component of dispersion. Biometrika, 82(1):81–91, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337629.

Blyth:1994:LDA

[Bly94] Stephen Blyth. Local divergence and association. Biometrika, 81(3):579–584, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2337128. See correction [Bly95].

Blyth:1995:ACL

[Bly95] S. Blyth. Amendments and corrections: “Local divergence and association”. Biometrika, 82(3):667, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337545. See [Bly94].

Becker:1993:MEA

[BM93] Niels G. Becker and Ian C. Marschner. A method for estimating the age-specific relative risk of HIV infection from AIDS incidence data. Biometrika, 80(1):165–178, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336766.

Bush:1996:SBM

[BM96] Christopher A. Bush and Steven N. MacEachern. A semiparametric Bayesian model for randomised block designs. Biometrika, 83(2):275–285, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337600.

Bassiakos:1991:GET

[BML91] Yiannis C. Bassiakos, Xiao-Li Meng, and Shaw-Hwa Lo. A general estimator of the treatment effect when the data are heavily censored. Biometrika, 78(4):741–748, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336925.

Bailey:1995:COA

[BMM95] R. A. Bailey, H. Monod, and J. P. Morgan. Construction and optimality of affine-resolvable designs. Biometrika, 82(1):187– 200, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337638.

Barndorff-Nielsen:1991:MSL

[BN91] O. E. Barndorff-Nielsen. Modified signed log likelihood ratio. Biometrika, 78(3):557–563, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337024.

Barndorff-Nielsen:1995:SIA

[BN95a] O. E. Barndorff-Nielsen. Stable and invariant adjusted profile likelihood and directed likelihood for curved exponential models. Biometrika, 82(3):489–499, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337528.

Burman:1995:GAT

[BN95b] P. Burman and D. Nolan. A general Akaike-type criterion for model selection in robust regression. Biometrika, 82(4):877–886, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337352.

Bagdonavicius:1997:TFS

[BN97] Vilijandas B. Bagdonavičius and Mikhail S. Nikulin. Transfer functionals and semiparametric regression models. Biometrika, 84(2):365–378, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337463.

Barndorff-Nielsen:1994:SIA

[BNC94] O. E. Barndorff-Nielsen and S. R. Chamberlin. Stable and invariant adjusted directed likelihoods. Biometrika, 81(3):485– 499, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337121.

Boswijk:1994:NAS

[BNL94] H. Peter Boswijk, Heinz Neudecker, and Shuang Zhe Liu. A note on the asymptotics of a stochastic vector difference equation. Biometrika, 81(1):216–218, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337063.

Barndorff-Nielsen:1993:NRB

[BNM93] Ole E. Barndorff-Nielsen and Peter McCullagh. A note on the relation between modified profile likelihood and the Cox–Reid adjusted profile likelihood. Biometrika, 80(2):321–328, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337202.

Brown:1991:DPN

[BO91] Philip J. Brown and Samuel D. Oman. Double points in nonlinear calibration. Biometrika, 78(1):33–43, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336893.

Bohning:1994:NTP

[Böh94] Dankmar Böhning. A note on a test for Poisson overdispersion. Biometrika, 81(2):418–419, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336974.

Bose:1996:SEI

[Bos96] Mausumi Bose. Some efficient incomplete block sequences. Biometrika, 83(4):956–961, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337302.

Banerjee:1999:EMN

[BP99] T. Banerjee and S. R. Paul. An extension of Morel–Nagaraj’s finite mixture distribution for modelling multinomial clustered data. Biometrika, 86(3):723–727, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673667.

Barnard:1999:SSD

[BR99a] John Barnard and Donald B. Rubin. Small-sample degrees of freedom with multiple imputation. Biometrika, 86(4):948–955, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673599.

Brooks:1999:QEM

[BR99b] S. P. Brooks and G. O. Roberts. On quantile estimation and Markov chain Monte Carlo convergence. Biometrika, 86(3): 710–717, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673665.

Bandeen-Roche:1996:MFT

[BRL96] Karen J. Bandeen-Roche and Kung-Yee Liang. Modelling failure-time associations in data with multiple levels of clustering. Biometrika, 83(1):29–39, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337430.

Brown:1990:NAD

[Bro90] Emery N. Brown. A note on the asymptotic distribution of the parameter estimates for the harmonic regression model. Biometrika, 77(3):653–656, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337006.

Burridge:1994:ODG

[BS94] Jim Burridge and Paola Sebastiani. D-optimal designs for generalised linear models with variance proportional to the square of the mean. Biometrika, 81(2):295–304, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336960.

Bohning:1996:ICD

[BSD96] Dankmar Bohning, Peter Schlattmann, and Ekkehart Dietz. Interval censored data: A note on the nonparametric maximum likelihood estimator of the distribution function. Biometrika, 83(2):462–466, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337616.

Bellavance:1995:NAA

[BT95] F. Bellavance and S. Tardif. A nonparametric approach to the analysis of three-treatment three-period crossover designs. Biometrika, 82(4):865–875, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337351.

Beran:1996:TCL

[BT96a] Jan Beran and Norma Terrin. Testing for a change of the long-memory parameter. Biometrika, 83(3):627–638, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337514.

See corrections [BT99].

Box:1996:PPC

[BT96b] George Box and John Tyssedal. Projective properties of certain orthogonal arrays. Biometrika, 83(4):950–955, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337301.

Bortot:1998:MEM

[BT98] Paola Bortot and Jonathan A. Tawn. Models for the extremes of Markov chains. Biometrika, 85(4):851–867, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337488.

Beran:1999:ACT

[BT99] Jan Beran and Norma Terrin. Amendments and corrections:“Testing for a change of the long-memory parameter” [Biometrika 83 (1996), no. 3, 627–638; (98a:62111)]. Biometrika, 86(1):233, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673552. See [BT96a].

Buckley:1991:DSS

[Buc91] M. J. Buckley. Detecting a smooth signal: Optimality of Cusum based procedures. Biometrika, 78(2):253–262, June 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337250.

Buckley:1994:FCD

[Buc94] M. J. Buckley. Fast computation of a discretized thin-plate smoothing spline for image data. Biometrika, 81(2):247–258, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336955.

Buhlmann:1999:DAP

[Büh99] Peter Bühlmann. Dynamic adaptive partitioning for nonlinear time series. Biometrika, 86(3):555–571, September 1999. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673654.

Buonaccorsi:1996:MEE

[Buo96] J. P. Buonaccorsi. A modified estimating equation approach to correcting for measurement error in regression. Biometrika, 83 (2):433–440, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337612.

Carroll:1997:SEM

[Car97] R. J. Carroll. Surprising effects of measurement error on an aggregate data estimator. Biometrika, 84(1):231–234, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337569.

Crisp:1994:NNL

[CB94] Adam Crisp and Jim Burridge. A note on nonregular likelihood functions in heteroscedastic regression models. Biometrika, 81 (3):585–587, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337129.

Cook:1997:TAR

[CB97] R. Dennis Cook and Efstathia Bura. Testing the adequacy of regression functions. Biometrika, 84(4):949–956, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337665.

Cordeiro:1994:NAD

[CBD94] Gauss M. Cordeiro, Denise A. Botter, and Silvia L. De Paula Ferrari. Nonnull asymptotic distributions of three classic criteria in generalised linear models. Biometrika, 81(4):709–720, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337074.

Chu:1995:NRE

[CC95] C. K. Chu and K. F. Cheng. Nonparametric regression estimates using misclassified binary responses. Biometrika, 82 (2):315–325, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337410.

Chambers:1992:PEF

[CDH92] R. L. Chambers, Alan H. Dorfman, and Peter Hall. Properties of estimators of the finite population distribution function. Biometrika, 79(3):577–582, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336788.

Chen:1994:WFP

[CDL94] Xiang-Hui Chen, Arthur P. Dempster, and Jun S. Liu. Weighted finite population sampling to maximize entropy. Biometrika, 81 (3):457–469, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337119.

Cordeiro:1991:MST

[CdPF91] Gauss M. Cordeiro and Silvia L. de Paula Ferrari. A modified score test statistic having chi-squared distribution to order n^{−1}. Biometrika, 78(3):573–582, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337026.

Carter:1992:CRS

[CES92] C. K. Carter, G. K. Eagleson, and B. W. Silverman. A comparison of the Reinsch and Speckman splines. Biometrika, 79 (1):81–91, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337149.

Cox:1991:RLD

[CF91] Trevor F. Cox and Gillian Ferry. Robust logistic discrimination. Biometrika, 78(4):841–849, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336936.

Caperaa:1997:NEP

[CFG97] P. Capéraà, A.-L. Fougères, and C. Genest. A nonparametric estimation procedure for bivariate extreme value copulas. Biometrika, 84(3):567–577, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337579.

Cheah:1994:MTE

[CFR94] P. K. Cheah, D. A. S. Fraser, and N. Reid. Multiparameter testing in exponential models: Third order approximations from likelihood. Biometrika, 81(2):271–278, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336957.

Carlin:1993:PLI

[CG93] Bradley P. Carlin and Alan E. Gelfand. Parametric likelihood inference for record breaking problems. Biometrika, 80(3):507– 515, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337172.

Chib:1998:AMP

[CG98] Siddhartha Chib and Edward Greenberg. Analysis of multivariate probit models. Biometrika, 85(2):347–361, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337362.

Chang:1995:VGS

[CGS95] Myron N. Chang, A. Lawrence Gould, and Steven M. Snapinn. P -values for group sequential testing. Biometrika, 82(3): 650–654, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337542.

Chernoff:1990:LOD

[CH90] Herman Chernoff and Yoel Haitovsky. Locally optimal design for comparing two probabilities from binomial data subject to misclassification. Biometrika, 77(4):797–805, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337102.

Chan:1993:MTE

[CH93] Y. M. Chan and Xuming He. On median-type estimators of direction for the von Mises–Fisher distribution. Biometrika, 80 (4):869–875, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336878.

Choi:1998:BRL

[CH98] Edwin Choi and Peter Hall. On bias reduction in local linear smoothing. Biometrika, 85(2):333–345, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337361.

Choi:1999:DSP

[CH99] Edwin Choi and Peter Hall. Data sharpening as a prelude to density estimation. Biometrika, 86(4):941–947, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673598.

Chaloner:1991:BRA

[Cha91a] Kathryn Chaloner. Bayesian residual analysis in the presence of censoring. Biometrika, 78(3):637–644, September 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337033.

Chan:1991:ODE

[Cha91b] Ling-Yau Chan. Optimal design for estimation of variance in nonparametric regression using first order differences. Biometrika, 78(4):926–929, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336947.

Chappell:1992:NLR

[Cha92a] Rick Chappell. A note on linear rank tests and Gill and Schumacher’s tests of proportionality. Biometrika, 79(1):199–201, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337161.

Chaudhuri:1992:NEV

[Cha92b] Arijit Chaudhuri. A note on estimating the variance of the regression estimator. Biometrika, 79(1):217–218, March 1992.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337166.

Chesher:1991:EME

[Che91] Andrew Chesher. The effect of measurement error. Biometrika, 78(3):451–462, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337015.

Chen:1996:ELC

[Che96] Song Xi Chen. Empirical likelihood confidence intervals for nonparametric density estimation. Biometrika, 83(2):329–341, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337604.

Cheng:1998:SHP

[Che98] Ching-Shui Cheng. Some hidden projection properties of orthogonal arrays with strength three. Biometrika, 85(2):491– 495, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337377.

Cheng:1999:TRL

[Che99] Qiansheng Cheng. On time-reversibility of linear processes. Biometrika, 86(2):483–486, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2673529.

Chiu:1990:WBS

[Chi90] Shean-Tsong Chiu. Why bandwidth selectors tend to choose smaller bandwidths, and a Remedy. Biometrika, 77(1):222– 226, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336068.

Chiu:1991:EDE

[Chi91] Shean-Tsong Chiu. The effect of discretization error on bandwidth selection for kernel density estimation. Biometrika, 78 (2):436–441, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337272.

Chiu:1992:ABS

[Chi92] Shean-Tsong Chiu. An automatic bandwidth selector for kernel density estimation. Biometrika, 79(4):771–782, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337233.

Chu:1995:MTP

[CHK95] Chia-Shang J. Chu, Kurt Hornik, and Chung-Ming Kuan. MO-SUM tests for parameter constancy. Biometrika, 82(3):603–617, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337537.

Christmann:1994:LMW

[Chr94] Andreas Christmann. Least median of weighted squares in logistic regression with large strata. Biometrika, 81(2):413–417, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336973.

Cai:1998:STH

[CHT98] Zongwu Cai, Clifford M. Hurvich, and Chih-Ling Tsai. Score tests for heteroscedasticity in wavelet regression. Biometrika, 85 (1):229–234, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337322.

Cheng:1999:HDP

[CHT99] Ming-Yen Cheng, Peter Hall, and Berwin A. Turlach. Highderivative parametric enhancements of nonparametric curve estimators. Biometrika, 86(2):417–428, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673522.

Cormack:1991:IPM

[CJ91] R. M. Cormack and P. E. Jupp. Inference for Poisson and multinomial models for capture-recapture experiments. Biometrika, 78(4):911–916, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336943.

Cavanaugh:1999:API

[CJ99] Joseph E. Cavanaugh and Wesley O. Johnson. Assessing the predictive influence of cases in a state space process. Biometrika, 86(1):183–190, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673546.

Christensen:1992:PDS

[CJP92] Ronald Christensen, Wesley Johnson, and Larry M. Pearson. Prediction diagnostics for spatial linear models. Biometrika, 79 (3):583–591, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336789.

Carter:1994:GSS

[CK94] C. K. Carter and R. Kohn. On Gibbs sampling for state space models. Biometrika, 81(3):541–553, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337125.

Carter:1996:MCM

[CK96] C. K. Carter and R. Kohn. Markov chain Monte Carlo in conditionally Gaussian state space models. Biometrika, 83(3): 589–601, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337511.

Chuang:1998:RMC

[CL98] Chin-Shan Chuang and Tze Leung Lai. Resampling methods for confidence intervals in group sequential trials. Biometrika, 85(2):317–332, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337360.

Chen:1999:TMC

[CL99a] Hua Yun Chen and Roderick Little. A test of missing completely at random for generalised estimating equations with missing data. Biometrika, 86(1):1–13, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673532.

Chen:1999:CCC

[CL99b] Kani Chen and Shaw-Hwa Lo. Case-cohort and case-control analysis with Cox’s model. Biometrika, 86(4):755–764, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673583.

Chen:1995:ATN

[CLT95] Rong Chen, Jun S. Liu, and Ruey S. Tsay. Additivity tests for nonlinear autoregression. Biometrika, 82(2):369–383, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337414.

Catchpole:1991:NSM

[CM91] E. A. Catchpole and B. J. T. Morgan. A note on Seber’s model for ring-recovery data. Biometrika, 78(4):917–919, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336944.

Catchpole:1997:DPR

[CM97] E. A. Catchpole and B. J. T. Morgan. Detecting parameter redundancy. Biometrika, 84(1):187–196, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337565.

Cooley:1998:CKP

[CM98] Craig A. Cooley and Steven N. MacEachern. Classification via kernel product estimators. Biometrika, 85(4):823–833, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337486.

Chai:1999:ODD

[CM99] Feng-Shun Chai and Rahul Mukerjee. Optimal designs for diallel crosses with specific combining abilities. Biometrika, 86(2):453– 458, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673525.

Catchpole:1998:EPR

[CMF98] E. A. Catchpole, B. J. T. Morgan, and S. N. Freeman. Estimation in parameter-redundant models. Biometrika, 85(2):462– 468, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337372.

Carroll:1999:NRP

[CMR99] Raymond J. Carroll, Jeffrey D. Maca, and David Ruppert. Nonparametric regression in the presence of measurement error. Biometrika, 86(3):541–554, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673653.

Chao:1993:SRE

[CMY93] Anne Chao, M.-C. Ma, and Mark C. K. Yang. Stopping rules and estimation for recapture debugging with unequal failure rates. Biometrika, 80(1):193–201, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336768.

Cribari-Neto:1995:SOA

[CNF95] Francisco Cribari-Neto and Silvia L. P. Ferrari. Second order asymptotics for score tests in generalised linear models. Biometrika, 82(2):426–432, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337420.

Coad:1991:STU

[Coa91] D. S. Coad. Sequential tests for an unstable response variable. Biometrika, 78(1):113–121, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336901.

Conniffe:1990:THE

[Con90] Denis Conniffe. Testing hypotheses with estimated scores. Biometrika, 77(1):97–106, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336052.

Constantine:1992:IPM

[Con92] Gregory M. Constantine. On the information and precision matrices of varietal contrasts. Biometrika, 79(1):214–216, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337165.

Corcoran:1998:BAE

[Cor98] Stephen A. Corcoran. Bartlett adjustment of empirical discrepancy statistics. Biometrika, 85(4):967–972, December 1998. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337498.

Cox:1990:DPC

[Cox90] D. R. Cox. Discussion of paper by C. B. Begg. Biometrika, 77 (3):483–484, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336984. See [Beg90b].

Cox:1993:UEE

[Cox93] D. R. Cox. Unbiased estimating equations derived from statistics that are functions of a parameter. Biometrika, 80(4):905– 909, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336882.

Cordeiro:1989:ILR

[CP89] Gauss M. Cordeiro and Gilberto A. Paula. Improved likelihood ratio statistics for exponential family nonlinear models. Biometrika, 76(1):93–100, March 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336372. See correction [CP91].

Cordeiro:1991:ACI

[CP91] Gauss M. Cordeiro and Gilberto A. Paula. Amendments and corrections: “Improved likelihood ratio statistics for exponential family nonlinear models” [Biometrika 76 (1989), no. 1, 93– 100; MR0991426 (90f:62067)]. Biometrika, 78(4):935, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336949.

See [CP89].

Cai:1995:EEH

[CP95] Jian Wen Cai and Ross L. Prentice. Estimating equations for hazard ratio parameters based on correlated failure time data. Biometrika, 82(1):151–164, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337635.

Clyde:1998:MSS

[CPV98] Merlise Clyde, Giovanni Parmigiani, and Brani Vidakovic. Multiple shrinkage and subset selection in wavelets. Biometrika, 85 (2):391–401, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337365.

Chen:1993:ELE

[CQ93] Jia Hua Chen and Jing Qin. Empirical likelihood estimation for finite populations and the effective usage of auxiliary information. Biometrika, 80(1):107–116, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336761.

Cox:1992:NDB

[CR92] D. R. Cox and N. Reid. A note on the difference between profile and modified profile likelihood. Biometrika, 79(2):408–411, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336853.

Carriere:1993:OTP

[CR93] Keumhee Chough Carrière and Gregory C. Reinsel. Optimal two-period repeated measurement designs with two or more treatments. Biometrika, 80(4):924–929, December 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336886.

Casella:1996:RBS

[CR96] George Casella and Christian P. Robert. Rao–Blackwellisation of sampling schemes. Biometrika, 83(1):81–94, March 1996. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337434.

Crowder:1990:SNT

[Cro90] Martin Crowder. On some nonregular tests for a modified Weibull model. Biometrika, 77(3):499–506, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336987.

Crowder:1995:UWC

[Cro95] Martin Crowder. On the use of a working correlation matrix in using generalised linear models for repeated measures. Biometrika, 82(2):407–410, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337417.

Cheng:1991:TRT

[CS91] Ching-Shui Cheng and David M. Steinberg. Trend robust twolevel factorial designs. Biometrika, 78(2):325–336, June 1991.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337259.

Chesher:1995:BCL

[CS95] Andrew Chesher and Richard J. Smith. Bartlett corrections to likelihood ratio tests. Biometrika, 82(2):433–436, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337421.

Chen:1998:MCM

[CS98] Ming-Hui Chen and Qi-Man Shao. Monte Carlo methods for Bayesian analysis of constrained parameter problems. Biometrika, 85(1):73–87, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337310.

Chandra:1992:MSF

[CSB92] K. Suresh Chandra, S. Sampath, and G. K. Balasubramani. Markov sampling for finite populations. Biometrika, 79(1):210– 213, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337164.

Chan:1998:LPL

[CT98] K. S. Chan and Ruey S. Tsay. Limiting properties of the least squares estimator of a continuous threshold autoregressive model. Biometrika, 85(2):413–426, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337367.

Currie:1996:NBJ

[Cur96] Iain D. Currie. A note on Buckley–James estimators for censored data. Biometrika, 83(4):912–915, December 1996. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337294.

Cheng:1991:UUM

[CV91a] Chi-Lun Cheng and John W. Van Ness. On the unreplicated ultrastructural model. Biometrika, 78(2):442–445, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337273.

Critchley:1991:IOM

[CV91b] Frank Critchley and Cecilia Vitiello. The influence of observations on misclassification probability estimates in linear discriminant analysis. Biometrika, 78(3):677–690, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337037.

Cox:1990:AML

[CW90] D. R. Cox and Nanny Wermuth. An approximation to maximum likelihood estimates in reduced models. Biometrika, 77(4): 747–761, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337098.

Cox:1992:RMM

[CW92] D. R. Cox and Nanny Wermuth. Response models for mixed binary and quantitative variables. Biometrika, 79(3):441–461, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336778.

Cook:1994:TRV

[CW94a] R. Dennis Cook and Sanford Weisberg. Transforming a response variable for linearity. Biometrika, 81(4):731–737, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337076.

Cox:1994:NQE

[CW94b] D. R. Cox and Nanny Wermuth. A note on the quadratic exponential binary distribution. Biometrika, 81(2):403–408, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336971.

Cox:1995:DCD

[CW95] D. R. Cox and Nanny Wermuth. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4): 688–689, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337330. See [Pea95a].

Coad:1996:CCI

[CW96] D. S. Coad and M. B. Woodroofe. Corrected confidence intervals after sequential testing with applications to survival analysis.

Biometrika, 83(4):763–777, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337281.

Cheng:1995:ATM

[CWY95] S. C. Cheng, L. J. Wei, and Z. Ying. Analysis of transformation models with censored data. Biometrika, 82(4):835–845, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337348.

Carey:1993:MMB

[CZD93] Vincent Carey, Scott L. Zeger, and Peter Diggle. Modelling multivariate binary data with alternating logistic regressions. Biometrika, 80(3):517–526, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337173.

Datta:1996:PPF

[Dat96] Gauri Sankar Datta. On priors providing frequentist validity of Bayesian inference for multiple parametric functions. Biometrika, 83(2):287–298, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337601.

Davison:1986:APL

[Dav86] A. C. Davison. Approximate predictive likelihood. Biometrika, 73(2):323–332, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2336208. See correction [Dav90b].

Davidian:1990:EVF

[Dav90a] M. Davidian. Estimation of variance functions in assays with possibly unequal replication and nonnormal data. Biometrika, 77(1):43–54, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336048.

Davison:1990:ACA

[Dav90b] A. C. Davison. Amendments and corrections: “Approximate predictive likelihood” [Biometrika 73 (1986), no. 2, 323–332; MR0855892 (87i:62014)]. Biometrika, 77(3):667, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337009.

See [Dav86].

Davison:1992:TEH

[Dav92] A. C. Davison. Treatment effect heterogeneity in paired data. Biometrika, 79(3):463–474, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336779.

Dawid:1992:E

[Daw92] A. P. Dawid. Editorial. Biometrika, 79(3):??, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336777.

Dawid:1993:E

[Daw93] A. P. Dawid. Editorial. Biometrika, 80(1):1, March 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336753.

Dawid:1995:DCD

[Daw95] A. P. Dawid. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):689–690, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337331. See

[Pea95a].

Darroch:1994:SAR

[DB94] J. N. Darroch and M. Borkent. Synergism, attributable risk and interaction for two binary exposure factors. Biometrika, 81 (2):259–270, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336956.

Day:1997:ABF

[DBL97] Roger Day, John Bryant, and Myrto Lefkopoulou. Adaptation of bivariate frailty models for prediction, with application to biological markers as prognostic indicators. Biometrika, 84(1): 45–56, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337554.

Deng:1991:ADU

[DC91] Lih-Yuan Deng and Raj S. Chhikara. On asymptotically designunbiased estimators of a finite population mean. Biometrika, 78 (1):189–195, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336909.

DeJong:1994:FLE

[DCCL94] Piet De Jong and Singfat Chu-Chun-Lin. Fast likelihood evaluation and prediction for nonstationary state space models. Biometrika, 81(1):133–142, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337056.

Dannenberg:1994:EWA

[DDM94] Olaf Dannenberg, Holger Dette, and Axel Munk. An extension of Welch’s approximate t-solution to comparative bioequivalence trials. Biometrika, 81(1):91–101, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337053.

Deshpande:1993:TST

[DDNN93] Jayant V. Deshpande, Avinash D. Dharmadhikari, and Uttara V. Naik-Nimbalkar. Two-sample tests based on cumulative incidence functions from coherent systems. Biometrika, 80(4): 877–883, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336879.

DeJong:1997:SST

[De 97] Piet De Jong. The scan sampler for time series models. Biometrika, 84(4):929–937, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337663.

DiCiccio:1992:MAC

[DE92] Thomas DiCiccio and Bradley Efron. More accurate confidence intervals in exponential families. Biometrika, 79(2):231– 245, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336835.

Dewanji:1992:NTC

[Dew92] Anup Dewanji. A note on a test for competing risks with missing failure type. Biometrika, 79(4):855–857, December 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337244.

DiCiccio:1991:AMA

[DF91] Thomas J. DiCiccio and Christopher A. Field. An accurate method for approximate conditional and Bayesian inference about linear regression models from censored data. Biometrika, 78(4):903–910, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336942.

Dellaportas:1999:MCM

[DF99] Petros Dellaportas and Jonathan J. Forster. Markov chain Monte Carlo model determination for hierarchical and graphical log-linear models. Biometrika, 86(3):615–633, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673658.

DiCiccio:1990:AMT

[DFF90] Thomas J. DiCiccio, Christopher A. Field, and D. A. S. Fraser. Approximations of marginal tail probabilities and inference for scalar parameters. Biometrika, 77(1):77–95, March 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336051.

Davidian:1993:NME

[DG93] Marie Davidian and A. Ronald Gallant. The nonlinear mixed effects model with a smooth random effects density. Biometrika, 80(3):475–488, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337169.

Datta:1995:PPF

[DG95] Gauri Sankar Datta and Jayanta Kumar Ghosh. On priors providing frequentist validity for Bayesian inference. Biometrika, 82(1):37–45, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337625.

DeLong:1994:ECS

[DGS94] David M. DeLong, George H. Guirguis, and Ying C. So. Efficient computation of subset selection probabilities with application to Cox regression. Biometrika, 81(3):607–611, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337133.

Do:1991:IRB

[DH91] Kim-Anh Do and Peter Hall. On importance resampling for the bootstrap. Biometrika, 78(1):161–167, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336906.

Davison:1992:BVB

[DH92] A. C. Davison and Peter Hall. On the bias and variability of bootstrap and cross-validation estimates of error rate in discrimination problems. Biometrika, 79(2):279–284, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336839.

Dette:1994:ODL

[DH94] Holger Dette and Linda M. Haines. E-optimal designs for linear and nonlinear models with two parameters. Biometrika, 81(4): 739–754, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337077.

Dykstra:1999:NID

[DHR99] Richard Dykstra, John Hewett, and Tim Robertson. Nonparametric, isotonic discriminant procedures. Biometrika, 86 (2):429–438, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673523.

Davison:1992:BL

[DHW92] A. C. Davison, D. V. Hinkley, and B. J. Worton. Bootstrap likelihoods. Biometrika, 79(1):113–130, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337152.

Dietrich:1991:MRL

[Die91] C. R. Dietrich. Modality of the restricted likelihood for spatial Gaussian random fields. Biometrika, 78(4):833–839, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336935.

Ding:1996:ANM

[Din96] Ye Ding. On the asymptotic normality of multinomial population size estimators with application to backcalculation of AIDS epidemic. Biometrika, 83(3):695–699, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337521.

Donoho:1994:ISA

[DJ94] David L. Donoho and Iain M. Johnstone. Ideal spatial adaptation by wavelet shrinkage. Biometrika, 81(3):425–455, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337118.

deJong:1995:SST

[dJS95] Piet de Jong and Neil Shephard. The simulation smoother for time series models. Biometrika, 82(2):339–350, June 1995. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337412.

Das:1991:COP

[DK91] Ashish Das and Sanpei Kageyama. A class of E-optimal proper efficiency-balanced designs. Biometrika, 78(3):693–696, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337039.

Durbin:1997:MCM

[DK97] J. Durbin and S. J. Koopman. Monte Carlo maximum likelihood estimation for non-Gaussian state space models. Biometrika, 84 (3):669–684, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337587.

deLeon:1991:OED

[dLA91] A. C. Ponce de Leon and A. C. Atkinson. Optimum experimental design for discriminating between two rival models in the presence of prior information. Biometrika, 78(3):601–608, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337029.

deLeeuw:1991:RRM

[dLvdH91] Jan de Leeuw and Peter G. M. van der Heijden. Reduced rank models for contingency tables. Biometrika, 78(1):229– 232, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336915.

Darroch:1990:STS

[DM90] J. N. Darroch and P. I. McCloud. Separating two sources of dependence in repeated influenza outbreaks. Biometrika, 77 (2):237–243, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336801.

DiCiccio:1991:AMT

[DM91] Thomas J. DiCiccio and Michael A. Martin. Approximations of marginal tail probabilities for a class of smooth functions with applications to Bayesian and conditional inference. Biometrika, 78(4):891–902, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336941.

DeFeo:1992:NLE

[DM92a] Patrick DeFeo and Raymond H. Myers. A new look at experimental design robustness. Biometrika, 79(2):375–380, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336848.

Dryden:1992:SSA

[DM92b] I. L. Dryden and K. V. Mardia. Size and shape analysis of landmark data. Biometrika, 79(1):57–68, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337147.

Drake:1995:NBD

[DM95] Christiana Drake and Allan McQuarrie. A note on the bias due to omitted confounders. Biometrika, 82(3):633–638, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337539.

Dey:1996:OBD

[DM96] Aloke Dey and Chand K. Midha. Optimal block designs for diallel crosses. Biometrika, 83(2):484–489, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337620.

Dawid:1998:FII

[DM98] A. Philip Dawid and Julia Mortera. Forensic identification with imperfect evidence. Biometrika, 85(4):835–849, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337487. See

corrections [DM99].

Dawid:1999:ACF

[DM99] A. Philip Dawid and Julia Mortera. Amendments and corrections:“Forensic identification with imperfect evidence” [Biometrika 85 (1998), no. 4, 835–849; (2000a:62285)]. Biometrika, 86(4):974, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2673603. See [DM98].

Denison:1998:BCA

[DMS98] David G. T. Denison, Bani K. Mallick, and Adrian F. M. Smith. A Bayesian CART algorithm. Biometrika, 85(2):363–377, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337363.

DiCiccio:1992:FAA

[DMY92] Thomas J. DiCiccio, Michael A. Martin, and G. Alastair Young. Fast and accurate approximate double bootstrap confidence intervals. Biometrika, 79(2):285–295, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336840.

DiCiccio:1993:AAC

[DMY93] Thomas J. DiCiccio, Michael A. Martin, and G. Alastair Young. Analytical approximations to conditional distribution functions. Biometrika, 80(4):781–790, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336869.

Dette:1999:OCR

[DO99] Holger Dette and Timothy E. O’Brien. Optimality criteria for regression models based on predicted variance. Biometrika, 86 (1):93–106, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673539.

Donegani:1991:APR

[Don91] Michel Donegani. An adaptive and powerful randomization test. Biometrika, 78(4):930–933, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336948.

Dougherty:1997:ECL

[Dou97] George G. Dougherty, Jr. Estimation of common linear functional relationships in k data sets. Biometrika, 84(1):103–110, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337559.

Didelez:1998:MLE

[DP98] Vanessa Didelez and Iris Pigeot. Maximum likelihood estimation in graphical models with missing values. Biometrika, 85(4): 960–966, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337497.

Droge:1998:MRA

[Dro98] Bernd Droge. Minimax regret analysis of orthogonal series regression estimation: Selection versus shrinkage. Biometrika, 85 (3):631–643, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337392.

DiCiccio:1993:BAA

[DS93] Thomas J. DiCiccio and Steven E. Stern. On Bartlett adjustments for approximate Bayesian inference. Biometrika, 80(4): 731–740, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336865.

Deshpande:1995:THP

[DS95]  Jayant V. Deshpande and Debasis Sengupta.  Testing

the hypothesis of proportional hazards in two populations. Biometrika, 82(2):251–261, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337405.

Deville:1998:UPS

[DT98] Jean-Claude Deville and Yves Tillé. Unequal probability sampling without replacement through a splitting method. Biometrika, 85(1):89–101, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337311.

Diks:1999:TSM

[DT99] Cees Diks and Howell Tong. A test for symmetries of multivariate probability distributions. Biometrika, 86(3):605–614, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673657.

Duan:1993:SAB

[Dua93] Naihua Duan. Sensitivity analysis for Box–Cox power transformation model: contrast parameters. Biometrika, 80(4):885–897, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336880.

Duffy:1990:CCR

[Duf90] Diane E. Duffy. On continuity-corrected residuals in logistic regression. Biometrika, 77(2):287–293, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336806.

Dupuis:1995:BEM

[Dup95] Jerome A. Dupuis. Bayesian estimation of movement and survival probabilities from capture-recapture data. Biometrika, 82 (4):761–772, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337343.

Dette:1996:ROE

[DW96] Holger Dette and Weng Kee Wong. Robust optimal extrapolation designs. Biometrika, 83(3):667–680, September 1996. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337517.

Dette:1998:BOD

[DW98] Holger Dette and Weng Kee Wong. Bayesian D-optimal designs on a fixed number of design points for heteroscedastic polynomial models. Biometrika, 85(4):869–882, December 1998. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337489.

Daniels:1991:SAS

[DY91] H. E. Daniels and G. A. Young. Saddlepoint approximation for the Studentized mean, with an application to the bootstrap. Biometrika, 78(1):169–179, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336907.

Emerson:1990:PEF

[EF90] Scott S. Emerson and Thomas R. Fleming. Parameter estimation following group sequential hypothesis testing. Biometrika, 77(4):875–892, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337110.

Efron:1993:BLC

[Efr93] Bradley Efron. Bayes and likelihood calculations from confidence intervals. Biometrika, 80(1):3–26, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336754.

Efron:1997:LHS

[Efr97] Bradley Efron. The length heuristic for simultaneous hypothesis tests. Biometrika, 84(1):143–157, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337562.

Eaton:1996:PAB

[EGS96] Morris L. Eaton, Alessandra Giovagnoli, and Paola Sebastiani. A predictive approach to the Bayesian design problem with application to normal regression models. Biometrika, 83(1):111– 125, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337436.

Ebrahimi:1990:TWS

[EH90] Nader Ebrahimi and Mohamed Habibullah. Testing whether the survival distribution is new better than used of specified age. Biometrika, 77(1):212–215, March 1990. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336065.

Eubank:1993:CCN

[EH93] R. L. Eubank and Jeffrey D. Hart. Commonality of Cusum, von Neumann and smoothing-based goodness-of-fit tests. Biometrika, 80(1):89–98, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336759.

Eales:1992:IMD

[EJ92] John D. Eales and Christopher Jennison. An improved method for deriving optimal one-sided group sequential tests. Biometrika, 79(1):13–24, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337143.

Ebrahimi:1996:CPH

[EK96] Nader Ebrahimi and S. N. U. A. Kirmani. A characterisation of the proportional hazards model through a measure of discrimination between two residual life distributions. Biometrika, 83 (1):233–235, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337445.

Engen:1997:SSC

[EL97] Steinar Engen and Magnar Lillegård. Stochastic simulations conditioned on sufficient statistics. Biometrika, 84(1):235–240, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337570.

Elster:1995:SCD

[EN95] Clemens Elster and Arnold Neumaier. Screening by conference designs. Biometrika, 82(3):589–602, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337536.

Ebrahimi:1990:RIL

[ES90a] Nader Ebrahimi and Ehsan S. Soofi. Relative information loss under Type II censored exponential data. Biometrika, 77 (2):429–435, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336827.

Eubank:1990:CFP

[ES90b] R. L. Eubank and Paul Speckman. Curve fitting by polynomialtrigonometric regression. Biometrika, 77(1):1–9, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336044.

Ekholm:1995:MRA

[ESM95] Anders Ekholm, Peter W. F. Smith, and John W. McDonald. Marginal regression analysis of a multivariate binary response. Biometrika, 82(4):847–854, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337349.

Farewell:1994:ERD

[FADD94] V. T. Farewell, O. O. Aalen, D. De Angelis, and N. E. Day. Estimation of the rate of diagnosis of HIV infection in HIV infected individuals. Biometrika, 81(2):287–294, June 1994. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336959.

Fahrmeir:1994:DMP

[Fah94] Ludwig Fahrmeir. Dynamic modelling and penalized likelihood estimation for discrete time survival data. Biometrika, 81(2):317–330, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336962.

Farebrother:1990:SHP

[Far90] R. W. Farebrother. Studies in the history of probability and statistics. XLII. Further details of contacts between Boscovich and Simpson in June 1760. Biometrika, 77(2):397–400, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336819.

Ferrari:1997:LPT

[FBCN97] Silvia L. P. Ferrari, Denise A. Botter, and Francisco Cribari-Neto. Local power of three classic criteria in generalised linear models with unknown dispersion. Biometrika, 84(2):482– 485, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337473.

Fu:1990:LCS

[FC90a] Yun-Xin Fu and R. N. Curnow. Locating a changed segment in a sequence of Bernoulli variables. Biometrika, 77(2):295– 304, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2336807. See correction [FC91].

Fu:1990:MLE

[FC90b] Yun-Xin Fu and R. N. Curnow. Maximum likelihood estimation of multiple change points. Biometrika, 77(3):563–573, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336995.

Fu:1991:ACL

[FC91] Yun-Xin Fu and R. N. Curnow. Amendments and corrections: “Locating a changed segment in a sequence of Bernoulli variables” [Biometrika 77 (1990), no. 2, 295–304; (91f:62038)]. Biometrika, 78(1):233, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336916. See [FC90a].

Frisen:1991:OS

[FD91] Marianne Frisén and Jacques De Maré. Optimal surveillance. Biometrika, 78(2):271–280, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337252.

Farrow:1993:BLM

[FG93] Malcolm Farrow and Michael Goldstein. Bayes linear methods for grouped multivariate repeated measurement studies with application to crossover trials. Biometrika, 80(1):39–59, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336756.

Firth:1991:MCN

[FGH91] D. Firth, J. Glosup, and D. V. Hinkley. Model checking with nonparametric curves. Biometrika, 78(2):245–252, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337249.

Fraser:1991:CIT

[FGS91] D. A. S. Fraser, I. Guttman, and M. S. Srivastava. Conditional inference for treatment and error in multivariate analysis. Biometrika, 78(3):565–572, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337025.

Fienberg:1995:DCD

[FGS95] Stephen E. Fienberg, Clark Glymour, and Peter Spirtes. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):690–692, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337332. See [Pea95a].

Firth:1991:QLM

[FH91] D. Firth and I. R. Harris. Quasi-likelihood for multiplicative random effects. Biometrika, 78(3):545–555, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337023.

Findley:1986:UMA

[Fin86] David F. Findley. The uniqueness of moving average representations with independent and identically distributed random variables for non-Gaussian stationary time series. Biometrika, 73 (2):520–521, August 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2336232. See correction [Fin90].

Findley:1990:ACU

[Fin90] David F. Findley. Amendments and corrections: “The uniqueness of moving average representations with independent and identically distributed random variables for non-Gaussian stationary time series” [Biometrika 73 (1986), no. 2, 520–521; MR0855914 (88e:62210)]. Biometrika, 77(1):235, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336071. See

[Fin86].

Firth:1993:BRM

[Fir93] David Firth. Bias reduction of maximum likelihood estimates. Biometrika, 80(1):27–38, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336755. See corrections [Fir95].

Firth:1995:ACB

[Fir95] David Firth. Amendments and corrections: “Bias reduction of maximum likelihood estimates” [Biometrika 80 (1993), no. 1, 27–38; (94b:62017)]. Biometrika, 82(3):667, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337546. See

[Fir93].

Faddy:1998:SSD

[FJ98] M. J. Faddy and M. C. Jones. Semiparametric smoothing for discrete data. Biometrika, 85(1):131–138, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337314.

Faraggi:1996:CRF

[FK96] David Faraggi and Edward L. Korn. Competing risks with frailty models when treatment affects only one failure type. Biometrika, 83(2):467–471, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337617.

Fahrmeir:1998:NMH

[FK98] Ludwig Fahrmeir and Artur Klinger. A nonparametric multiplicative hazard model for event history analysis. Biometrika, 85(3):581–592, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337388.

Fitzmaurice:1993:LBM

[FL93] Garrett M. Fitzmaurice and Nan M. Laird. A likelihood-based method for analysing longitudinal binary responses. Biometrika, 80(1):141–151, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336764.

Fraser:1990:NLR

[FLR90] D. A. S. Fraser, H. S. Lee, and N. Reid. Nonnormal linear regression; an example of significance levels in high dimensions. Biometrika, 77(2):333–341, June 1990. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336811.

Flury:1990:PP

[Flu90] Bernhard A. Flury. Principal points. Biometrika, 77(1):33– 41, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336047.

Frigessi:1997:CCM

[FMS97] Arnoldo Frigessi, Fabio Martinelli, and Julian Stander. Computational complexity of Markov chain Monte Carlo methods for finite Markov random fields. Biometrika, 84(1):1–18, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337551.

Fong:1990:BAS

[Fon90] Wai Kwan Fong. A Bayesian approach to successive sampling with partial replacement of units on two occasions. Biometrika, 77(2):383–388, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336817.

Fernandez:1994:CML

[FOS94] Carmen Fernández, Jacek Osiewalski, and Mark F. J. Steel. The continuous multivariate location-scale model revisited: A tale of robustness. Biometrika, 81(3):588–594, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337130.

Fujikoshi:1991:SCG

[FR91] Yasunori Fujikoshi and C. Radhakrishna Rao. Selection of covariables in the growth curve model. Biometrika, 78(4):779–785, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336929.

Ferretti:1996:URB

[FR96] Nelida Ferretti and Juan Romo. Unit root bootstrap tests for AR (1) models. Biometrika, 83(4):849–860, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337288.

Frangakis:1999:ACI

[FR99] Constantine E. Frangakis and Donald B. Rubin. Addressing complications of intention-to-treat analysis in the combined presence of all-or-none treatment-noncompliance and subsequent missing outcomes. Biometrika, 86(2):365–379, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673518.

Fraser:1990:TPO

[Fra90] D. A. S. Fraser. Tail probabilities from observed likelihoods. Biometrika, 77(1):65–76, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336050.

Freedman:1995:DCD

[Fre95] David Freedman. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):692–693, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337333.

See [Pea95a].

Fraser:1999:SGF

[FRW99] D. A. S. Fraser, N. Reid, and J. Wu. A simple general formula for tail probabilities for frequentist and Bayesian inference. Biometrika, 86(2):249–264, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673509.

Frydman:1995:NEM

[Fry95] Halina Frydman. Nonparametric estimation of a Markov “illness-death” process from interval-censored observations, with application to diabetes survival data. Biometrika, 82(4): 773–789, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337344.

Flehinger:1998:SCR

[FRY98] Betty J. Flehinger, Benjamin Reiser, and Emmanuel Yashchin. Survival with competing risks and masked causes of failures. Biometrika, 85(1):151–164, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337316.

Farewell:1990:DPC

[FS90] V. T. Farewell and D. A. Sprott. Discussion of paper by C. B. Begg. Biometrika, 77(3):479–480, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336981. See [Beg90b].

Fujikoshi:1997:MAM

[FS97] Yasunori Fujikoshi and Kenichi Satoh. Modified AIC and C_{p} in multivariate linear regression. Biometrika, 84(3):707–716, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337590.

Fernandez:1999:MSR

[FS99] Carmen Fernandez and Mark F. J. Steel. Multivariate Student– t regression models: pitfalls and inference. Biometrika, 86(1): 153–167, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673544.

Fukuchi:1999:SMS

[Fuk99] Jun-Ichiro Fukuchi. Subsampling and model selection in time series analysis. Biometrika, 86(3):591–604, September 1999. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673656.

Foster:1998:AC

[FV98] Dean P. Foster and Rakesh V. Vohra. Asymptotic calibration. Biometrika, 85(2):379–390, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337364.

Fan:1998:EEC

[FY98] Jianqing Fan and Qiwei Yao. Efficient estimation of conditional variance functions in stochastic regression. Biometrika, 85(3): 645–660, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337393.

Fan:1996:ECD

[FYT96] Jianqing Fan, Qiwei Yao, and Howell Tong. Estimation of conditional densities and sensitivity measures in nonlinear dynamical systems. Biometrika, 83(1):189–206, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337441.

Fine:1998:LTM

[FYW98] J. P. Fine, Z. Ying, and L. J. Wei. On the linear transformation model for censored data. Biometrika, 85(4):980–986, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337500.

Gabriel:1998:GBR

[Gab98] K. Ruben Gabriel. Generalised bilinear regression. Biometrika, 85(3):689–700, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337396.

Gamerman:1992:DAS

[Gam92] Dani Gamerman. A dynamic approach to the statistical analysis of point processes. Biometrika, 79(1):39–50, March 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337145.

Gamerman:1998:MCM

[Gam98] Dani Gamerman. Markov chain Monte Carlo for dynamic generalised linear models. Biometrika, 85(1):215–227, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337321.

Gelfand:1990:NOE

[GD90] A. E. Gelfand and S. R. Dalal. A note on overdispersed exponential families. Biometrika, 77(1):55–64, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336049.

Gange:1996:SMC

[GD96] Stephen J. Gange and David L. DeMets. Sequential monitoring of clinical trials with correlated responses. Biometrika, 83 (1):157–167, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337439.

Ghosh:1997:TAC

[GD97] D. K. Ghosh and J. Divecha. Two associate class partially balanced incomplete block designs and partial diallel crosses.

Biometrika, 84(1):245–248, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337572.

Gasser:1990:CWK

[GE90] Theo Gasser and Joachim Engel. The choice of weights in kernel regression estimation. Biometrika, 77(2):377–381, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336816.

Gu:1999:MGC

[GFG99] Minggao Gu, Dean Follmann, and Nancy L. Geller. Monitoring a general class of two-sample survival statistics with applications. Biometrika, 86(1):45–57, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673535.

Gentleman:1994:MLI

[GG94] Robert Gentleman and Charles J. Geyer. Maximum likelihood for interval censored data: Consistency and computation. Biometrika, 81(3):618–623, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337135.

Garvan:1997:NPD

[GG97] Cynthia Wilson Garvan and Malay Ghosh. Noninformative priors for dispersion models. Biometrika, 84(4):976–982, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337669.

Gelfand:1998:MCM

[GG98] Alan E. Gelfand and Sujit K. Ghosh. Model choice: A minimum posterior predictive loss approach. Biometrika, 85(1): 1–11, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337305.

Giudici:1999:DGG

[GG99] Paolo Giudici and Peter J. Green. Decomposable graphical Gaussian model determination. Biometrika, 86(4):785–801, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673585.

Gillespie:1992:CBF

[GGI92] Brenda Gillespie, John A. Gillespie, and Boris Iglewicz. A comparison of the bias in four versions of the product-limit estimator. Biometrika, 79(1):149–155, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337155.

Genest:1995:SEP

[GGR95] C. Genest, K. Ghoudi, and L.-P. Rivest. A semiparametric estimation procedure of dependence parameters in multivariate families of distributions. Biometrika, 82(3):543–552, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337532.

Gay:1990:CRF

[GH90a] Roger Gay and C. C. Heyde. On a class of random field models which allows long range dependence. Biometrika, 77(2):401– 403, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336820.

Gombay:1990:ADM

[GH90b] Edit Gombay and Lajos Horváth. Asymptotic distributions of maximum likelihood tests for change in the mean. Biometrika, 77(2):411–414, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336823.

Ghosal:1999:PMP

[Gho99] Subhashis Ghosal. Probability matching priors for non-regular cases. Biometrika, 86(4):956–964, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673600.

Gelfand:1991:NBB

[GK91] Alan E. Gelfand and Lynn Kuo. Nonparametric Bayesian bioassay including ordered polytomous response. Biometrika, 78(3): 657–666, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337035.

Godambe:1993:OEU

[GK93] V. P. Godambe and S. Kunte. Optimal estimation under biased allocation of treatments. Biometrika, 80(4):797–806, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336871.

Gupta:1994:OCD

[GK94] Sudhir Gupta and Sanpei Kageyama. Optimal complete diallel crosses. Biometrika, 81(2):420–424, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336975.

Gastwirth:1998:DSS

[GKR98] Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. Dual and simultaneous sensitivity analysis for matched pairs. Biometrika, 85(4):907–920, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337492.

Gerami:1992:CDS

[GL92] A. Gerami and S. M. Lewis. Comparing dual with single treatments in block designs. Biometrika, 79(3):603–610, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336791.

Gu:1994:NIR

[GL94] Minggao Gu and Jiang Lu. A note on identifiability of the regression parameter and validity of the partial likelihood approach in general relative risk regression. Biometrika, 81(4): 802–806, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337085.

Gallant:1997:ESD

[GL97] A. Ronald Gallant and Jonathan R. Long. Estimating stochastic differential equations efficiently by minimum chi-squared. Biometrika, 84(1):125–141, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337561.

Gleser:1992:NAF

[Gle92] Leon Jay Gleser. A note on the analysis of familial data. Biometrika, 79(2):412–415, June 1992. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336854. See corrections [Gle96].

Gleser:1996:CNA

[Gle96] Leon Jay Gleser. Corrections: “A note on the analysis of familial data” [Biometrika 79 (1992), no. 2, 412–415; (93f:62145)]. Biometrika, 83(1):247, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337449. See [Gle92].

Glidden:1999:CAG

[Gli99] David V. Glidden. Checking the adequacy of the gamma frailty model for multivariate failure times. Biometrika, 86(2):381– 393, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673519.

Goffinet:1992:TNM

[GLL92] Bruno Goffinet, Patrice Loisel, and Beatrice Laurent. Testing in normal mixture models when the proportions are known. Biometrika, 79(4):842–846, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337241.

Glonek:1996:CRM

[Glo96] G. F. V. Glonek. A class of regression models for multivariate categorical responses. Biometrika, 83(1):15–28, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337429.

Gilbert:1999:MLE

[GLV99] Peter B. Gilbert, Subhash R. Lele, and Yehuda Vardi. Maximum likelihood estimation in semiparametric selection bias models with application to AIDS vaccine trials. Biometrika, 86(1):27– 43, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673534.

Godambe:1991:OEF

[God91] V. P. Godambe. Orthogonality of estimating functions and nuisance parameters. Biometrika, 78(1):143–151, March 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336904.

Goldstein:1991:NMM

[Gol91] Harvey Goldstein. Nonlinear multilevel models, with an application to discrete response data. Biometrika, 78(1):45–51, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336894.

Goudie:1990:LBS

[Gou90] I. B. J. Goudie. A likelihood-based stopping rule for recapture debugging. Biometrika, 77(1):203–206, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336063.

Gower:1990:TDB

[Gow90] J. C. Gower. Three-dimensional biplots. Biometrika, 77(4): 773–785, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337100.

Gower:1992:GB

[Gow92] J. C. Gower. Generalized biplots. Biometrika, 79(3):475–493, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336780.

Geoffard:1995:ECC

[GP95] Pierre-Yves Geoffard and Tomas Philipson. The empirical content of canonical models of infectious diseases: The proportional hazard specification. Biometrika, 82(1):101–111, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337631.

Gutierrez-Pena:1997:MCP

[GP97] E. Gutiérrez-Peña. Moments for the canonical parameter of an exponential family under a conjugate distribution. Biometrika, 84(3):727–732, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337592.

Goetghebeur:1990:MLR

[GR90] Els Goetghebeur and Louise Ryan. A modified log rank test for competing risks with missing failure type. Biometrika, 77 (1):207–211, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336064.

George:1992:CRE

[GR92] Edward I. George and Christian P. Robert. Capture-recapture estimation via Gibbs sampling. Biometrika, 79(4):677–683, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337223.

Giron:1994:BKF

[GR94] F. J. Girón and J. C. Rojano. Bayesian Kalman filtering with elliptically contoured errors. Biometrika, 81(2):390–395, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336969.

Goetghebeur:1995:ACR

[GR95] Els Goetghebeur and Louise Ryan. Analysis of competing risks survival data when some failure types are missing. Biometrika, 82(4):821–833, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337347.

Goutis:1998:MCG

[GR98] Constantinos Goutis and Christian P. Robert. Model choice in generalised linear models: a Bayesian approach via Kullback– Leibler projections. Biometrika, 85(1):29–37, March 1998. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337307.

Grambsch:1994:SRT

[Gra94a] Patricia M. Grambsch. Simple robust tests for scale differences in paired data. Biometrika, 81(2):359–372, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336966.

Gray:1994:KMI

[Gra94b] Robert J. Gray. A kernel method for incorporating information on disease progression in the analysis of survival. Biometrika, 81 (3):527–539, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337124.

Green:1995:RJM

[Gre95] Peter J. Green. Reversible jump Markov chain Monte Carlo computation and Bayesian model determination. Biometrika, 82(4):711–732, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337340.

Grieve:1994:FNS

[Gri94] Andrew P. Grieve. A further note on sampling to locate rare defectives with strong prior evidence. Biometrika, 81(4):787–789, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337082.

Grieve:1995:EBA

[Gri95] Andrew P. Grieve. Extending a Bayesian analysis of the twoperiod crossover to accommodate missing data. Biometrika, 82 (2):277–286, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337407.

Guilkey:1991:ITU

[GS91] David K. Guilkey and Peter Schmidt. An interpolation test for a unit root in the presence of drift. Biometrika, 78(4):809–815, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336932.

Glad:1995:BAS

[GS95] Ingrid K. Glad and Giovanni Sebastiani. A Bayesian approach to synthetic magnetic resonance imaging. Biometrika, 82(2): 237–250, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337404.

Gelfand:1995:EPN

[GSC95] Alan E. Gelfand, Sujit K. Sahu, and Bradley P. Carlin. Efficient parametrisations for normal linear mixed models. Biometrika, 82(3):479–488, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337527.

Grambsch:1994:PHT

[GT94] Patricia M. Grambsch and Terry M. Therneau. Proportional hazards tests and diagnostics based on weighted residuals. Biometrika, 81(3):515–526, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337123. See correction [GT95].

Grambsch:1995:ACP

[GT95] Patricia M. Grambsch and Terry M. Therneau. Amendments and corrections: “Proportional hazards tests and diagnostics based on weighted residuals” [Biometrika 81 (1994), no. 3, 515–526]. Biometrika, 82(3):668, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337547. See [GT94].

Guay:1995:CMM

[Gua95] M. Guay. Curvature measures for multiresponse regression models. Biometrika, 82(2):411–417, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337418.

Goffinet:1996:OIS

[GW96] Bruno Goffinet and Daniel Wallach. Optimised importance sampling quantile estimation. Biometrika, 83(4):791–800, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337283.

Goldstein:1998:AEB

[GW98] Michael Goldstein and David A. Wooff. Adjusting exchangeable beliefs. Biometrika, 85(1):39–54, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337308.

Galbraith:1994:SNE

[GZW94] John W. Galbraith and Victoria Zinde-Walsh. A simple noniterative estimator for moving average models. Biometrika, 81 (1):143–155, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337057.

Galbraith:1997:SSA

[GZW97] John W. Galbraith and Victoria Zinde-Walsh. On some simple, autoregression-based estimation and identification techniques for ARMA models. Biometrika, 84(3):685–696, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337588.

Haddad:1995:CFL

[Had95] John N. Haddad. On the closed form of the likelihood function of the first order moving average model. Biometrika, 82(1):232– 234, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337644.

Hall:1989:TUR

[Hal89] Alastair Hall. Testing for a unit root in the presence of moving average errors. Biometrika, 76(1):49–56, March 1989. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336368. See correction

[Hal90a].

Hall:1990:ACT

[Hal90a] Alastair Hall. Amendments and corrections: “Testing for a unit root in the presence of moving average errors” [Biometrika 76 (1989), no. 1, 49–56; MR0991422 (90g:62218)]. Biometrika, 77(1):235, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336072. See [Hal89].

Hall:1990:BVB

[Hal90b] Peter Hall. On the bias of variable bandwidth curve estimators. Biometrika, 77(3):529–535, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336991.

Heitjan:1994:IGI

[Hei94] Daniel F. Heitjan. Ignorability in general incomplete-data models. Biometrika, 81(4):701–708, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337073.

Henderson:1990:PLR

[Hen90] Robin Henderson. A problem with the likelihood ratio test for a change-point hazard rate model. Biometrika, 77(4):835–843, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337106.

Hirst:1990:EIM

[HFC90] D. J. Hirst, I. Ford, and F. Critchley. An empirical investigation of methods for interval estimation of the log odds ratio in discriminant analysis. Biometrika, 77(3):609–615, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336999.

Hoover:1990:RLC

[HG90] Donald R. Hoover and Frank M. Guess. Response linked censoring: Modelling and estimation. Biometrika, 77(4):893–896, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337111.

Herrmann:1992:CBK

[HGK92] Eva Herrmann, Theo Gasser, and Alois Kneip. Choice of bandwidth for kernel regression when residuals are correlated. Biometrika, 79(4):783–795, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337234.

Holder:1993:MBC

[HH93] Daniel J. Holder and Francis Hsuan. Moment-based criteria for determining bioequivalence. Biometrika, 80(4):835–846, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336875.

Hall:1995:BRB

[HHJ95] Peter Hall, Joel L. Horowitz, and Bing-Yi Jing. On blocking rules for the bootstrap with dependent data. Biometrika, 82(3): 561–574, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337534.

Hsu:1994:CIA

[HHLR94] Jason C. Hsu, J. T. Gene Hwang, Hung-Kung Liu, and Stephen J. Ruberg. Confidence intervals associated with tests for bioequivalence. Biometrika, 81(1):103–114, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337054.

Hill:1990:GFM

[Hil90] Joe R. Hill. A general framework for model-based statistics. Biometrika, 77(1):115–126, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336054.

Hirotsu:1991:ACT

[Hir91] C. Hirotsu. An approach to comparing treatments based on repeated measures. Biometrika, 78(3):583–594, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337027.

Hardle:1996:SSV

[HK96] W. Härdle and A. Korostelev. Search for significant variables in nonparametric additive regression. Biometrika, 83(3):541–549, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337507.

Hirotsu:1992:MCP

[HKH92] Chihiro Hirotsu, Satoshi Kuriki, and Anthony J. Hayter. Multiple comparison procedures based on the maximal component of the cumulative chi-squared statistic. Biometrika, 79(2):381– 392, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336849.

Hall:1990:AOD

[HKT90] Peter Hall, J. W. Kay, and D. M. Titterington. Asymptotically optimal difference-based estimation of variance in nonparametric regression. Biometrika, 77(3):521–528, September 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336990.

Hanfelt:1995:ALR

[HL95] John J. Hanfelt and Kung-Yee Liang. Approximate likelihood ratios for general estimating functions. Biometrika, 82(3):461– 477, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337526.

Hu:1996:ETL

[HL96] X. Joan Hu and Jerald F. Lawless. Estimation from truncated lifetime data with supplementary information on covariates and censoring times. Biometrika, 83(4):747–761, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337280.

Hsu:1997:HBS

[HL97] John S. J. Hsu and Tom Leonard. Hierarchical Bayesian semiparametric procedures for logistic regression. Biometrika, 84(1): 85–93, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337557.

Huang:1998:NEJ

[HL98] Yijian Huang and Thomas A. Louis. Nonparametric estimation of the joint distribution of survival time and mark variables. Biometrika, 85(4):785–798, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337483.

Hu:1999:IAU

[HL99] X. Joan Hu and Stephen W. Lagakos. Interim analyses using repeated confidence bands. Biometrika, 86(3):517–529, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673651.

Hall:1990:VEN

[HM90] Peter Hall and J. S. Marron. On variance estimation in nonparametric regression. Biometrika, 77(2):415–419, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336824.

Henderson:1991:RPR

[HM91] Robin Henderson and Alvin Milner. On residual plots for relative risk regression. Biometrika, 78(3):631–636, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337032.

Heyde:1993:CQL

[HM93] C. C. Heyde and R. Morton. On constrained quasi-likelihood estimation. Biometrika, 80(4):755–761, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336867.

Heyde:1998:MRG

[HM98] C. C. Heyde and R. Morton. Multiple roots in general estimating equations. Biometrika, 85(4):954–959, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337496.

Hall:1995:PLS

[HMT95] Peter Hall, J. Stephen Marron, and D. M. Titterington. On partial local smoothing rules for curve estimation. Biometrika, 82(3):575–587, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337535.

Heathcote:1990:LSE

[HN90] C. R. Heathcote and D. F. Nicholls. Least-squares estimation of the contact rate in models for the spread of infectious diseases. Biometrika, 77(1):161–168, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336058.

Ha:1993:BEC

[HN93] Eunho Ha and H. Joseph Newton. The bias of estimators of causal spatial autoregressive processes. Biometrika, 80(1):242– 245, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336773.

Harvill:1995:SAD

[HN95] Jane L. Harvill and H. Joseph Newton. Saddlepoint approximations for the difference of order statistics. Biometrika, 82 (1):226–231, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337643.

Hong:1996:TIB

[Hon96] Yongmiao Hong. Testing for independence between two covariance stationary time series. Biometrika, 83(3):615–625, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337513.

Horn:1990:RQE

[Hor90] Paul S. Horn. Robust quantile estimators for skewed populations. Biometrika, 77(3):631–636, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337002.

Hsu:1996:GMH

[HP96a] Li Hsu and Ross L. Prentice. A generalisation of the Mantel– Haenszel test to bivariate failure time data. Biometrika, 83(4): 905–911, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337293.

Hsu:1996:ASD

[HP96b] Li Hsu and Ross L. Prentice. On assessing the strength of dependency between failure time variates. Biometrika, 83(3):491–506, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337503.

Hall:1998:NDT

[HPT98] Peter Hall, Byeong U. Park, and Berwin A. Turlach. A note on design transformation and binning in nonparametric curve estimation. Biometrika, 85(2):469–476, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337373.

Hall:1999:PIB

[HPT99] Peter Hall, Liang Peng, and Nader Tajvidi. On prediction intervals based on predictive likelihood or bootstrap methods. Biometrika, 86(4):871–880, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673591.

Hsu:1999:DEU

[HPZF99] Li Hsu, Ross L. Prentice, Lue Ping Zhao, and Juan Juan Fan. On dependence estimation using correlated failure time data from case-control family studies. Biometrika, 86(4):743–753, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673582.

Harvill:1999:NTN

[HR99] Jane L. Harvill and Bonnie K. Ray. A note on tests for nonlinearity in a vector time series. Biometrika, 86(3):728–734, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673668.

Hoover:1998:NSE

[HRWY98] Donald R. Hoover, John A. Rice, Colin O. Wu, and Li-Ping Yang. Nonparametric smoothing estimates of time-varying coefficient models with longitudinal data. Biometrika, 85(4):809– 822, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337485.

Heller:1990:CER

[HS90] Glenn Heller and Jeffrey S. Simonoff. A comparison of estimators for regression with a censored response variable. Biometrika, 77(3):515–520, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336989.

Hills:1993:DPI

[HS93] Susan E. Hills and Adrian F. M. Smith. Diagnostic plots for improved parameterization in Bayesian inference. Biometrika, 80(1):61–74, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336757.

He:1997:LRA

[HS97] Xuming He and Liji Shen. Linear regression after spline transformation. Biometrika, 84(2):474–481, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337472.

Hauschke:1996:NWA

[HSH96] Dieter Hauschke, Volker W. Steinijans, and Ludwig A. Hothorn. A note on Welch’s approximate t-solution to bioequivalence assessment. Biometrika, 83(1):236–237, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337446.

Hsieh:1996:TMT

[Hsi96] Fushing Hsieh. A transformation model for two survival curves: An empirical process approach. Biometrika, 83(3):519–528, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337505.

Hall:1991:ODB

[HSJM91] Peter Hall, Simon J. Sheather, M. C. Jones, and J. S. Marron. On optimal data-based bandwidth selection in kernel density estimation. Biometrika, 78(2):263–269, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337251.

Hurvich:1990:IEK

[HST90] Clifford M. Hurvich, Robert Shumway, and Chih-Ling Tsai. Improved estimators of Kullback–Leibler information for autoregressive model selection in small samples. Biometrika, 77(4): 709–719, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337094.

Hurvich:1991:BCA

[HT91] Clifford M. Hurvich and Chih-Ling Tsai. Bias of the corrected AIC criterion for underfitted regression and time series models. Biometrika, 78(3):499–509, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337019.

Hjellvik:1995:NTL

[HT95a] Vidar Hjellvik and Dag Tjøstheim. Nonparametric tests of linearity for time series. Biometrika, 82(2):351–368, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337413.

Hurvich:1995:RRC

[HT95b] Clifford M. Hurvich and Chih-Ling Tsai. Relative rates of convergence for efficient model selection criteria in linear regression. Biometrika, 82(2):418–425, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337419.

Hu:1996:ESD

[HT96] Ping Hu and Anastasios A. Tsiatis. Estimating the survival distribution when ascertainment of vital status is subject to delay. Biometrika, 83(2):371–380, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337607.

Hall:1997:ECI

[HT97] Peter Hall and Berwin A. Turlach. Enhancing convolution and interpolation methods for nonparametric regression. Biometrika, 84(4):779–790, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337651.

Hurvich:1998:CAH

[HT98] Clifford M. Hurvich and Chih-Ling Tsai. A crossvalidatory AIC for hard wavelet thresholding in spatially adaptive function estimation. Biometrika, 85(3):701–710, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337397.

Hjellvik:1999:MPI

[HT99] Vidar Hjellvik and Dag Tjøstheim. Modelling panels of intercorrelated autoregressive time series. Biometrika, 86(3):573–590, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673655.

Hu:1997:SCS

[Hu97] Inchi Hu. Strong consistency in stochastic regression models via posterior covariance matrices. Biometrika, 84(3):744–749, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337595.

Hu:1998:SDN

[Hu98] Inchi Hu. On sequential designs in nonlinear problems. Biometrika, 85(2):496–503, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337378.

Harrison:1991:DLM

[HW91] Jeff Harrison and Mike West. Dynamic linear model diagnostics. Biometrika, 78(4):797–808, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336931.

Hall:1993:PBC

[HW93] Peter Hall and Andrew Wood. On the performance of boxcounting estimators of fractal dimension. Biometrika, 80(1): 246–252, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336774.

Hwang:1997:VTI

[HW97] J. T. Gene Hwang and Weizhen Wang. The validity of the test of individual equivalence ratios. Biometrika, 84(4):893–900, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337659.

Hu:1995:BBE

[HZ95] Fei Fang Hu and James V. Zidek. A bootstrap based on the estimating equations of the linear model. Biometrika, 82(2):263– 275, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337406.

Hu:1997:PCT

[HZ97] Ping Hu and Marvin Zelen. Planning clinical trials to evaluate early detection programmes. Biometrika, 84(4):817–829, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337654.

Imbens:1995:DCD

[IR95] Guido W. Imbens and Donald B. Rubin. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82 (4):694–695, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337334. See [Pea95a].

Jones:1990:SCU

[JA90] Richard H. Jones and Lynn M. Ackerson. Serial correlation in unequally spaced longitudinal data. Biometrika, 77(4):721–731, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337095.

Jones:1992:NTM

[JC92] Michael P. Jones and John Crowley. Nonparametric tests of the Markov model for survival data. Biometrika, 79(3):513–522, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336782.

Jefferys:1990:REW

[Jef90] William H. Jefferys. Robust estimation when more than one variable per equation of condition has error. Biometrika, 77(3): 597–607, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336998.

Jensen:1992:MSL

[Jen92] J. L. Jensen. The modified signed likelihood statistic and saddlepoint approximations. Biometrika, 79(4):693–703, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337225.

Jandhyala:1999:CDB

[JF99] Venkata K. Jandhyala and Stergios B. Fotopoulos. Capturing the distributional behaviour of the maximum likelihood estimator of a changepoint. Biometrika, 86(1):129–140, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673542.

Jing:1994:BSA

[JFR94] B.-Y. Jing, A. Feuerverger, and J. Robinson. On the bootstrap saddlepoint approximations. Biometrika, 81(1):211–215, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337062.

Jensen:1991:GCA

[JJL91] Søren Tolver Jensen, Søren Johansen, and Steffen L. Lauritzen. Globally convergent algorithms for maximizing a likelihood function. Biometrika, 78(4):867–877, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336939.

Jorgensen:1999:SSM

[JLCSS99] Bent Jørgensen, Søren Lundbye-Christensen, Peter Xue-Kun Song, and Li Sun. A state space model for multivariate longitudinal count data. Biometrika, 86(1):169–181, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673545.

Jones:1995:SBR

[JLN95] M. C. Jones, O. Linton, and J. P. Nielsen. A simple bias reduction method for density estimation. Biometrika, 82(2):327– 338, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337411.

Jewell:1993:FCP

[JN93] Nicholas P. Jewell and Jens P. Nielsen. A framework for consistent prediction rules based on markers. Biometrika, 80(1):153– 164, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336765.

Jones:1991:KDE

[Jon91] M. C. Jones. Kernel density estimation for length biased data. Biometrika, 78(3):511–519, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337020.

Jones:1996:LDF

[Jon96] M. C. Jones. The local dependence function. Biometrika, 83(4): 899–904, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337292.

Jones:1997:CSR

[Jon97] Michael P. Jones. A class of semiparametric regressions for the accelerated failure time model. Biometrika, 84(1):73–84, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337556.

Jorgensen:1993:IFI

[Jor93] Murray A. Jorgensen. Influence functions for iteratively defined statistics. Biometrika, 80(2):253–265, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337197.

Jacroux:1990:CTF

[JR90] Michael Jacroux and Rita Saha Ray. On the construction of trend-free run orders of treatments. Biometrika, 77(1):187– 191, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336061.

John:1992:BEF

[JS92] J. A. John and Deborah J. Street. Bounds for the efficiency factor of row-column designs. Biometrika, 79(3):658–661, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336798.

See corrections [JS93].

John:1993:ACB

[JS93] J. A. John and Deborah J. Street. Amendments and corrections: “Bounds for the efficiency factor of row-column designs” [Biometrika 79 (1992), no. 3, 658–661; (94a:62118)]. Biometrika, 80(3):712–713, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337194. See [JS92].

Jones:1998:EFP

[JSAHM98] M. C. Jones, M. Samiuddin, A. H. Al-Harbey, and T. A. H. Maatouk. The edge frequency polygon. Biometrika, 85(1):235– 239, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337323.

Jennison:1991:ECS

[JT91] Christopher Jennison and Bruce W. Turnbull. Exact calculations for sequential t, χ^{2} and F tests. Biometrika, 78(1):133– 141, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336903.

Jung:1996:RAL

[Jun96] Sin-Ho Jung. Regression analysis for long-term survival rate. Biometrika, 83(1):227–232, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337444.

James:1994:REV

[JVDW94] A. T. James, W. N. Venables, I. B. Dry, and J. T. Wiskich. Random effects and variances as a synthesis of nonlinear regression analyses of mitochondrial electron transport. Biometrika, 81(2):219–235, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336953.

Kakizawa:1996:HOM

[Kak96] Yoshihide Kakizawa. Higher order monotone Bartlett-type adjustment for some multivariate test statistics. Biometrika, 83(4): 923–927, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337296.

Kass:1990:DTL

[Kas90] Robert E. Kass. Data-translated likelihood and Jeffreys’s rules. Biometrika, 77(1):107–114, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336053.

Kavalieris:1991:NEA

[Kav91] L. Kavalieris. A note on estimating autoregressive-moving average order. Biometrika, 78(4):920–922, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336945.

Kohn:1992:NSR

[KAW92] Robert Kohn, Craig F. Ansley, and Chi-Ming Wong. Nonparametric spline regression with autoregressive moving average errors. Biometrika, 79(2):335–346, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336844.

Kang:1990:AIM

[KB90] Gunseog Kang and Douglas M. Bates. Approximate inferences in multiresponse regression analysis. Biometrika, 77(2):321– 331, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336810.

Kano:1990:CSA

[KBB90] Yutaka Kano, Maia Berkane, and Peter M. Bentler. Covariance structure analysis with heterogeneous kurtosis parameters. Biometrika, 77(3):575–585, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336996.

Kuk:1992:MMC

[KC92] Anthony Y. C. Kuk and Chen-Hsin Chen. A mixture model combining logistic regression with proportional hazards regression. Biometrika, 79(3):531–541, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336784.

Kempthorne:1990:DPC

[Kem90] Oscar Kempthorne. Discussion of paper by C. B. Begg. Biometrika, 77(3):481–483, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336983. See [Beg90b].

Kosorok:1993:USF

[KF93] Michael R. Kosorok and Thomas R. Fleming. Using surrogate failure time data to increase cost effectiveness in clinical trials. Biometrika, 80(4):823–833, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336874.

Kim:1991:BST

[Kim91] Dongcheol Kim. A Bayesian significance test of the stationarity of regression parameters. Biometrika, 78(3):667–675, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337036.

Kim:1992:PSM

[Kim92] Seong-Ju Kim. A practical solution to the multivariate Behrens– Fisher problem. Biometrika, 79(1):171–176, March 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337157.

Konishi:1996:GIC

[KK96] Sadanori Konishi and Genshiro Kitagawa. Generalised information criteria in model selection. Biometrika, 83(4):875–890, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337290.

Kim:1998:TSG

[KL98] Jinheum Kim and Seung-Yeoun Lee. Two-sample goodnessof-fit tests for additive risk models with censored observations. Biometrika, 85(3):593–603, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337389.

Kleffe:1993:POS

[Kle93] J. Kleffe. Parent-offspring and sibling correlation estimation based on MINQUE theory. Biometrika, 80(2):393–404, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337209.

Khattree:1994:EIC

[KN94] Ravindra Khattree and Dayanand N. Naik. Estimation of interclass correlation under circular covariance. Biometrika, 81 (3):612–617, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337134.

Koenker:1994:QSS

[KNP94] Roger Koenker, Pin Ng, and Stephen Portnoy. Quantile smoothing splines. Biometrika, 81(4):673–680, December 1994. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337070.

Komaki:1993:SSM

[Kom93] Fumiyasu Komaki. State-space modelling of time series sampled from continuous processes with pulses. Biometrika, 80(2):417– 429, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337211.

Komaki:1996:APP

[Kom96] Fumiyasu Komaki. On asymptotic properties of predictive distributions. Biometrika, 83(2):299–313, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337602.

Koopman:1993:DSS

[Koo93] Siem Jan Koopman. Disturbance smoother for state space models. Biometrika, 80(1):117–126, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336762.

Korn:1993:EDF

[Kor93] Edward L. Korn. On estimating the distribution function for quality of life in cancer clinical trials. Biometrika, 80(3):535– 542, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337175.

Kosorok:1999:TSQ

[Kos99] Michael R. Kosorok. Two-sample quantile tests under general conditions. Biometrika, 86(4):909–921, December 1999. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673594. See correction

[Kos01].

Kosorok:2001:ACT

[Kos01] Michael R. Kosorok. Amendments and corrections: “Twosample quantile tests under general conditions” [Biometrika 86 (1999), no. 4, 909–921; 1741986]. Biometrika, 88(4):1201, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/88/4/1201. See [Kos99].

Konishi:1992:PCA

[KR92] Sadanori Konishi and C. Radhakrishna Rao. Principal component analysis for multivariate familial data. Biometrika, 79(3): 631–641, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336794.

Keen:1991:AVI

[KS91] K. J. Keen and M. S. Srivastava. The asymptotic variance of the interclass correlation coefficient. Biometrika, 78(1):225– 228, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336914.

Koopman:1992:EST

[KS92] Siem Jan Koopman and Neil Shephard. Exact score for time series models in state space form. Biometrika, 79(4):823–826, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337237.

Kwan:1996:FSD

[KS96] Andy C. C. Kwan and Ah-Boon Sim. On the finite-sample distribution of modified portmanteau tests for randomness of a Gaussian time series. Biometrika, 83(4):938–943, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337299.

Kong:1997:RCA

[KS97] Fan Hui Kong and Eric Slud. Robust covariate-adjusted logrank tests. Biometrika, 84(4):847–862, December 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337656.

Knowles:1991:CRS

[KSZ91] Mark Knowles, David Siegmund, and Heping Zhang. Confidence regions in semilinear regression. Biometrika, 78(1):15–31, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336892.

Kauermann:1999:MDU

[KT99] Göran Kauermann and Gerhard Tutz. On model diagnostics using varying coefficient models. Biometrika, 86(1):119– 128, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673541.

Kowalski:1997:RCE

[KTDMB97] Jeanne Kowalski, Xin M. Tu, Roger S. Day, and José R. Mendoza-Blanco. On the rate of convergence of the ECME algorithm for multiple regression models with t-distributed errors. Biometrika, 84(2):269–281, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337456.

Kuk:1990:ASQ

[Kuk90] Anthony Y. C. Kuk. Asking sensitive questions indirectly. Biometrika, 77(2):436–438, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336828.

Kuk:1993:KME

[Kuk93] Anthony Y. C. Kuk. A kernel method for estimating finite population distribution functions using auxiliary information. Biometrika, 80(2):385–392, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337208.

Kumazawa:1993:CNT

[Kum93] Yoshiki Kumazawa. Comments on ‘A New Test for Mean Residual Life Times’ by Ahmad. Biometrika, 80(2):473–474, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337218.

See [Ahm92].

Kunert:1991:CDT

[Kun91] Joachim Kunert. Cross-over designs for two treatments and correlated errors. Biometrika, 78(2):315–324, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337258.

Kuonen:1999:SAD

[Kuo99] D. Kuonen. Saddlepoint approximations for distributions of quadratic forms in normal variables. Biometrika, 86(4):929–935, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673596.

Kushner:1997:OET

[Kus97] H. B. Kushner. Optimality and efficiency of two-treatment repeated measurements designs. Biometrika, 84(2):455–468, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337470.

See corrections [Kus99].

Kushner:1999:ACO

[Kus99] H. B. Kushner. Amendments and corrections:“Optimality and efficiency of two-treatment repeated measurement designs” [Biometrika 84 (1997), no. 2, 455–468; (98k:62134)]. Biometrika, 86(1):234, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673554. See [Kus97].

Keaney:1994:IAB

[KW94] K. M. Keaney and L. J. Wei. Interim analyses based on median survival times. Biometrika, 81(2):279–286, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336958.

Kwan:1997:FRF

[KW97] Andy C. C. Kwan and Yangru Wu. Further results on the finite-sample distribution of Monti’s portmanteau test for the adequacy of an ARMA(p, q) model. Biometrika, 84(3):733–736, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337593.

Lam:1998:CTE

[Lam98] K. F. Lam. A class of tests for the equality of k cause-specific hazard rates in a competing risks model. Biometrika, 85(1):179– 188, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337318.

Lang:1992:OOI

[Lan92] Joseph B. Lang. Obtaining the observed information matrix for the Poisson log linear model with incomplete data. Biometrika, 79(2):405–407, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336852.

Lavine:1995:ALQ

[Lav95] M. Lavine. On an approximate likelihood for quantiles. Biometrika, 82(1):220–222, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337641.

Langholz:1995:CMS

[LB95] Bryan Langholz and Ornulf Borgan. Counter-matching: A stratified nested case-control sampling method. Biometrika, 82 (1):69–79, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337628.

Lewis:1991:GBR

[LD91] S. M. Lewis and A. M. Dean. On general balance in row-column designs. Biometrika, 78(3):595–600, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337028.

Lunn:1998:NGC

[LD98] A. D. Lunn and S. J. Davies. A note on generating correlated binary variables. Biometrika, 85(2):487–490, June 1998. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337376.

Lee:1990:MAS

[Lee90] Sik Yum Lee. Multilevel analysis of structural equation models. Biometrika, 77(4):763–772, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337099.

Lee:1997:SSS

[Lee97] Thomas C. M. Lee. A simple span selector for periodogram smoothing. Biometrika, 84(4):965–969, December 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337667.

Lefrancois:1991:DIO

[Lef91] Bernard Lefrançois. Detecting over-influential observations in time series. Biometrika, 78(1):91–99, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336899.

Lenk:1991:TPB

[Len91] Peter J. Lenk. Towards a practicable Bayesian nonparametric density estimator. Biometrika, 78(3):531–543, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337022.

Levin:1990:SCC

[Lev90] Bruce Levin. The saddlepoint correction in conditional logistic likelihood analysis. Biometrika, 77(2):275–285, June 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336805.

Lin:1994:CBS

[LFW94] D. Y. Lin, T. R. Fleming, and L. J. Wei. Confidence bands for survival curves under the proportional hazards model. Biometrika, 81(1):73–81, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337051.

Linton:1996:EAR

[LH96] O. B. Linton and W. Härdle. Estimation of additive regression models with known links. Biometrika, 83(3):529–540, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337506.

Liu:1999:UEF

[LH99] Aiyi Liu and W. J. Hall. Unbiased estimation following a group sequential test. Biometrika, 86(1):71–78, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673537.

Li:1991:SLM

[Li91a] W. K. Li. Some Lagrange multiplier tests for seasonal differencing. Biometrika, 78(2):381–387, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337265.

Li:1991:TMA

[Li91b] W. K. Li. Testing model adequacy for some Markov regression models for time series. Biometrika, 78(1):83–89, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336898.

Li:1992:ASE

[Li92] W. K. Li. On the asymptotic standard errors of residual autocorrelations in nonlinear time series modelling. Biometrika, 79 (2):435–437, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336859.

Li:1993:DFQ

[Li93] Bing Li. A deviance function for the quasi-likelihood method. Biometrika, 80(4):741–753, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336866.

Li:1995:PMI

[Li95] Hongyi Li. The power of modified instrumental variables tests for unit roots. Biometrika, 82(3):660–666, September 1995. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337544.

Laud:1996:PSP

[LI96a] Purushottam W. Laud and Joseph G. Ibrahim. Predictive specification of prior model probabilities in variable selection. Biometrika, 83(2):267–274, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337599.

Lipsitz:1996:CMI

[LI96b] Stuart R. Lipsitz and Joseph G. Ibrahim. A conditional model for incomplete covariates in parametric regression models. Biometrika, 83(4):916–922, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337295.

Li:1998:OEE

[Li98] Bing Li. An optimal estimating equation based on the first three cumulants. Biometrika, 85(1):103–114, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337312.

Lieberman:1994:LAM

[Lie94a] Offer Lieberman. A Laplace approximation to the moments of a ratio of quadratic forms. Biometrika, 81(4):681–690, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337071.

Lieberman:1994:SAL

[Lie94b] Offer Lieberman. Saddlepoint approximation for the least squares estimator in first-order autoregression. Biometrika, 81 (4):807–811, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2337086. See corrections [Lie96].

Lieberman:1996:ACS

[Lie96] Offer Lieberman. Amendments and corrections: “Saddlepoint approximation for the least squares estimator in firstorder autoregression” [Biometrika 81 (1994), no. 4, 807–811; (96f:62150)]. Biometrika, 83(1):247, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337450. See [Lie94b].

Lieberman:1998:ULE

[Lie98] Offer Lieberman. From unbiased linear estimating equations to unbiased estimators. Biometrika, 85(1):244–250, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337325.

Lin:1991:NST

[Lin91] D. Y. Lin. Nonparametric sequential testing in clinical trials with incomplete multivariate observations. Biometrika, 78(1): 123–131, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336902.

Lin:1992:SLR

[Lin92] D. Y. Lin. Sequential log rank tests adjusting for covariates with the accelerated life model. Biometrika, 79(3):523–529, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336783.

Lin:1997:VCT

[Lin97a] Xihong Lin. Variance component testing in generalised linear models with random effects. Biometrika, 84(2):309–326, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337459.

Linton:1997:EEA

[Lin97b] Oliver B. Linton. Efficient estimation of additive nonparametric regression models. Biometrika, 84(2):469–473, June 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337471.

Liseo:1993:ENP

[Lis93] Brunero Liseo. Elimination of nuisance parameters with reference priors. Biometrika, 80(2):295–304, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337200.

Little:1994:CPM

[Lit94] Roderick J. A. Little. A class of pattern-mixture models for normal incomplete data. Biometrika, 81(3):471–483, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337120.

Liu:1996:PTM

[Liu96a] Jun S. Liu. Peskun’s theorem and a modified discrete-state Gibbs sampler. Biometrika, 83(3):681–682, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337518.

Liu:1996:MSM

[Liu96b] Shu-Ing Liu. Model selection for multiperiod forecasts. Biometrika, 83(4):861–873, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337289.

Liu:1996:CUC

[Liu96c] Wei Liu. The construction of upper confidence bounds on the range of several location parameters following a sequential range test. Biometrika, 83(1):223–226, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337443.

Liu:1998:IMC

[Liu98] Chuanhai Liu. Information matrix computation from conditional information via normal approximation. Biometrika, 85 (4):973–979, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337499.

Levin:1990:BBC

[LK90] Bruce Levin and Fanhui Kong. Bartlett’s bias correction to the profile score function is a saddlepoint correction. Biometrika, 77 (1):219–221, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336067.

Lee:1996:RST

[LKT96] Sandra J. Lee, Kyungmann Kim, and Anastasios A. Tsiatis. Repeated significance testing in longitudinal clinical trials. Biometrika, 83(4):779–789, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337282.

Lin:1992:NST

[LL92] D. Y. Lin and P. Y. Liu. Nonparametric sequential tests against ordered alternatives in multiple-armed clinical trials. Biometrika, 79(2):420–425, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336856.

Lipsitz:1990:FDM

[LLH90] Stuart R. Lipsitz, Nan M. Laird, and David P. Harrington. Finding the design matrix for the marginal homogeneity model. Biometrika, 77(2):353–358, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336813.

Lipsitz:1991:GEE

[LLH91] Stuart R. Lipsitz, Nan M. Laird, and David P. Harrington. Generalized estimating equations for correlated binary data: Using the odds ratio as a measure of association. Biometrika, 78 (1):153–160, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336905.

Lloyd:1992:TCH

[Llo92a] Chris J. Lloyd. Testing capture homogeneity in a recapture model. Biometrika, 79(3):555–561, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336786.

Lloyd:1992:MME

[Llo92b] Christopher J. Lloyd. Modified martingale estimation for recapture experiments with heterogeneous capture probabilities. Biometrika, 79(4):833–836, December 1992. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337239.

Lloyd:1994:EMM

[Llo94] C. J. Lloyd. Efficiency of martingale methods in recapture studies. Biometrika, 81(2):305–315, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336961.

Lloyd:1996:APR

[Llo96a] C. J. Lloyd. Accurate pivotals from recapture experiments with time variation. Biometrika, 83(3):700–708, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337522.

Lloyd:1996:EAR

[Llo96b] C. J. Lloyd. Estimating attributable response as a function of a continuous risk factor. Biometrika, 83(3):563–573, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337509.

Lazar:1998:EPC

[LM98] Nicole Lazar and Per Aslak Mykland. An evaluation of the power and conditionality properties of empirical likelihood. Biometrika, 85(3):523–534, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337383.

Lazar:1999:ELP

[LM99] Nicole A. Lazar and Per Aslak Mykland. Empirical likelihood in the presence of nuisance parameters. Biometrika, 86(1):203– 211, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673548.

Linton:1995:KME

[LN95] Oliver Linton and Jens Perch Nielsen. A kernel method of estimating structured nonparametric regression based on marginal integration. Biometrika, 82(1):93–100, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337630.

Loader:1991:IHR

[Loa91] Clive R. Loader. Inference for a hazard rate change point. Biometrika, 78(4):749–757, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336926.

Lohr:1995:OBD

[Loh95] Sharon L. Lohr. Optimal Bayesian design of experiments for the one-way random effects model. Biometrika, 82(1):175– 186, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337637.

Lin:1998:AHR

[LOY98] D. Y. Lin, David Oakes, and Zhiliang Ying. Additive hazards regression with current status data. Biometrika, 85(2):289– 298, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337358.

Liu:1993:HMH

[LP93] Qing Liu and Donald A. Pierce. Heterogeneity in Mantel– Haenszel-type models. Biometrika, 80(3):543–556, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337176.

Liu:1994:NGH

[LP94] Qing Liu and Donald A. Pierce. A note on Gauss–Hermite quadrature. Biometrika, 81(3):624–629, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337136.

Liu:1994:EAS

[LR94] Chuanhai Liu and Donald B. Rubin. The ECME algorithm: A simple extension of EM and ECM with faster monotone convergence. Biometrika, 81(4):633–648, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337067.

Liu:1998:ESE

[LR98] Chuanhai Liu and Donald B. Rubin. Ellipsoidally symmetric extensions of the general location model for mixed categorical and continuous data. Biometrika, 85(3):673–688, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337395.

Lee:1993:BDA

[LRW93] Chu-In Charles Lee, Tim Robertson, and F. T. Wright. Bounds on distributions arising in order restricted inferences with restricted weights. Biometrika, 80(2):405–416, June 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337210.

Lin:1996:CTF

[LRW96] D. Y. Lin, J. M. Robins, and L. J. Wei. Comparing two failure time distributions in the presence of dependent censoring. Biometrika, 83(2):381–393, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337608.

Liu:1998:PEA

[LRW98] Chuanhai Liu, Donald B. Rubin, and Ying Nian Wu. Parameter expansion to accelerate EM: the PX–EM algorithm. Biometrika, 85(4):755–770, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337481.

Lin:1999:NEG

[LSY99] D. Y. Lin, W. Sun, and Zhiliang Ying. Nonparametric estimation of the gap time distributions for serial events with censored data. Biometrika, 86(1):59–70, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673536.

Liang:1992:EBC

[LT92] Kung-Yee Liang and Daniel Tsou. Empirical Bayes and conditional inference with many nuisance parameters. Biometrika, 79(2):261–270, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336837.

Ledford:1996:SNI

[LT96] Anthony W. Ledford and Jonathan A. Tawn. Statistics for near independence in multivariate extreme values. Biometrika, 83 (1):169–187, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337440.

Luo:1997:LRT

[LTC97] Xiaolong Luo, Bruce W. Turnbull, and Larry C. Clark. Likelihood ratio tests for a changepoint with survival data. Biometrika, 84(3):555–565, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337578.

Langholz:1990:TEM

[LTCR90] Bryan Langholz, Duncan Thomas, Tsunjen Chen, and Phillip Rhodes. Tests of effect in 1: J matched case-control studies based on a modified likelihood approach. Biometrika, 77(4): 897–900, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337112.

Li:1999:SIQ

[LTW99] Gang Li, Ram C. Tiwari, and Martin T. Wells. Semiparametric inference for a quantile comparison function with applications to receiver operating characteristic curves. Biometrika, 86(3): 487–502, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673649.

Lucas:1994:PLA

[Luc94a] Thomas W. Lucas. Priors and likelihoods that are arbitrarily sensitive to outlying observations. Biometrika, 81(2):385– 389, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336968.

Luceno:1994:FAE

[Luc94b] Alberto Luceño. A fast algorithm for the exact likelihood of stationary and partially nonstationary vector autoregressivemoving average processes. Biometrika, 81(3):555–565, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337126.

Luceno:1996:FLA

[Luc96] Alberto Luceño. A fast likelihood approximation for vector general linear processes with long series: Application to fractional differencing. Biometrika, 83(3):603–614, September 1996. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337512.

Luceno:1997:EMV

[Luc97] Alberto Luceno. Estimation of missing values in possibly partially nonstationary vector time series. Biometrika, 84(2):495– 499, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337476.

Liu:1994:CSG

[LWK94] Jun S. Liu, Wing Hung Wong, and Augustine Kong. Covariance structure of the Gibbs sampler with applications to the comparisons of estimators and augmentation schemes. Biometrika, 81 (1):27–40, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337047.

Lin:1993:CCM

[LWY93] D. Y. Lin, L. J. Wei, and Z. Ying. Checking the Cox model with cumulative sums of martingale-based residuals. Biometrika, 80 (3):557–572, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337177.

Lin:1998:AFT

[LWY98] D. Y. Lin, L. J. Wei, and Zhiliang Ying. Accelerated failure time models for counting processes. Biometrika, 85(3):605–618, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337390.

Lin:1993:SNE

[LY93] D. Y. Lin and Zhiliang Ying. A simple nonparametric estimator of the bivariate survival function under univariate censoring. Biometrika, 80(3):573–581, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337178.

Lin:1994:SAA

[LY94] D. Y. Lin and Zhiliang Ying. Semiparametric analysis of the additive risk model. Biometrika, 81(1):61–71, March 1994. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337050.

Lee:1999:NLR

[LY99] Stephen M. S. Lee and G. Alastair Young. Nonparametric likelihood ratio confidence intervals. Biometrika, 86(1):107– 118, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673540.

Ma:1997:ELF

[Ma97] Chunsheng Ma. On the exact likelihood function of a multivariate autoregressive moving average model. Biometrika, 84(4): 957–964, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337666.

Malani:1995:MRR

[Mal95] Hina Mehta Malani. A modification of the redistribution to the right algorithm using disease markers. Biometrika, 82(3): 515–526, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337530.

Meshkani:1992:EBE

[MB92a] M. R. Meshkani and L. Billard. Empirical Bayes estimators for a finite Markov chain. Biometrika, 79(1):185–193, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337159.

Monahan:1992:PLB

[MB92b] John F. Monahan and Dennis D. Boos. Proper likelihoods for Bayesian analysis. Biometrika, 79(2):271–278, June 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336838.

Monod:1993:TFD

[MB93] H. Monod and R. A. Bailey. Two-factor designs balanced for the neighbour effect of one factor. Biometrika, 80(3):643–659, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337184.

Mukerjee:1991:BTA

[MC91] Rahul Mukerjee and Tapas K. Chandra. Bartlett-type adjustment for the conditional likelihood ratio statistic of Cox and Reid. Biometrika, 78(2):365–372, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337263.

McCullagh:1992:CIC

[McC92] Peter McCullagh. Conditional inference and Cauchy models. Biometrika, 79(2):247–259, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336836.

McCullagh:1994:EMQ

[McC94] Peter McCullagh. Exponential mixtures and quadratic exponential families. Biometrika, 81(4):721–729, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337075. See

corrections [McC96].

McCullagh:1996:ACE

[McC96] Peter McCullagh. Amendments and corrections: “Exponential mixtures and quadratic exponential families” [Biometrika 81 (1994), no. 4, 721–729; (96d:62014)]. Biometrika, 83(1): 248, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337451. See [McC94].

Mukerjee:1993:FVP

[MD93] Rahul Mukerjee and Dipak K. Dey. Frequentist validity of posterior quantiles in the presence of a nuisance parameter: Higher order asymptotics. Biometrika, 80(3):499–505, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337171.

McClean:1995:NML

[MD95a] Sally McClean and Colum Devine. A nonparametric maximum likelihood estimator for incomplete renewal data. Biometrika, 82(4):791–803, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337345.

McCloud:1995:ACR

[MD95b] P. I. McCloud and J. N. Darroch. An analysis of categorical repeated measurements in the presence of an external factor. Biometrika, 82(1):47–55, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337626.

Martin:1992:RFC

[ME92] R. J. Martin and J. A. Eccleston. Recursive formulae for constructing block designs with dependent errors. Biometrika, 79 (2):426–430, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336857.

Martin:1998:VBC

[ME98] R. J. Martin and J. A. Eccleston. Variance-balanced changeover designs for dependent observations. Biometrika, 85(4): 883–892, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337490.

Muller:1996:BCF

[MEW96] Peter Müller, Alaattin Erkanli, and Mike West. Bayesian curve fitting using multivariate normal mixtures. Biometrika, 83(1): 67–79, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337433.

Mallick:1994:GLM

[MG94] Bani K. Mallick and Alan E. Gelfand. Generalized linear models with unknown link functions. Biometrika, 81(2):237–245, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336954.

Mukerjee:1997:SOP

[MG97] Rahul Mukerjee and Malay Ghosh. Second-order probability matching priors. Biometrika, 84(4):970–975, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337668.

Mijares:1990:NAB

[Mij90a] Tito A. Mijares. The normal approximation to the Bartlett– Nanda–Pillai trace test in multivariate analysis. Biometrika, 77 (1):230–233, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336070.

Mijares:1990:NAL

[Mij90b] Tito A. Mijares. On the normal approximation to the Lawley– Hotelling trace criterion. Biometrika, 77(2):443, June 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336830.

Mitchell:1993:TLV

[Mit93] Heather Mitchell. Testing for linearity versus convexity in regression using a contrast-type test. Biometrika, 80(3):695–697, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337189.

Mardia:1991:RST

[MK91] K. V. Mardia and J. T. Kent. Rao score tests for goodness of fit and independence. Biometrika, 78(2):355–363, June 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337262.

Mardia:1996:KSD

[MKGL96] K. V. Mardia, J. T. Kent, C. R. Goodall, and J. A. Little. Kriging and splines with derivative information. Biometrika, 83 (1):207–221, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2337442. See corrections [MKGL98].

Mardia:1998:ACK

[MKGL98] K. V. Mardia, J. T. Kent, C. R. Goodall, and J. A. Little. Amendments and corrections: “Kriging and splines with derivative information” [Biometrika 83 (1996), no. 1, 207– 221; (98d:65017)]. Biometrika, 85(2):505, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337379. See [MKGL96].

Molenberghs:1997:ALO

[MKL97] G. Molenberghs, M. G. Kenward, and E. Lesaffre. The analysis of longitudinal ordinal data with nonrandom drop-out. Biometrika, 84(1):33–44, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337553.

Murphy:1995:PPL

[ML95] Susan Murphy and Bing Li. Projected partial likelihood and its application to longitudinal data. Biometrika, 82(2):399– 406, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337416.

Madigan:1990:ERA

[MM90] David Madigan and Krzysztof Mosurski. An extension of the results of Asmussen and Edwards on collapsibility in contingency tables. Biometrika, 77(2):315–319, June 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336809. See correc-

tions [MM99].

Marzec:1991:TED

[MM91] Leszek Marzec and Paweł Marzec. On testing the equality in dispersion of two probability distributions. Biometrika, 78(4): 923–925, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336946.

Mammen:1997:MRK

[MM97a] E. Mammen and J. S. Marron. Mass recentred kernel smoothers. Biometrika, 84(4):765–777, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337650.

Marzec:1997:FCR

[MM97b] Leszek Marzec and Paweł Marzec. On fitting Cox’s regression model with time-dependent coefficients. Biometrika, 84(4): 901–908, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337660.

Madigan:1999:ACE

[MM99] David Madigan and Krzysztof Mosurski. Amendments and corrections:“An extension of the results of Asmussen and Edwards on collapsibility in contingency tables” [Biometrika 77 (1990), no. 2, 315–319; (91i:62075)]. Biometrika, 86(4):973, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673602.

See [MM90].

Mentre:1997:ODR

[MMB97] France Mentré, Alain Mallet, and Doha Baccar. Optimal design in random-effects regression models. Biometrika, 84(2):429– 442, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337468.

Moreau:1992:WLR

[MMLH92] Thierry Moreau, Jean Maccario, Joseph Lellouch, and Catherine Huber. Weighted log rank statistics for comparing two distributions. Biometrika, 79(1):195–198, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337160.

Mudholkar:1992:TVN

[MMS92] Govind S. Mudholkar, Maria McDermott, and Deo Kumar Srivastava. A test of p-variate normality. Biometrika, 79(4): 850–854, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337243.

Nam:1993:BCM

[mN93a] Jun mo Nam. Bias-corrected maximum likelihood estimator of a log common odds ratio. Biometrika, 80(3):688–694, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337188.

Morel:1993:FMD

[MN93b] Jorge G. Morel and Neerchal K. Nagaraj. A finite mixture distribution for modelling multinomial extra variation. Biometrika, 80(2):363–371, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337206.

Marshall:1997:NMA

[MO97] Albert W. Marshall and Ingram Olkin. A new method for adding a parameter to a family of distributions with application to the exponential and Weibull families. Biometrika, 84(3): 641–652, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2337585. See correction [MO05].

Marshall:2005:ACN

[MO05] Albert W. Marshall and Ingram Olkin. Amendments and corrections: “A new method for adding a parameter to a family of distributions with application to the exponential and Weibull families” [Biometrika 84 (1997), no. 3, 641–652; MR1603936]. Biometrika, 92(2):505, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://biomet.oxfordjournals.org/content/92/2/505.1;
 http://www.jstor.org/stable/20441205. See [MO97].

Monti:1994:PRA

[Mon94] Anna Clara Monti. A proposal for a residual autocorrelation test in linear models. Biometrika, 81(4):776–780, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337080.

Monti:1997:ELC

[Mon97] Anna Clara Monti. Empirical likelihood confidence regions in time series models. Biometrika, 84(2):395–405, June 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337465.

Morton:1991:AEM

[Mor91] Richard Morton. Analysis of extra-multinomial data derived from extra-Poisson variables conditional on their total. Biometrika, 78(1):1–6, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336890.

Morgenthaler:1992:LAD

[Mor92] Stephan Morgenthaler. Least-absolute-deviations fits for generalized linear models. Biometrika, 79(4):747–754, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337230.

Morton:1993:URC

[Mor93] R. Morton. Ultrastructural relationships and canonical variates analysis. Biometrika, 80(4):915–918, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336884.

McCulloch:1992:BFN

[MR92a] Robert E. McCulloch and Peter E. Rossi. Bayes factors for nonlinear hypotheses and likelihood distributions. Biometrika, 79(4):663–676, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337222.

Meng:1992:PLR

[MR92b] Xiao-Li Meng and Donald B. Rubin. Performing likelihood ratio tests with multiply-imputed data sets. Biometrika, 79(1):103– 111, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337151.

Meng:1993:MLE

[MR93a] Xiao-Li Meng and Donald B. Rubin. Maximum likelihood estimation via the ECM algorithm: A general framework. Biometrika, 80(2):267–278, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337198.

Monti:1993:RBE

[MR93b] Anna Clara Monti and Elvezio Ronchetti. On the relationship between empirical likelihood and empirical saddlepoint approximation for multivariate M-estimators. Biometrika, 80(2):329– 338, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337203.

Muller:1997:BSM

[MR97] Peter Müller and Kathryn Roeder. A Bayesian semiparametric model for case-control studies with errors in variables. Biometrika, 84(3):523–537, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337576.

Mukerjee:1999:PPM

[MR99] Rahul Mukerjee and N. Reid. On a property of probability matching priors: Matching the alternative coverage probabilities. Biometrika, 86(2):333–340, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673515.

Moreno-Rebollo:1999:IDS

[MRMRMP99] J. L. Moreno-Rebollo, A. Muñoz-Reyes, and J. Muñoz-Pichardo. Influence diagnostic in survey sampling: Conditional bias. Biometrika, 86(4):923–928, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673595.

Mangat:1990:ARR

[MS90a] N. S. Mangat and Ravindra Singh. An alternative randomized response procedure. Biometrika, 77(2):439–442, June 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336829.

Marx:1990:PCE

[MS90b] Brian D. Marx and Eric P. Smith. Principal component estimation for generalized linear regression. Biometrika, 77(1):23– 31, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336046.

Mukerjee:1990:OEF

[MS90c] Rahul Mukerjee and S. Sengupta. Optimal estimation of a finite population mean in the presence of linear trend. Biometrika, 77 (3):625–630, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337001.

Malec:1992:BMC

[MS92a] Donald Malec and J. Sedransk. Bayesian methodology for combining the results from different experiments when the specifications for pooling are uncertain. Biometrika, 79(3):593–601, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336790.

McLeish:1992:PLF

[MS92b] D. L. McLeish and Christopher G. Small. A projected likelihood function for semiparametric models. Biometrika, 79(1):93– 102, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337150.

McKeague:1994:PPA

[MS94] Ian W. McKeague and Peter D. Sasieni. A partly parametric additive risk model. Biometrika, 81(3):501–514, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337122.

M:1997:ACS

[MS97] Mohammad Salehi M. and George A. F. Seber. Adaptive cluster sampling with networks selected without replacement. Biometrika, 84(1):209–219, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337567.

Martinussen:1999:SAR

[MS99] Torben Martinussen and Thomas H. Scheike. A semiparametric additive regression model for longitudinal data. Biometrika, 86 (3):691–702, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673663.

Mukerjee:1997:OPD

[Muk97] Rahul Mukerjee. Optimal partial diallel crosses. Biometrika, 84 (4):939–948, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337664.

Muller:1991:SOK

[Mül91] Hans-Georg Müller. Smooth optimum kernel estimators near endpoints. Biometrika, 78(3):521–530, September 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337021.

Meng:1999:SED

[MV99] Xiao-Li Meng and David A. Van Dyk. Seeking efficient data augmentation schemes via conditional and marginal augmentation. Biometrika, 86(2):301–320, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673513.

Muller:1990:NAC

[MW90] Hans-Georg Müller and Jane-Ling Wang. Nonparametric analysis of changes in hazard rates for censored survival data: An alternative to change-point models. Biometrika, 77(2):305– 314, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336808.

McKnight:1992:TIR

[MW92] Barbara McKnight and Jürgen Wahrendorf. Tumour incidence rate alternatives and the cause-of-death test for carcinogenicity. Biometrika, 79(1):131–138, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337153.

Mardia:1994:SAP

[MW94] K. V. Mardia and A. N. Walder. Shape analysis of paired landmark data. Biometrika, 81(1):185–196, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337060.

Mallick:1997:CIS

[MW97] Bani K. Mallick and Stephen G. Walker. Combining information from several experiments with nonparameter priors. Biometrika, 84(3):697–706, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337589.

Muller:1997:LHR

[MWC97] Hans-Georg Müller, Jane-Ling Wang, and William B. Capra. From lifetables to hazard rates: The transformation approach. Biometrika, 84(4):881–892, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337658.

Madigan:1997:BME

[MY97] David Madigan and Jeremy C. York. Bayesian methods for estimation of the size of a closed population. Biometrika, 84(1): 19–31, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337552.

Maller:1992:EPI

[MZ92] R. A. Maller and S. Zhou. Estimating the proportion of immunes in a censored sample. Biometrika, 79(4):731–739, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337228.

Newbold:1993:BSA

[NA93] Paul Newbold and Christos Agiakloglou. Bias in the sample autocorrelations of fractional noise. Biometrika, 80(3):698–702, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337190.

Nagelkerke:1991:NGD

[Nag91] N. J. D. Nagelkerke. A note on a general definition of the coefficient of determination. Biometrika, 78(3):691–692, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337038.

Nakamura:1990:CSF

[Nak90] Tsuyoshi Nakamura. Corrected score function for errors-in-variables models: Methodology and application to generalized linear models. Biometrika, 77(1):127–137, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336055.

Nayak:1991:ENC

[Nay91] Tapan K. Nayak. Estimating the number of component processes of a superimposed process. Biometrika, 78(1):75–81, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336897.

Neuhaus:1999:BEL

[Neu99] John M. Neuhaus. Bias and efficiency loss due to misclassified responses in binary regression. Biometrika, 86(4):843–855, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673589.

Newton:1996:BPL

[New96] Michael A. Newton. Bootstrapping phylogenies: Large deviations and dispersion effects. Biometrika, 83(2):315–328, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337603.

Neuenschwander:1995:CCV

[NF95] Beat E. Neuenschwander and Bernard D. Flury. Common canonical variates. Biometrika, 82(3):553–560, September 1995.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337533.

Neuhaus:1992:EMD

[NHK92] J. M. Neuhaus, W. W. Hauck, and J. D. Kalbfleisch. The effects of mixture distribution misspecification when fitting mixedeffects logistic models. Biometrika, 79(4):755–762, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337231.

Nielsen:1997:BCU

[Nie97] B. Nielsen. Bartlett correction of the unit root test in autoregressive models. Biometrika, 84(2):500–504, June 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337477.

Nielsen:1999:LRT

[Nie99a] B. Nielsen. The likelihood-ratio test for rank in bivariate canonical correlation analysis. Biometrika, 86(2):279–288, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673511.

Nielsen:1999:SEH

[Nie99b] Jens Perch Nielsen. Super-efficient hazard estimation based on high-quality marker information. Biometrika, 86(1):227– 232, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673551.

Neuhaus:1993:GAA

[NJ93] John M. Neuhaus and Nicholas P. Jewell. A geometric approach to assess bias due to omitted covariates in generalized linear models. Biometrika, 80(4):807–815, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336872.

Neuhaus:1996:EEB

[NL96] J. M. Neuhaus and M. L. Lesperance. Estimation efficiency in a binary mixed-effects model setting. Biometrika, 83(2):441– 446, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337613.

Nadeau:1998:IMC

[NL98] C. Nadeau and J. F. Lawless. Inference for means and covariances of point processes through estimating functions. Biometrika, 85(4):893–906, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337491.

Natarajan:1995:NEP

[NM95] Ranjini Natarajan and Charles E. McCulloch. A note on the existence of the posterior distribution for a class of mixed models for binomial responses. Biometrika, 82(3):639–643, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337540.

Nott:1999:PLM

[NR99] David J. Nott and Tobias Rydén. Pairwise likelihood methods for inference in image models. Biometrika, 86(3):661–676, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673661.

Newton:1999:RAN

[NZ99] Michael A. Newton and Yunlei Zhang. A recursive algorithm for nonparametric analysis with missing data. Biometrika, 86 (1):15–26, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673533.

OBrien:1992:NQD

[O’B92] Timothy E. O’Brien. A note on quadratic designs for nonlinear regression models. Biometrika, 79(4):847–849, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337242.

Oakes:1994:SIM

[OC94] David Oakes and Lu Cui. On semiparametric inference for modulated renewal processes. Biometrika, 81(1):83–90, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337052.

Oakes:1990:NRL

[OD90] David Oakes and Tamraparni Dasu. A note on residual life. Biometrika, 77(2):409–410, June 1990. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336822.

Odaki:1993:IFD

[Oda93] Mitsuhiro Odaki. On the invertibility of fractionally differenced ARIMA processes. Biometrika, 80(3):703–709, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337191.

Ozturk:1997:GWC

[ÖH97] Ömer Öztürk and Thomas P. Hettmansperger. Generalised weighted Cramér–von Mises distance estimators. Biometrika, 84(2):283–294, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337457.

Omori:1993:IRE

[OJ93] Yasuhiro Omori and Richard A. Johnson. The influence of random effects on the unconditional hazard rate and survival functions. Biometrika, 80(4):910–914, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336883.

Ogata:1991:MLE

[OK91] Yosihiko Ogata and Koiti Katsura. Maximum likelihood estimates of the fractal dimension for random spatial patterns. Biometrika, 78(3):463–474, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337016.

Oakes:1992:FIB

[OM92] David Oakes and Amita K. Manatunga. Fisher information for a bivariate extreme value distribution. Biometrika, 79(4): 827–832, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337238.

Oman:1991:MEM

[Oma91] Samuel D. Oman. Multiplicative effects in mixed model analysis of variance. Biometrika, 78(4):729–739, December 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336924.

Oman:1998:CRS

[Oma98] Samuel D. Oman. Calibration with random slopes. Biometrika, 85(2):439–449, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337369.

ONeill:1992:ERN

[O’N92] Terence J. O’Neill. Error rates of non-Bayes classification rules and the robustness of Fisher’s linear discriminant function. Biometrika, 79(1):177–184, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337158.

OQuigley:1998:SRC

[OR98] John O’Quigley and Ethan Reiner. A stopping rule for the continual reassessment method. Biometrika, 85(3):741–748, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337400.

Orme:1998:IST

[Orm98] Chris D. Orme. On the insensitivity of the score test for heterogeneity to omitted covariates in multivariate failure time models. Biometrika, 85(2):457–461, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337371.

Osiewalski:1993:RBI

[OS93] Jacek Osiewalski and Mark F. J. Steel. Robust Bayesian inference in l_{q}-spherical models. Biometrika, 80(2):456–460, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337215.

Oh:1991:AML

[OSS91] Chang-Hyuck Oh, Norman C. Severo, and John Slivka. Approximations to the maximum likelihood estimate in some pure birth processes. Biometrika, 78(2):295–299, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337255.

Paik:1996:QLR

[Pai96] Myunghee Cho Paik. Quasi-likelihood regression models with missing covariates. Biometrika, 83(4):825–834, December 1996.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337286.

Pankratz:1993:DTO

[Pan93] Alan Pankratz. Detecting and treating outliers in dynamic regression models. Biometrika, 80(4):847–854, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336876.

Park:1997:NRB

[Par97a] Hyo-Il Park. A note on the relation between two forms of linear rank statistics for right censored and grouped data. Biometrika, 84(4):987–988, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337671.

Parmigiani:1997:TME

[Par97b] Giovanni Parmigiani. Timing medical examinations via intensity functions. Biometrika, 84(4):803–816, December 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337653.

Patel:1991:AID

[Pat91] H. I. Patel. Analysis of incomplete data from a clinical trial with repeated measurements. Biometrika, 78(3):609–619, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337030.

Paul:1990:MLE

[Pau90] S. R. Paul. Maximum likelihood estimation of intraclass correlation in the analysis of familial data: Estimating equation approach. Biometrika, 77(3):549–555, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336993.

Pauler:1998:SCR

[Pau98] Donna K. Pauler. The Schwarz criterion and related methods for normal linear models. Biometrika, 85(1):13–27, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337306.

Pawitan:1996:AEC

[Paw96] Yudi Pawitan. Automatic estimation of the cross-spectrum of a bivariate time series. Biometrika, 83(2):419–432, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337611.

Prentice:1992:CSF

[PC92] R. L. Prentice and J. Cai. Covariance and survivor function estimation using censored multivariate failure time data. Biometrika, 79(3):495–512, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336781. See correction

[PC93].

Prentice:1993:ACC

[PC93] R. L. Prentice and J. Cai. Amendments and corrections: “Covariance and survivor function estimation using censored multivariate failure time data” [Biometrika 79 (1992), no. 3, 495–512; (93j:62239)]. Biometrika, 80(3):711–712, September 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337193. See [PC92].

Philipson:1997:ESS

[PD97a] Tomas Philipson and Jeffrey DeSimone. Experiments and subject sampling. Biometrika, 84(3):619–630, September 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337583.

Pinheiro:1997:ERB

[PD97b] José C. Pinheiro and David L. DeMets. Estimating and reducing bias in group sequential designs with Gaussian independent increment structure. Biometrika, 84(4):831–845, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337655.

Palma:1999:SAI

[PD99] Wilfredo Palma and Guido Del Pino. Statistical analysis of incomplete long-range dependent data. Biometrika, 86(4):965– 972, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673601.

Paulino:1995:BMC

[PdBP95] Carlos Daniel Mimoso Paulino and Carlos Alberto de Braganca Pereira. Bayesian methods for categorical data under informative general censoring. Biometrika, 82(2):439–446, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337423.

Pearl:1995:CDE

[Pea95a]  Judea Pearl.  Causal diagrams for empirical research.

Biometrika, 82(4):669–688, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2337329. See discussion [CW95,

Daw95, FGS95, Fre95, IR95, Rob95, Ros95, Sha95, Sob95] and rejoinder [Pea95b].

Pearl:1995:RDC

[Pea95b] Judea Pearl. Rejoinder to discussions of “Causal Diagrams for Empirical Research”. Biometrika, 82(4):702–710, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337339.

See [Pea95a].

Pepe:1992:IUS

[Pep92] Margaret Sullivan Pepe. Inference using surrogate outcome data and a validation sample. Biometrika, 79(2):355–365, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336846.

Pepe:1997:RMF

[Pep97] Margaret Sullivan Pepe. A regression modelling framework for receiver operating characteristic curves in medical diagnostic testing. Biometrika, 84(3):595–608, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337581.

Percival:1995:EWV

[Per95] Donald P. Percival. On estimation of the wavelet variance. Biometrika, 82(3):619–631, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337538.

Perron:1999:BAR

[Per99] François Perron. Beyond accept-reject sampling. Biometrika, 86(4):803–813, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673586.

Pantula:1993:LSD

[PF93] Sastry G. Pantula and Wayne A. Fuller. The large sample distribution of the roots of the second order autoregressive polynomial. Biometrika, 80(4):919–923, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336885.

Pawitan:1991:EBC

[PG91] Y. Pawitan and A. K. Gangopadhyay. Efficient bias corrected nonparametric spectral estimation. Biometrika, 78(4):825–832, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336934.

Pena:1993:CPM

[PG93] Daniel Peña and Irwin Guttman. Comparing probabilistic methods for outlier detection in linear models. Biometrika, 80 (3):603–610, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337181.

Prentice:1997:RHR

[PH97] Ross L. Prentice and Li Hsu. Regression on hazard ratios and cross ratios in multivariate failure time analysis. Biometrika, 84(2):349–363, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337462.

Pigeot:1990:CAE

[Pig90] Iris Pigeot. A class of asymptotically efficient noniterative estimators of a common odds ratio. Biometrika, 77(2):420– 423, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336825.

Pfeffermann:1991:PUR

[PK91] Danny Pfeffermann and Abba M. Krieger. Poststratification using regression estimators when information on strata means and sizes is missing. Biometrika, 78(2):409–419, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337268.

Papageorgiou:1998:OSD

[PK98] Ioulia Papageorgiou and K. X. Karakostas. On optimal sampling designs for autocorrelated finite populations. Biometrika, 85(2):482–486, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337375.

Pukkila:1990:IAM

[PKK90] Tarmo Pukkila, Sergio Koreisha, and Arto Kallinen. The identification of ARMA models. Biometrika, 77(3):537–548, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336992.

Pocock:1990:DPC

[Poc90] S. J. Pocock. Discussion of paper by C. B. Begg. Biometrika, 77 (3):480–481, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336982. See [Beg90b].

Polson:1991:RPM

[Pol91] Nicholas G. Polson. A representation of the posterior mean for a location model. Biometrika, 78(2):426–430, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337270.

Poskitt:1987:MHR

[Pos87] D. S. Poskitt. A modified Hannan–Rissanen strategy for mixed autoregressive-moving average order determination. Biometrika, 74(4):781–790, December 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336472. See correction [Pos90].

Poskitt:1990:ACM

[Pos90] D. S. Poskitt. Amendments and corrections: “A modified Hannan–Rissanen strategy for mixed autoregressive-moving average order determination” [Biometrika 74 (1987), no. 4, 781– 790; MR0919846 (90f:62291)]. Biometrika, 77(3):667, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337010.

See [Pos87].

Pourahmadi:1999:JMC

[Pou99] Mohsen Pourahmadi. Joint mean-covariance models with applications to longitudinal data: Unconstrained parameterisation. Biometrika, 86(3):677–690, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673662.

Peddada:1992:JVE

[PP92] Shyamal Das Peddada and Girish Patwardhan. Jackknife variance estimators in linear models. Biometrika, 79(3):654–657, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336797.

Pierce:1994:HOA

[PP94] Donald A. Pierce and Dawn Peters. Higher-order asymptotics and the likelihood principle: One-parameter models. Biometrika, 81(1):1–10, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337045. See corrections [PP95].

Pierce:1995:ACH

[PP95] Donald A. Pierce and Dawn Peters. Amendments and corrections: “Higher-order asymptotics and the likelihood principle: one-parameter models” [Biometrika 81 (1994), no. 1, 1– 10; (95a:62011)]. Biometrika, 82(1):235, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337645. See [PP94].

Pierce:1999:IET

[PP99] Donald A. Pierce and Dawn Peters. Improving on exact tests by approximate conditioning. Biometrika, 86(2):265–277, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673510.

Preisser:1996:DDG

[PQ96] John S. Preisser and Bahjat F. Qaqish. Deletion diagnostics for generalised estimating equations. Biometrika, 83(3):551–562, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337508.

Pukelsheim:1992:ERA

[PR92] Friedrich Pukelsheim and Sabine Rieder. Efficient rounding of approximate designs. Biometrika, 79(4):763–770, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337232.

Pikounis:1994:LRT

[PR94a] V. Bill Pikounis and P. V. Rao. Linear rank tests for the nonresponders problem with censored data. Biometrika, 81(4): 691–699, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337072.

Polson:1994:BFD

[PR94b] Nicholas G. Polson and Gareth O. Roberts. Bayes factors for discrete observations from diffusion processes. Biometrika, 81 (1):11–26, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337046.

Proschan:1997:CPF

[Pro97] Michael A. Proschan. Conditional power with Fisher’s least significant difference procedure. Biometrika, 84(1):197–208, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337566.

Proschan:1999:PSF

[Pro99] Michael A. Proschan. Properties of spending function boundaries. Biometrika, 86(2):466–473, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673527.

Pericchi:1995:NBI

[PS95a] Luis R. Pericchi and Bruno Sansó. A note on bounded influence in Bayesian analysis. Biometrika, 82(1):223–225, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337642.

Prentice:1995:ADS

[PS95b] Ross L. Prentice and Lianne Sheppard. Aggregate data studies of disease risk factors. Biometrika, 82(1):113–125, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337632.

Pfanzagl:1996:SHP

[PS96] J. Pfanzagl and O. Sheynin. Studies in the history of probability and statistics. XLIV. A forerunner of the t-distribution. Biometrika, 83(4):891–898, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337291.

Penzer:1997:ELA

[PS97a] Jeremy Penzer and Brian Shea. The exact likelihood of an autoregressive-moving average model with incomplete data. Biometrika, 84(4):919–928, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337662.

Pickands:1997:EMG

[PS97b] James Pickands III and Robert A. Stine. Estimation for an M/G/∞ queue with incomplete information. Biometrika, 84 (2):295–308, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337458.

Paik:1997:UCP

[PT97] Myunghee Cho Paik and Wei-Yann Tsai. On using the Cox proportional hazards model with missing covariates. Biometrika, 84(3):579–593, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337580.

Polson:1990:PDB

[PW90] Nick Polson and Larry Wasserman. Prior distributions for the bivariate binomial. Biometrika, 77(4):901–904, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337113.

Pistone:1996:GCG

[PW96] Giovanni Pistone and Henry P. Wynn. Generalised confounding with Gröbner bases. Biometrika, 83(3):653–666, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337516.

Parzen:1994:RMB

[PWY94] M. I. Parzen, L. J. Wei, and Z. Ying. A resampling method based on pivotal estimating functions. Biometrika, 81(2):341– 350, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336964.

Pettit:1990:MEO

[PY90] L. I. Pettit and K. D. S. Young. Measuring the effect of observations on Bayes factors. Biometrika, 77(3):455–466, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336979.

Quinn:1991:FET

[QF91] B. G. Quinn and J. M. Fernandes. A fast efficient technique for the estimation of frequency. Biometrika, 78(3):489–497, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337018.

Qian:1996:GLM

[QGG96] Guoqi Qian, George Gabor, and R. P. Gupta. Generalised linear model selection by the predictive least quasi-deviance criterion. Biometrika, 83(1):41–54, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337431.

Qin:1998:ICC

[Qin98] Jing Qin. Inferences for case-control and semiparametric two-sample density ratio models. Biometrika, 85(3):619–630, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337391.

Quinn:1991:EFP

[QT91] B. G. Quinn and P. J. Thomson. Estimating the frequency of a periodic function. Biometrika, 78(1):65–74, March 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336896.

Quinn:1999:FET

[Qui99] B. G. Quinn. A fast efficient technique for the estimation of frequency: Interpretation and generalisation. Biometrika, 86 (1):213–220, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673549.

Qin:1997:GFT

[QZ97] Jing Qin and Biao Zhang. A goodness-of-fit test for logistic regression models based on case-control data. Biometrika, 84 (3):609–618, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337582.

Qaqish:1997:CCS

[QZC97] Bahjat F. Qaqish, Haibo Zhou, and Jianwen Cai. On casecontrol sampling of clustered data. Biometrika, 84(4):983–986, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337670.

Rabinowitz:1993:EDA

[Rab93] Daniel Rabinowitz. Estimating Durbin’s approximation. Biometrika, 80(3):671–680, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337186.

Rabinowitz:1997:NEE

[Rab97] Daniel Rabinowitz. A note on efficient estimation from casecontrol data. Biometrika, 84(2):486–488, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337474.

Raftery:1996:ABF

[Raf96] Adrian E. Raftery. Approximate Bayes factors and accounting for model uncertainty in generalised linear models. Biometrika, 83(2):251–266, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337598.

Rahiala:1999:RCA

[Rah99] Markku Rahiala. Random coefficient autoregressive models for longitudinal data. Biometrika, 86(3):718–722, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673666.

Reschenhofer:1991:LTG

[RB91] E. Reschenhofer and I. M. Bomze. Length tests for goodness of fit. Biometrika, 78(1):207–216, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336911. See correction [RB92].

Reschenhofer:1992:ACL

[RB92] E. Reschenhofer and I. M. Bomze. Amendments and corrections: “Length tests for goodness of fit” [Biometrika 78 (1991), no. 1, 207–216; (93a:62079)]. Biometrika, 79(4):859, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337245. See

[RB91].

Reeves:1991:ECV

[Ree91] Gillian K. Reeves. Estimation of contrast variances in linear models. Biometrika, 78(1):7–14, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336891.

Reeves:1993:GFT

[Ree93] Gillian K. Reeves. Goodness-of-fit tests in two-state processes. Biometrika, 80(2):431–442, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337212.

Roeder:1998:MHF

[REKB98] Kathryn Roeder, Michael Escobar, Joseph B. Kadane, and Ivan Balazs. Measuring heterogeneity in forensic databases using hierarchical Bayes models. Biometrika, 85(2):269–287, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337357.

Rue:1999:BOI

[RH99] Håvard Rue and Merrilee A. Hurn. Bayesian object identification. Biometrika, 86(3):649–660, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673660.

Raghunathan:1993:ABD

[RI93] T. E. Raghunathan and Yoichi Ii. Analysis of binary data from a multicentre clinical trial. Biometrika, 80(1):127–139, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336763.

Ringrose:1996:ACR

[Rin96] Trevor J. Ringrose. Alternative confidence regions for canonical variate analysis. Biometrika, 83(3):575–587, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337510.

Rivest:1994:SPW

[Riv94] Louis-Paul Rivest. Statistical properties of Winsorized means for skewed distributions. Biometrika, 81(2):373–383, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336967.

Rivest:1997:DPC

[Riv97] Louis-Paul Rivest. A decentred predictor for circular–circular regression. Biometrika, 84(3):717–726, September 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337591.

Rotnitzky:1990:HTR

[RJ90] Andrea Rotnitzky and Nicholas P. Jewell. Hypothesis testing of regression parameters in semiparametric generalized linear models for cluster correlated data. Biometrika, 77(3):485–497, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336986.

Rao:1990:EDF

[RKM90] J. N. K. Rao, J. G. Kovar, and H. J. Mantel. On estimating distribution functions and quantiles from survey data using auxiliary information. Biometrika, 77(2):365–375, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336815.

Rathouz:1999:RSN

[RL99] Paul J. Rathouz and Kung-Yee Liang. Reducing sensitivity to nuisance parameters in semiparametric models: A quasi-score method. Biometrika, 86(4):857–869, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673590.

Ramgopal:1993:NBB

[RLS93] P. Ramgopal, P. W. Laud, and A. F. M. Smith. Nonparametric Bayesian bioassay with prior constraints on the shape of the potency curve. Biometrika, 80(3):489–498, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337170. See

corrections [RLS94].

Ramgopal:1994:ACN

[RLS94] P. Ramgopal, P. W. Laud, and A. F. M. Smith. Amendments and corrections: “Nonparametric Bayesian bioassay with prior constraints on the shape of the potency curve” [Biometrika 80 (1993), no. 3, 489–498; (94g:62089)]. Biometrika, 81(3):631, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337137. See [RLS93].

Rao:1990:OCS

[RN90] J. N. K. Rao and A. K. Nigam. Optimal controlled sampling designs. Biometrika, 77(4):807–814, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337103.

Robins:1992:ETD

[Rob92] James Robins. Estimation of the time-dependent accelerated failure time model in the presence of confounding factors. Biometrika, 79(2):321–334, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336843.

Robins:1995:DCD

[Rob95] James M. Robins. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):695–698, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337335.

See [Pea95a].

Rocke:1986:ORS

[Roc86] David M. Rocke. Outlier resistance in small samples. Biometrika, 73(1):175–181, April 1986. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336283. See correction [Roc90].

Rocke:1990:ACO

[Roc90] David M. Rocke. Amendments and corrections: “Outlier resistance in small samples” [Biometrika 73 (1986), no. 1, 175– 181; MR0836445 (88b:62069)] by D. M. Rocke. Biometrika, 77 (1):235–236, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336073. See [Roc86].

Rom:1990:SRT

[Rom90] Dror M. Rom. A sequentially rejective test procedure based on a modified Bonferroni inequality. Biometrika, 77(3):663–665, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337008.

Rosenbaum:1992:DBC

[Ros92] Paul R. Rosenbaum. Detecting bias with confidence in observational studies. Biometrika, 79(2):367–374, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336847.

Rosenbaum:1995:DCD

[Ros95] Paul R. Rosenbaum. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):698–699, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337336.

See [Pea95a].

Reilly:1995:MSM

[RP95] Marie Reilly and Margaret Sullivan Pepe. A mean score method for missing and auxiliary covariate data in regression models. Biometrika, 82(2):299–314, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337409.

Rabinowitz:1993:LCB

[RR93] Daniel Rabinowitz and Louise Ryan. Lower confidence bounds for time to cure. Biometrika, 80(3):681–687, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337187.

Rotnitzky:1995:SRE

[RR95] Andrea Rotnitzky and James M. Robins. Semiparametric regression estimation in the presence of dependent censoring. Biometrika, 82(4):805–820, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337346.

Rao:1992:JVE

[RS92] J. N. K. Rao and J. Shao. Jackknife variance estimation with survey data under hot deck imputation. Biometrika, 79(4): 811–822, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337236.

Rao:1995:VEU

[RS95] J. N. K. Rao and R. R. Sitter. Variance estimation under two-phase sampling with application to imputation for missing data. Biometrika, 82(2):453–460, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337425.

Rao:1999:MBR

[RS99] J. N. K. Rao and Jun Shao. Modified balanced repeated replication for complex survey data. Biometrika, 86(2):403–415, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673521.

Robins:1992:SEA

[RT92a] James Robins and Anastasios A. Tsiatis. Semiparametric estimation of an accelerated failure time model with timedependent covariates. Biometrika, 79(2):311–319, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336842.

Rubin:1992:CEM

[RT92b] Donald B. Rubin and Neal Thomas. Characterizing the effect of matching using linear propensity score methods with normal distributions. Biometrika, 79(4):797–809, December 1992. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337235.

Roberts:1996:GCC

[RT96] G. O. Roberts and R. L. Tweedie. Geometric convergence and central limit theorems for multidimensional Hastings and Metropolis algorithms. Biometrika, 83(1):95–110, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337435.

Ray:1997:BSK

[RT97] Bonnie K. Ray and Ruey S. Tsay. Bandwidth selection for kernel regression with long-range dependent errors. Biometrika, 84(4): 791–802, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337652.

Rabinowitz:1995:RIC

[RTA95] Daniel Rabinowitz, Anastasios Tsiatis, and Jorge Aragon. Regression with interval-censored data. Biometrika, 82(3):501– 513, September 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337529.

Rudolfer:1990:MCM

[Rud90] Stephan M. Rudolfer. A Markov chain model of extrabinomial variation. Biometrika, 77(2):255–264, June 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336803. See correction

[Rud91].

Rudolfer:1991:CMC

[Rud91] Stephan M. Rudolfer. Corrections: “A Markov chain model of extrabinomial variation” [Biometrika 77 (1990), no. 2, 255–264; (91f:62028)]. Biometrika, 78(4):935, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336950. See [Rud90].

Russell:1991:CGC

[Rus91] K. G. Russell. The construction of good change-over designs when there are fewer units than treatments. Biometrika, 78 (2):305–313, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337257.

Ruiz-Velasco:1991:AEL

[RV91] S. Ruiz-Velasco. Asymptotic efficiency of logistic regression relative to linear discriminant analysis. Biometrika, 78(2):235– 243, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337248.

Roverato:1998:IMA

[RW98] Alberto Roverato and Joe Whittaker. The Isserlis matrix and its application to non-decomposable graphical Gaussian models. Biometrika, 85(3):711–725, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337398.

Robbins:1991:EMT

[RZ91] Herbert Robbins and Cun-Hui Zhang. Estimating a multiplicative treatment effect under biased allocation. Biometrika, 78 (2):349–354, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337261.

Stafford:1993:SAS

[SA93] James E. Stafford and David F. Andrews. A symbolic algorithm for studying adjustments to the profile likelihood. Biometrika, 80(4):715–730, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2336864.

Samuelsen:1997:PAA

[Sam97] Sven Ove Samuelsen. A pseudolikelihood approach to analysis of nested case-control studies. Biometrika, 84(2):379–394, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337464.

Sasieni:1996:PEH

[Sas96] Peter D. Sasieni. Proportional excess hazards. Biometrika, 83 (1):127–141, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337437.

Satten:1996:RBI

[Sat96] Glen A. Satten. Rank-based inference in the proportional hazards model for interval censored data. Biometrika, 83(2):355– 370, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337606.

Stewart:1991:SUO

[SB91] Frances P. Stewart and Ralph A. Bradley. Some universally optimal row-column designs with empty nodes. Biometrika, 78 (2):337–348, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337260.

Stefanski:1996:SED

[SB96] L. A. Stefanski and J. M. Bay. Simulation extrapolation deconvolution of finite population cumulative distribution function estimators. Biometrika, 83(2):407–417, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337610.

Sun:1998:RPP

[SB98] D. Sun and J. O. Berger. Reference priors with partial information. Biometrika, 85(1):55–71, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337309.

Sartori:1999:DMP

[SBSP99] N. Sartori, R. Bellio, A. Salvan, and L. Pace. The directed modified profile likelihood in models with many nuisance parameters. Biometrika, 86(3):735–742, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673669.

Solomon:1992:NCV

[SC92] P. J. Solomon and D. R. Cox. Nonlinear component of variance models. Biometrika, 79(1):1–11, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337142.

StLaurent:1993:LLI

[SC93] Roy T. St. Laurent and R. Dennis Cook. Leverage, local influence and curvature in nonlinear regression. Biometrika, 80 (1):99–106, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336760.

Samuel-Cahn:1996:SIB

[SC96] Ester Samuel-Cahn. Is the Simes improved Bonferroni procedure conservative? Biometrika, 83(4):928–933, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337297. See

correction [Ano04].

Schemper:1990:EVP

[Sch90a] M. Schemper. The explained variation in proportional hazards regression. Biometrika, 77(1):216–218, March 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336066. See correc-

tions [Sch94].

Schott:1990:CMP

[Sch90b] James R. Schott. Canonical mean projections and confidence regions in canonical variate analysis. Biometrika, 77(3):587–596, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336997.

Schall:1991:EGL

[Sch91a] Robert Schall. Estimation in generalized linear models with random effects. Biometrika, 78(4):719–727, December 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336923.

Schott:1991:STC

[Sch91b] James R. Schott. Some tests for common principal component subspaces in several groups. Biometrika, 78(4):771–777, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336928.

Schemper:1992:FRE

[Sch92] M. Schemper. Further results on the explained variation in proportional hazards regression. Biometrika, 79(1):202–204, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337162.

Schafer:1993:LAP

[Sch93] D. W. Schafer. Likelihood analysis for probit regression with measurement errors. Biometrika, 80(4):899–904, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336881.

Schemper:1994:ACE

[Sch94] M. Schemper. Amendments and corrections: “The explained variation in proportional hazards regression” [Biometrika 77 (1990), no. 1, 216–218; (91a:62272)]. Biometrika, 81(3):631, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337138. See [Sch90a].

Schott:1997:AEC

[Sch97] James R. Schott. Asymptotics of eigenprojections of correlation matrices with some applications in principal components analysis. Biometrika, 84(2):327–337, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337460.

Schott:1999:PCP

[Sch99] James R. Schott. Partial common principal component subspaces. Biometrika, 86(4):899–908, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673593.

Schall:1992:NRB

[SD92] Robert Schall and Timothy T. Dunne. A note on the relationship between parameter collinearity and local influence. Biometrika, 79(2):399–404, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336851.

Skouras:1999:EPF

[SD99a] Kostas Skouras and A. Philip Dawid. On efficient probability forecasting systems. Biometrika, 86(4):765–784, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673584.

Sutradhar:1999:ERE

[SD99b] Brajendra C. Sutradhar and Kalyan Das. On the efficiency of regression estimators in generalised linear models for longitudinal data. Biometrika, 86(2):459–465, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673526.

Samiuddin:1990:NKD

[SES90] M. Samiuddin and G. M. El-Sayyad. On nonparametric kernel density estimates. Biometrika, 77(4):865–874, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337109.

Severini:1990:CPL

[Sev90] Thomas A. Severini. Conditional properties of likelihood-based significance tests. Biometrika, 77(2):343–352, June 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336812.

Severini:1992:CRL

[Sev92] Thomas A. Severini. Conditional robustness in location estimation. Biometrika, 79(1):69–79, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337148.

Severini:1993:LAP

[Sev93] Thomas A. Severini. Local ancillarity in the presence of a nuisance parameter. Biometrika, 80(2):305–320, June 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337201.

Severini:1994:AEN

[Sev94] Thomas A. Severini. On the approximate elimination of nuisance parameters by conditioning. Biometrika, 81(4):649–661, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337068.

Severini:1998:AMP

[Sev98a] Thomas A. Severini. An approximation to the modified profile likelihood function. Biometrika, 85(2):403–411, June 1998. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337366.

Severini:1998:LFI

[Sev98b] Thomas A. Severini. Likelihood functions for inference in the presence of a nuisance parameter. Biometrika, 85(3):507–522, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337382.

Severini:1999:EAL

[Sev99] Thomas A. Severini. An empirical adjustment to the likelihood ratio statistic. Biometrika, 86(2):235–247, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673508.

Shi:1998:NTR

[SF98] Peide Shi and Wing-Kam Fung. A note on transforming a response variable for linearity. Biometrika, 85(3):749–754, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337401.

Shaw:1997:ETC

[SG97] Frank H. Shaw and Charles J. Geyer. Estimation and testing in constrained covariance component models. Biometrika, 84 (1):95–102, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337558.

Seifert:1993:NER

[SGW93] Burkhardt Seifert, Theo Gasser, and Andreas Wolf. Nonparametric estimation of residual variance revisited. Biometrika, 80 (2):373–383, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337207.

Sharples:1990:IAO

[Sha90] Linda D. Sharples. Identification and accommodation of outliers in general hierarchical models. Biometrika, 77(3):445–453, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336978.

Shafer:1995:DCD

[Sha95] Glenn Shafer. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):699–700, December 1995.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337337. See

[Pea95a].

Shephard:1994:PNG

[She94] Neil Shephard. Partial non-Gaussian state space. Biometrika, 81(1):115–131, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337055.

Shen:1998:POR

[She98] Xiaotong Shen. Proportional odds regression and sieve maximum likelihood estimation. Biometrika, 85(1):165–177, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337317.

Shi:1995:FIM

[Shi95] Daoji Shi. Fisher information for a multivariate extreme value distribution. Biometrika, 82(3):644–649, September 1995. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337541.

Shi:1997:LIP

[Shi97] Lei Shi. Local influence in principal components analysis. Biometrika, 84(1):175–186, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337564.

Shih:1998:GFT

[Shi98] Joanna H. Shih. A goodness-of-fit test for association in a bivariate survival model. Biometrika, 85(1):189–200, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337319.

Signorini:1991:SSP

[Sig91] David F. Signorini. Sample size for Poisson regression. Biometrika, 78(2):446–450, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337274.

Silvapulle:1992:RTI

[Sil92] Mervyn J. Silvapulle. Robust tests of inequality constraints and one-sided hypotheses in the linear model. Biometrika, 79(3):

621–630, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336793.

Silliman:1997:NCW

[Sil97] Nancy Paul Silliman. Nonparametric classes of weight functions to model publication bias. Biometrika, 84(4):909–918, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337661.

Sitter:1993:BRR

[Sit93] R. R. Sitter. Balanced repeated replications based on orthogonal multi-arrays. Biometrika, 80(1):211–221, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336770.

Schill:1993:LAC

[SJDT93] W. Schill, K.-H. Jöckel, K. Drescher, and J. Timm. Logistic analysis in case-control studies under validation sampling. Biometrika, 80(2):339–352, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337204.

Slud:1997:STS

[SK97] Eric V. Slud and Edward L. Korn. Semiparametric two-sample tests in clinical trials with a post-randomisation response indicator. Biometrika, 84(1):221–230, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337568.

Skates:1993:SAC

[Ska93] Steven James Skates. On secant approximations to cumulative distribution functions. Biometrika, 80(1):223–235, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336771.

Spiekerman:1996:CMC

[SL96] C. F. Spiekerman and D. Y. Lin. Checking the marginal Cox model for correlated failure time data. Biometrika, 83(1):143– 156, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337438.

Slate:1999:AMN

[Sla99] Elizabeth H. Slate. Assessing multivariate nonnormality using univariate distributions. Biometrika, 86(1):191–202, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673547.

Slud:1991:REL

[Slu91] Eric Slud. Relative efficiency of the log rank test within a multiplicative intensity model. Biometrika, 78(3):621–630, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337031.

Small:1993:NNS

[SM93] Christopher G. Small and D. J. Murdoch. Nonparametric Neyman–Scott problems: telescoping product methods. Biometrika, 80(4):763–779, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336868.

Smith:1991:BAM

[Smi91] Philip J. Smith. Bayesian analyses for a multiple capturerecapture model. Biometrika, 78(2):399–407, June 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337267.

Smith:1994:NR

[Smi94] Richard L. Smith. Nonregular regression. Biometrika, 81(1): 173–183, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337059.

Smyth:1994:NMC

[Smy94] Gordon K. Smyth. A note on modelling cross-correlations: Hyperbolic secant regression. Biometrika, 81(2):396–402, June 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336970.

Shen:1996:CCR

[SO96] Larry Z. Shen and John O’Quigley. Consistency of continual reassessment method under model misspecification. Biometrika, 83(2):395–405, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337609.

Shoji:1998:SME

[SO98] Isao Shoji and Tohru Ozaki. A statistical method of estimation and simulation for systems of stochastic differential equations. Biometrika, 85(1):240–243, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337324.

So:1999:TSA

[So99] Mike K. P. So. Time series with additive noise. Biometrika, 86 (2):474–482, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673528.

Sobel:1995:DCD

[Sob95] Michael E. Sobel. Discussion of “Causal Diagrams for Empirical Research” by J. Pearl. Biometrika, 82(4):700–702, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337338.

See [Pea95a].

Schafer:1996:LAE

[SP96] Daniel W. Schafer and Kathleen G. Purdy. Likelihood analysis for errors-in-variables regression with replicate measurements. Biometrika, 83(4):813–824, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337285.

Shephard:1997:LAN

[SP97] Neil Shephard and Michael K. Pitt. Likelihood analysis of non-Gaussian measurement time series. Biometrika, 84(3):653–667, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337586.

Spady:1991:SAR

[Spa91] Richard H. Spady. Saddlepoint approximations for regression models. Biometrika, 78(4):879–889, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336940.

Shin:1996:TOG

[SPP96] Dong Wan Shin, Chul Gyu Park, and Taesung Park. Testing for ordered group effects with repeated measurements. Biometrika, 83(3):688–694, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337520.

Srivenkataramana:1991:ATP

[ST91] T. Srivenkataramana and D. S. Tracy. Amendment: “Twophase sampling for selection with probability proportional to size in sample surveys” [Biometrika 76 (1989), no. 4, 818–821; MR1041430 (91c:62009)]. Biometrika, 78(1):233, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336917.

Skaug:1993:NTS

[ST93] Hans Julius Skaug and Dag Tjøstheim. A nonparametric test of serial independence based on the empirical distribution function. Biometrika, 80(3):591–602, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337180.

Stoffer:1998:MSC

[ST98] David S. Stoffer and David E. Tyler. Matching sequences: Cross-spectral analysis of categorical time series. Biometrika, 85 (1):201–213, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337320.

Solomon:1999:OTV

[ST99] P. J. Solomon and J. M. G. Taylor. Orthogonality and transformations in variance components models. Biometrika, 86(2):289– 300, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673512.

Stanghellini:1997:ISF

[Sta97] Elena Stanghellini. Identification of a single-factor model using graphical Gaussian rules. Biometrika, 84(1):241–244, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337571.

Smith:1997:MCM

[STC97] Richard L. Smith, Jonathan A. Tawn, and Stuart G. Coles. Markov chain models for threshold exceedances. Biometrika, 84 (2):249–268, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337455.

Stenger:1990:AMR

[Ste90a] H. Stenger. Asymptotic minimaxity of the ratio strategy. Biometrika, 77(2):389–395, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336818.

Stern:1990:CPC

[Ste90b] Hal Stern. A continuum of paired comparisons models. Biometrika, 77(2):265–273, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336804.

Stein:1991:NCE

[Ste91] Michael L. Stein. A new class of estimators for the reduced second moment measure of point processes. Biometrika, 78 (2):281–286, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337253.

Stein:1993:AOE

[Ste93] Michael L. Stein. Asymptotically optimal estimation for the reduced second moment measure of point processes. Biometrika, 80(2):443–449, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337213.

Sinha:1994:MML

[STH94] Debajyoti Sinha, Martin A. Tanner, and W. J. Hall. Maximization of the marginal likelihood of grouped survival data. Biometrika, 81(1):53–60, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337049.

Stigler:1992:SHP

[Sti92] Stephen M. Stigler. Studies in the history of probability and statistics. XLIII. Karl Pearson and quasi-independence.

Biometrika, 79(3):563–575, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336787.

Stoffer:1993:SAC

[STM93] David S. Stoffer, David E. Tyler, and Andrew J. McDougall. Spectral analysis for categorical time series: Scaling and the spectral envelope. Biometrika, 80(3):611–622, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337182.

Sun:1999:PDH

[STS99] Dongchu Sun, Robert K. Tsutakawa, and Paul L. Speckman. Posterior distribution of hierarchical models using CAR(1) distributions. Biometrika, 86(2):341–350, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673516.

Sugden:1993:PES

[Sug93] R. A. Sugden. Partial exchangeability and survey sampling inference. Biometrika, 80(2):451–455, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337214.

Sun:1991:SLE

[Sun91] J. Sun. Significance levels in exploratory projection pursuit. Biometrika, 78(4):759–769, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336927.

Sundberg:1994:PES

[Sun94] Rolf Sundberg. Precision estimation in sample survey inference: A criterion for choice between variance estimators. Biometrika, 81(1):157–172, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337058.

Scott:1991:FMD

[SW91] David W. Scott and M. P. Wand. Feasibility of multivariate density estimates. Biometrika, 78(1):197–205, March 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336910.

Stute:1994:JEK

[SW94] Winfried Stute and Jane-Ling Wang. The jackknife estimate of a Kaplan–Meier integral. Biometrika, 81(3):602–606, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337132.

Scott:1997:FRM

[SW97] A. J. Scott and C. J. Wild. Fitting regression models to casecontrol data by maximum likelihood. Biometrika, 84(1):57– 71, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337555.

Sweeting:1992:PBA

[Swe92] Trevor J. Sweeting. Parameter-based asymptotics. Biometrika, 79(2):219–230, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336834.

Sweeting:1995:BAA

[Swe95a] Trevor J. Sweeting. A Bayesian approach to approximate conditional inference. Biometrika, 82(1):25–36, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337624.

Sweeting:1995:FBL

[Swe95b] Trevor J. Sweeting. A framework for Bayesian and likelihood approximations in statistics. Biometrika, 82(1):1–23, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337623.

Schaid:1990:OTS

[SWT90] Daniel J. Schaid, Sam Wieand, and Terry M. Therneau. Optimal two-stage screening designs for survival comparisons. Biometrika, 77(3):507–513, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336988.

Sun:1996:FVP

[SY96] Dongchu Sun and Keying Ye. Frequentist validity of posterior quantiles for a two-parameter exponential family. Biometrika, 83(1):55–65, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337432.

Smith:1995:RSC

[SZ95] Peter J. Smith and J. Zhang. Renovated scatterplots for censored data. Biometrika, 82(2):447–452, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337424.

Shen:1999:PEP

[SZ99] Yu Shen and Marvin Zelen. Parametric estimation procedures for screening programmes: Stable and nonstable disease models for multimodality case finding. Biometrika, 86(3):503–515, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673650.

Tarone:1979:TGF

[Tar79] R. E. Tarone. Testing the goodness of fit of the binomial distribution. Biometrika, 66(3):585–590, December 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335179. See correction

[Tar90].

Tarone:1990:ACT

[Tar90] R. E. Tarone. Amendments and corrections: “Testing the Goodness of fit of the Binomial Distribution”. Biometrika, 77(3):668, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337011. See [Wis68, Tar79].

Tawn:1990:MME

[Taw90] Jonathan A. Tawn. Modelling multivariate extreme value distributions. Biometrika, 77(2):245–253, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336802.

Tsiatis:1995:SMP

[TBK95] Anastasios A. Tsiatis, Hélène Boucher, and Kyungmann Kim. Sequential methods for parametric survival models. Biometrika, 82(1):165–173, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337636.

Tsai:1998:NNE

[TC98] Wei-Yann Tsai and John Crowley. A note on nonparametric estimators of the bivariate survival function under univariate censoring. Biometrika, 85(3):573–580, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337387.

Therneau:1990:MBR

[TGF90] Terry M. Therneau, Patricia M. Grambsch, and Thomas R. Fleming. Martingale-based residuals for survival models. Biometrika, 77(1):147–160, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336057.

Thompson:1991:NRT

[Tho91a] G. L. Thompson. A note on the rank transform for interactions. Biometrika, 78(3):697–701, September 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337040. See correction

[Tho93].

Thompson:1991:SAC

[Tho91b] Steven K. Thompson. Stratified adaptive cluster sampling. Biometrika, 78(2):389–397, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337266.

Thompson:1993:ACN

[Tho93] Georgia Lee Thompson. Amendments and corrections: “A note on the rank transform for interactions” [Biometrika 78 (1991), no. 3, 697–701; (92k:62127)]. Biometrika, 80(3):711, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337192.

See [Tho91a].

Tille:1996:EPU

[Til96] Yves Tillé. An elimination procedure for unequal probability sampling without replacement. Biometrika, 83(1):238–241, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337447.

Tierney:1989:AMD

[TKK89] Luke Tierney, Robert E. Kass, and Joseph B. Kadane. Approximate marginal densities of nonlinear functions. Biometrika, 76 (3):425–433, September 1989. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2336109. See correction [TKK91].

Tierney:1991:ACA

[TKK91] Luke Tierney, Robert E. Kass, and Joseph B. Kadane. Amendments and corrections: “Approximate marginal densities of nonlinear functions” [Biometrika 76 (1989), no. 3, 425–433; MR1040637 (91c:62015)]. Biometrika, 78(1):233–234, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336918.

See [TKK89].

Thompson:1991:NES

[TKT91] A. M. Thompson, J. W. Kay, and D. M. Titterington. Noise estimation in signal restoration using regularization. Biometrika, 78(3):475–488, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337017.

Troxel:1998:MMA

[TLH98] Andrea B. Troxel, Stuart R. Lipsitz, and David P. Harrington. Marginal models for the analysis of longitudinal measurements with nonignorable non-monotone missing data. Biometrika, 85 (3):661–672, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337394.

Tu:1995:IAP

[TLP95] Xin Ming Tu, Eugene Litvak, and Marcello Pagano. On the informativeness and accuracy of pooled testing in estimating prevalence of a rare disease: Application to HIV screening. Biometrika, 82(2):287–297, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337408.

Tsai:1990:TAI

[Tsa90] Wei-Yann Tsai. Testing the assumption of independence of truncation time and failure time. Biometrika, 77(1):169–177, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336059.

Tsai:1994:NJI

[Tsa94] Chih-Ling Tsai. A note on Jorgensen’s iteratively defined statistics. Biometrika, 81(4):781–786, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337081.

Tosteson:1990:DLR

[TW90] Tor D. Tosteson and James H. Ware. Designing a logistic regression study using surrogate measures for exposure and outcome. Biometrika, 77(1):11–21, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336045.

Todd:1997:CIC

[TW97] Susan Todd and John Whitehead. Confidence interval calculation for a sequential clinical trial of binary responses. Biometrika, 84(3):737–743, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337594.

Todd:1996:PIE

[TWF96] Susan Todd, John Whitehead, and Karen M. Facey. Point and interval estimation following a sequential clinical trial. Biometrika, 83(2):453–461, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337615.

Tiao:1993:RML

[TX93] George C. Tiao and Daming Xu. Robustness of maximum likelihood estimates for multi-step predictions: The exponential smoothing case. Biometrika, 80(3):623–641, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337183.

Uddin:1990:SSC

[Udd90] Nizam Uddin. Some series constructions for minimal size equineighboured balanced incomplete block designs with nested rows and columns. Biometrika, 77(4):829–833, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337105.

Uddin:1990:SCB

[UM90] Nizam Uddin and John P. Morgan. Some constructions for balanced incomplete block designs with nested rows and columns. Biometrika, 77(1):193–202, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336062.

Uddin:1997:EBD

[UM97] Nizam Uddin and John P. Morgan. Efficient block designs for settings with spatially correlated errors. Biometrika, 84(2):443– 454, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337469. See corrections [UM99].

Uddin:1999:ACE

[UM99] N. Uddin and J. P. Morgan. Amendments and corrections:“Efficient block designs for settings with spatially correlated errors” [Biometrika 84 (1997), no. 2, 443–454; (99c:62212)]. Biometrika, 86(1):233, 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). See [UM97].

Uzunogullari:1992:CHR

[UW92] Ülkü Uzunoğullari and Jane-Ling Wang. A comparison of hazard rate estimators for left truncated and right censored data. Biometrika, 79(2):297–310, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336841.

Vecchia:1991:TPA

[VB91] A. V. Vecchia and R. Ballerini. Testing for periodic autocorrelations in seasonal time series data. Biometrika, 78(1):53– 63, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336895.

Venkatraman:1996:DFP

[VB96] E. S. Venkatraman and Colin B. Begg. A distribution-free procedure for comparing receiver operating characteristic curves from a paired experiment. Biometrika, 83(4):835–848, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337287.

VanEs:1990:KEW

[VH90] Bert Van Es and Adriaan Hoogendoorn. Kernel estimation in Wicksell’s corpuscle problem. Biometrika, 77(1):139–145, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336056.

VanderLaan:1998:LEE

[VH98] Mark J. Van der Laan and Alan E. Hubbard. Locally efficient estimation of the survival distribution with right-censored data and covariates when collection of data is delayed. Biometrika, 85(4):771–783, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337482.

Vidoni:1995:SPD

[Vid95] Paolo Vidoni. A simple predictive density based on the p^{∗}- formula. Biometrika, 82(4):855–863, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337350.

Vidoni:1998:NME

[Vid98] Paolo Vidoni. A note on modified estimative prediction limits and distributions. Biometrika, 85(4):949–953, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337495.

Visser:1996:NEB

[Vis96] Michael Visser. Nonparametric estimation of the bivariate survival function with an application to vertically transmitted AIDS. Biometrika, 83(3):507–518, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337504.

VanDerLaan:1997:EEL

[VJP97] Mark J. Van Der Laan, Nicholas P. Jewell, and Derick R. Peterson. Efficient estimation of the lifetime and disease onset distribution. Biometrika, 84(3):539–554, September 1997. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337577.

Vlachonikolis:1990:PDC

[Vla90] I. G. Vlachonikolis. Predictive discrimination and classification with mixed binary and continuous variables. Biometrika, 77(3): 657–662, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337007.

Vonesh:1996:NUL

[Von96] Edward F. Vonesh. A note on the use of Laplace’s approximation for nonlinear mixed-effects models. Biometrika, 83(2):447– 452, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337614.

Velu:1990:EPO

[VR90] Raja Velu and M. Bhaskara Rao. Estimation of parent-offspring correlation. Biometrika, 77(3):557–562, September 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336994.

Vach:1993:LRI

[VS93] Werner Vach and Martin Schumacher. Logistic regression with incompletely observed categorical covariates: A comparison of three approaches. Biometrika, 80(2):353–362, June 1993. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337205.

Venter:1995:NGC

[VS95] J. H. Venter and J. L. J. Snyman. A note on the generalised cross-validation criterion in linear model selection. Biometrika, 82(1):215–219, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337640.

VanDerMeulen:1996:THM

[VV96] E. A. Van Der Meulen and N. J. Van Sittert. Testing homogeneity of many within-subject variances using small samples. Biometrika, 83(3):683–687, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337519.

Wakaki:1990:CLQ

[Wak90] Hirofumi Wakaki. Comparison of linear and quadratic discriminant functions. Biometrika, 77(1):227–229, March 1990. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336069.

Walden:1992:APP

[Wal92] A. T. Walden. Asymptotic percentage points for Siegel’s test statistic for compound periodicities. Biometrika, 79(2):438– 440, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336860.

Wang:1991:GIC

[Wan91] You-Gan Wang. Gittins indices and constrained allocation in clinical trials. Biometrika, 78(1):101–111, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336900.

Wang:1992:TPA

[Wan92] Suojin Wang. Tail probability approximations in the first-order noncircular autoregression. Biometrika, 79(2):431–434, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336858.

Wang:1993:SEF

[Wan93] Suojin Wang. Saddlepoint expansions in finite population problems. Biometrika, 80(3):583–590, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337179.

Wang:1995:JSC

[Wan95] Yazhen Wang. Jump and sharp cusp detection by wavelets. Biometrika, 82(2):385–397, June 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337415.

Wang:1996:HRA

[Wan96] Mei-Cheng Wang. Hazards regression analysis for lengthbiased data. Biometrika, 83(2):343–354, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337605.

Wand:1999:OAS

[Wan99] M. P. Wand. On the optimal amount of smoothing in penalised spline regression. Biometrika, 86(4):936–940, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673597.

Watkins:1990:PEL

[Wat90] A. J. Watkins. On polynomial expansions for likelihoods of spatial data with finite range correlation function. Biometrika, 77(2):404–408, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336821.

Weiss:1992:GCS

[WC92a] Robert E. Weiss and R. Dennis Cook. A graphical case statistic for assessing posterior influence. Biometrika, 79(1):51–55, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337146.

Wilson:1992:CRE

[WC92b] Richard M. Wilson and Mark F. Collins. Capture-recapture estimation with samples of size one using frequency data. Biometrika, 79(3):543–553, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336785.

Wang:1993:REL

[WC93] C. Y. Wang and R. J. Carroll. On robust estimation in logistic case-control studies. Biometrika, 80(1):237–241, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336772.

Wang:1999:HOA

[WC99a] Suojin Wang and Raymond J. Carroll. High-order accurate methods for retrospective sampling problems. Biometrika, 86 (4):881–897, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673592.

Wikle:1999:DRA

[WC99b] Christopher K. Wikle and Noel Cressie. A dimension-reduced approach to space-time Kalman filtering. Biometrika, 86(4):

815–829, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673587.

Wang:1996:NCM

[WD96a] Naisyin Wang and Marie Davidian. A note on covariate measurement error in nonlinear mixed effects models. Biometrika, 83(4):801–812, December 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337284.

Wang:1996:NEF

[WD96b] Suojin Wang and Alan H. Dorfman. A new estimator for the finite population distribution function. Biometrika, 83(3):639– 652, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337515.

Wells:1994:NKE

[Wel94] Martin T. Wells. Nonparametric kernel estimation in counting processes with explanatory variables. Biometrika, 81(4): 795–801, December 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337084.

West:1991:KDE

[Wes91] Mike West. Kernel density estimation and marginalization consistency. Biometrika, 78(2):421–425, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337269.

West:1997:TSD

[Wes97] Mike West. Time series decomposition. Biometrika, 84(2):489– 494, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337475.

Wellman:1991:IDL

[WG91] Joyce M. Wellman and Richard F. Gunst. Influence diagnostics for linear measurement error models. Biometrika, 78(2):373– 380, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337264.

Wagner:1995:MEM

[WG95] Udo Wagner and Alois L. J. Geyer. A maximum entropy method for inverting Laplace transforms of probability density functions. Biometrika, 82(4):887–892, December 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337353.

Wang:1999:STM

[WHD99] Weizhen Wang, J. T. Gene Hwang, and Anirban Dasgupta. Statistical tests for multivariate bioequivalence. Biometrika, 86 (2):395–402, June 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673520.

Whittemore:1995:LRF

[Whi95] Alice S. Whittemore. Logistic regression of family data from case-control studies. Biometrika, 82(1):57–67, March 1995. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337627. See correction

[Whi97].

Whittemore:1997:ACL

[Whi97] Alice S. Whittemore. Amendments and corrections: “Logistic Regression of Family Data From Case-Control Studies. Biometrika, 84(4):989–990, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337673. See [Whi95].

Wolpert:1998:PGR

[WI98] Robert L. Wolpert and Katja Ickstadt. Poisson/gamma random field models for spatial statistics. Biometrika, 85(2):251– 267, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337356.

Willmot:1990:CPO

[Wil90] Gordon E. Willmot. On the construction of a parameter orthogonal to the mean. Biometrika, 77(2):424–428, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336826.

Wild:1991:FPR

[Wil91] C. J. Wild. Fitting prospective regression models to case-control data. Biometrika, 78(4):705–717, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336922.

Wisniewski:1968:THB

[Wis68] T. K. M. Wisniewski. Testing for homogeneity of a binomial series. Biometrika, 55(2):426–428, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334888. See correc-

tions [Tar90, Wis90].

Wisniewski:1990:ACT

[Wis90] T. K. M. Wisniewski. Amendments and corrections: Testing for Homogeneity of a Binomial Series. Biometrika, 77(3):668, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2337012. See [Wis68].

Williams:1996:NOL

[WJ96] E. R. Williams and J. A. John. A note on optimality in lattice square designs. Biometrika, 83(3):709–713, September 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337523.

Westfall:1997:BPB

[WJU97] Peter H. Westfall, Wesley O. Johnson, and Jessica M. Utts. A Bayesian perspective on the Bonferroni adjustment. Biometrika, 84(2):419–427, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337467.

Wong:1992:LEP

[WL92] Wing Hung Wong and Bing Li. Laplace expansion for posterior densities of nonlinear functions of parameters. Biometrika, 79 (2):393–398, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336850.

Waterman:1996:PSM

[WL96] Richard P. Waterman and Bruce G. Lindsay. Projected score methods for approximating conditional scores. Biometrika, 83 (1):1–13, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337428.

Wong:1997:TTA

[WL97a] C. S. Wong and W. K. Li. Testing for threshold autoregression with conditional heteroscedasticity. Biometrika, 84(2):407– 418, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337466.

Wong:1997:MCH

[WL97b] Heung Wong and W. K. Li. On a multivariate conditional heteroscedastic model. Biometrika, 84(1):111–123, March 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337560.

Walden:1995:EBM

[WMP95] A. T. Walden, E. J. McCoy, and D. B. Percival. The effective bandwidth of a multitaper spectral estimator. Biometrika, 82 (1):201–214, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337639.

Wallenstein:1994:PSS

[WNG94] Sylvan Wallenstein, Joseph Naus, and Joseph Glaz. Power of the scan statistic in detecting a changed segment in a Bernoulli sequence. Biometrika, 81(3):595–601, August 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337131.

Wolff:1990:NBC

[Wol90] Rodney C. L. Wolff. A note on the behaviour of the correlation integral in the presence of a time series. Biometrika, 77(4): 689–697, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337092.

Wolfinger:1993:LAN

[Wol93] Russ Wolfinger. Laplace’s approximation for nonlinear mixed models. Biometrika, 80(4):791–795, December 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336870.

Wong:1991:BAL

[Won91] M. Y. Wong. Bartlett adjustment to the likelihood ratio statistic for testing several slopes. Biometrika, 78(1):221–224, March 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2336913.

Wong:1992:UAC

[Won92] Weng Kee Wong. A unified approach to the construction of minimax designs. Biometrika, 79(3):611–619, September 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336792.

Woodroofe:1992:EAS

[Woo92] Michael Woodroofe. Estimation after sequential testing: A simple approach for a truncated sequential probability ratio test. Biometrika, 79(2):347–353, June 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336845.

Williams:1992:ESC

[WP92] Paige L. Williams and Christopher J. Portier. Explicit solutions for constrained maximum likelihood estimators in survival/sacrifice experiments. Biometrika, 79(4):717–729, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337227.

Wang:1998:LST

[WR98] Naisyin Wang and James M. Robins. Large-sample theory for parametric multiple imputation procedures. Biometrika, 85(4): 935–948, December 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337494.

Wright:1992:NSL

[Wri92] Tommy Wright. A note on sampling to locate rare defectives with strong prior evidence. Biometrika, 79(4):685–691, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337224.

Willmot:1994:NPO

[WS94] G. E. Willmot and D. A. Sprott. A note on parameters orthogonal to the mean. Biometrika, 81(2):409–412, June 1994.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2336972.

Walker:1999:MFD

[WS99] Stephen G. Walker and David A. Stephens. A multivariate family of distributions on (0, ∞)^{p}. Biometrika, 86(3):703–709, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673664.

Wei:1990:IAR

[WSL90] L. J. Wei, John Q. Su, and John M. Lachin. Interim analyses with repeated measurements in a sequential clinical trial. Biometrika, 77(2):359–364, June 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2336814.

Wei:1990:CCB

[WT90] Greg C. G. Wei and Martin A. Tanner. Calculating the content and boundary of the highest posterior density region via data augmentation. Biometrika, 77(3):649–652, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337005.

Wu:1991:BRR

[Wu91] C. F. J. Wu. Balanced repeated replications based on mixed orthogonal arrays. Biometrika, 78(1):181–188, March 1991. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2336908.

Wu:1993:CSD

[Wu93] C. F. J. Wu. Construction of supersaturated designs through partially aliased interactions. Biometrika, 80(3):661–669, September 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337185.

Walden:1990:ESB

[WW90] A. T. Walden and R. E. White. Estimating the statistical bandwidth of a time series. Biometrika, 77(4):699–707, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337093.

Weinberg:1993:PAC

[WW93] Clarice R. Weinberg and Sholom Wacholder. Prospective analysis of case-control data under general multiplicative-intercept risk models. Biometrika, 80(2):461–465, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337216.

Wang:1997:NEB

[WW97] Weijing Wang and Martin T. Wells. Nonparametric estimators of the bivariate survival function under simplified censoring conditions. Biometrika, 84(4):863–880, December 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337657.

Wang:1998:NES

[WW98] Weijing Wang and Martin T. Wells. Nonparametric estimation of successive duration times under dependent censoring. Biometrika, 85(3):561–572, September 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337386.

Wei:1990:LRA

[WYL90] L. J. Wei, Z. Ying, and D. Y. Lin. Linear regression analysis of censored survival data based on rank tests. Biometrika, 77(4): 845–851, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337107.

Wu:1993:MAD

[WZ93] C.-F. J. Wu and Run Chu Zhang. Minimum aberration designs with two-level and four-level factors. Biometrika, 80(1):203– 209, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2336769.

Xiang:1999:CCC

[XL99] Anny Hui Xiang and Bryan Langholz. Comparison of casecontrol to full cohort analyses under model misspecification. Biometrika, 86(1):221–226, March 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2673550.

Xia:1999:EPL

[XTL99] Yingcun Xia, Howell Tong, and W. K. Li. On extended partially linear single-index models. Biometrika, 86(4):831–842, December 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673588.

Yao:1993:TCP

[Yao93] Qi Wei Yao. Tests for change-points with epidemic alternatives. Biometrika, 80(1):179–191, March 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336767.

Yatracos:1991:NTP

[Yat91] Yannis G. Yatracos. A note on Tukey’s polyefficiency. Biometrika, 78(3):702–703, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337041.

Ye:1991:NPI

[YB91] Keying Ye and James O. Berger. Noninformative priors for inferences in exponential regression models. Biometrika, 78(3): 645–656, September 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337034.

Yuen:1997:TFS

[YB97] K. C. Yuen and M. D. Burke. A test of fit for a semiparametric additive risk model. Biometrika, 84(3):631–639, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337584.

Young:1990:BB

[YD90] G. A. Young and H. E. Daniels. Bootstrap bias. Biometrika, 77 (1):179–185, March 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2336060.

Yip:1996:ICR

[YHL96] Paul S. F. Yip, Richard M. Huggins, and D. Y. Lin. Inference for capture-recapture experiments in continuous time with variable capture rates. Biometrika, 83(2):477–483, June 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337619.

Ying:1990:LRS

[Yin90] Zhiliang Ying. Linear rank statistics for truncated data. Biometrika, 77(4):909–914, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337115.

Yao:1998:AIR

[YWH98] Q. Yao, L. J. Wei, and J. W. Hogan. Analysis of incomplete repeated measurements with dependent censoring times. Biometrika, 85(1):139–149, March 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337315.

Ying:1992:PSP

[YWL92] Z. Ying, L. J. Wei, and J. S. Lin. Prediction of survival probability based on a linear regression model. Biometrika, 79(1):205– 209, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337163.

Zelen:1993:OSE

[Zel93] Marvin Zelen. Optimal scheduling of examinations for the early detection of disease. Biometrika, 80(2):279–293, June 1993. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337199. See

corrections [Zel96].

Zelen:1996:ACO

[Zel96] Marvin Zelen. Amendments and corrections: “Optimal scheduling of examinations for the early detection of disease” [Biometrika 80 (1993), no. 2, 279–293; (94e:62111)]. Biometrika, 83(1):249, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2337452. See [Zel93].

Zelen:1991:THB

[ZH91] M. Zelen and Y. Haitovsky. Testing hypotheses with binary data subject to misclassification errors: Analysis and experimental design. Biometrika, 78(4):857–865, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2336938.

Zhang:1992:IAV

[Zha92] Ping Zhang. Inference after variable selection in linear regression models. Biometrika, 79(4):741–746, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337229.

Zhao:1994:SAH

[Zha94] Lue Ping Zhao. Segregation analysis of human pedigrees using estimating equations. Biometrika, 81(1):197–209, March 1994. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2337061.

Zhang:1999:CSG

[Zha99] Biao Zhang. A chi-squared goodness-of-fit test for logistic regression models based on case-control data. Biometrika, 86(3): 531–539, September 1999. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673652.

Zhao:1998:CAA

[ZHH^{+}98] Lue Ping Zhao, Li Hsu, Sarah Holte, Yan Chen, Filemon Quiaoit, and Ross L. Prentice. Combined association and aggregation analysis of data from case-control family studies. Biometrika, 85(2):299–315, June 1998. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337359.

Zhou:1992:ECL

[Zho92] Mai Zhou. M-estimation in censored linear models. Biometrika, 79(4):837–841, December 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2337240.

Zimmerman:1991:CDB

[Zim91] Dale Zimmerman. Censored distance-based intensity estimation of spatial point processes. Biometrika, 78(2):287–294, June 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337254.

Zheng:1995:EMS

[ZK95] Ming Zheng and John P. Klein. Estimates of marginal survival for dependent competing risks based on an assumed copula.

Biometrika, 82(1):127–138, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2337633.

Zucker:1990:WLR

[ZL90] David M. Zucker and Edward Lakatos. Weighted log rank type statistics for comparing survival curves when there is a time lag in the effectiveness of treatment. Biometrika, 77(4):853–864, December 1990. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337108.

Zhao:1990:CBR

[ZP90] Lue Ping Zhao and Ross L. Prentice. Correlated binary regression using a quadratic exponential model. Biometrika, 77(3): 642–648, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2337004.

Zhou:1995:ACD

[ZP95] Haibo Zhou and Margaret Sullivan Pepe. Auxiliary covariate data in failure time regression. Biometrika, 82(1):139– 149, March 1995. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2337634.

Zhang:1991:RMT

[ZPB91] Ji Zhang, Sastry G. Pantula, and Dennis D. Boos. Robust methods for testing the pattern of a single covariance matrix. Biometrika, 78(4):787–795, December 1991. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2336930.

Zhao:1997:CED

[ZT97] Hongwei Zhao and Anastasios A. Tsiatis. A consistent estimator for the distribution of quality adjusted survival time. Biometrika, 84(2):339–348, June 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/www.jstor.org/stable/2337461.